### v1

# Universal Time-Series Representation Learning: A Survey

Abstract.
Time-series data exists in every corner of real-world systems and services, ranging from satellites in the sky to wearable devices on human bodies. Learning representations by extracting and inferring valuable information from these time series is crucial for understanding the complex dynamics of particular phenomena and enabling informed decisions. With the learned representations, we can perform numerous downstream analyses more effectively. Among several approaches, deep learning has demonstrated remarkable performance in extracting hidden patterns and features from time-series data without manual feature engineering. This survey first presents a novel taxonomy based on three fundamental elements in designing state-of-the-art universal representation learning methods for time series. According to the proposed taxonomy, we comprehensively review existing studies and discuss their intuitions and insights into how these methods enhance the quality of learned representations. Finally, as a guideline for future studies, we summarize commonly used experimental setups and datasets and discuss several promising research directions. An up-to-date corresponding resource is available at https://github.com/itouchz/awesome-deep-time-series-representations.
time series, representation learning, neural networks, temporal modeling
ccs: Computing methodologies Neural networksccs: Computing methodologies Learning latent representationsccs: Mathematics of computing Time series analysisccs: Information systems Data mining
1. Introduction
1.1. Background
A time series is a sequence of data points recorded in chronological order, reflecting the complex dynamics of particular variables or phenomena over time. Time-series data can represent various meaningful information across application domains at different time points, enabling informed decision-making and predictions, such as sensor readings in the Internet of Things (Fathy et al., 2018; Cook et al., 2019), measurements in cyber-physical systems (Giraldo et al., 2018; Luo et al., 2021), fluctuation in stock markets (Ge et al., 2022; Cao, 2022), and human activity on wearable devices (Chen et al., 2021a; Gu et al., 2021). However, to extract and understand meaningful information from such complicated observations, we need a mechanism to represent these time series, which leads to the emergence of time-series representation research. Based on the new representations, we can effectively perform various downstream time-series analyses (Esling and Agon, 2012), e.g., forecasting (Lim and Zohren, 2021), classification (Ismail Fawaz et al., 2019), regression (Tan et al., 2021a), and anomaly detection (Choi et al., 2021). Fig. 1 depicts the basic concept of representation methods for time-series data.
Figure 1. Basic concept of time-series representation methods.
Early attempts (Sharma et al., 2020) represent time series using piecewise linear methods (e.g., piecewise aggregate approximation), symbolic-based methods (e.g., symbolic aggregate approximation), feature-based methods (e.g., shapelets), or transformation-based methods (e.g., discrete wavelet transform). These traditional time-series representation methods are known to be time-consuming and less effective because of their dependence on domain knowledge and the poor generality of predefined priors. Since the quality of representations significantly affects the downstream task performance, many studies propose to learn the meaningful time-series representations automatically (Zerveas et al., 2021; Yue et al., 2022; Yang and Hong, 2022). The main goal of these studies is to obtain high-quality learned representations of time series that capture valuable information within the data and unveil the underlying dynamics of the corresponding systems or phenomena. Among several approaches, neural networks or deep learning (DL) have demonstrated unprecedented performance in extracting hidden patterns and features from a wide range of data, including time series, without the requirement of manual feature engineering.
Given the sequential nature of time series, recurrent neural networks (RNN) and their variants, such as long short-term memory and gated recurrent unit, are considered a popular choice for capturing temporal dependencies in time series (Lalapura et al., 2021; Mao and Sejdić, 2022). Nevertheless, recurrent-based networks are complex and computationally expensive. Another line of work adopts one-dimensional convolutional neural networks (CNN) to improve the computational efficiency with the parallel processing of convolutional operations (Iwana and Uchida, 2021). Even though RNN- and CNN-based models are shown to be good at capturing temporal dependencies, they cannot explicitly model the relationship between different variables within a multivariate time series. Many studies propose to use attention-based networks or graph neural networks to jointly learn the temporal dependencies in each variable and the correlations between different variables in multivariate time series using attention mechanisms or graph structures (Wen et al., 2023; Jin et al., 2023a). Despite the significant progress in architectural design, time series can be collected irregularly or have missing values caused by sensor malfunctions in real-world scenarios, making the commonly used neural networks inefficient due to the adversarial side-effect during the imputation process. Consequently, recent research integrates neural ordinary differential equations into existing networks such that the models can produce continuous hidden states, thereby being robust to irregular time series (Rubanova et al., 2019a; Sun et al., 2020).
In addition, the reliability and efficacy of DL-based methods are generally contingent upon the availability of sufficiently well-annotated data, commonly known as supervised learning. Time-series data, however, is naturally continuous-valued, contains high levels of noise, and has less intuitively discernible visual patterns. In contrast to human-recognizable patterns in images or texts, time-series data can have inconsistent semantic meanings in real-world settings across application domains. As a result, obtaining a well-annotated time series is inevitably inefficient and considerably more challenging even for domain experts due to the convoluted dynamics of the time-evolving observations collected from diverse sensors or wearable devices with different frequencies. For example, we can collect a large set of sensor signals from a smart factory, while only a few of them can be annotated by domain experts. To circumvent the laborious annotation process and reduce the reliance on labeled instances, there has been a growing interest in unsupervised and self-supervised learning approaches using self-generated labels from various pretext tasks without relying on human annotation (Eldele et al., 2023; Zhang et al., 2023a; Ma et al., 2023; Meng et al., 2023).
While unsupervised and self-supervised representation learning share the same objective to extract latent representations from intricate raw time series without the human-annotated labels, their underlying mechanism is different. Unsupervised learning methods (Meng et al., 2023) usually adopt autoencoders and sequence-to-sequence models to learn meaningful representations using reconstruction-based learning objectives. However, accurately reconstructing the complex time-series data remains challenging, especially with high-frequency signals. On the contrary, self-supervised learning methods (Zhang et al., 2023a) leverage pretext tasks to autonomously generate labels by utilizing intrinsic information derived from the unlabeled data. Lately, pretext tasks with contrasting loss (also known as contrastive learning) have been proposed to enhance learning efficiency through discriminative pre-training with self-generated supervised signals. Contrastive learning aims to bring similar samples closer while pushing dissimilar samples apart in the feature space. These pretext tasks are self-generated challenges the model learns to solve from the unlabeled data, thereby being able to produce meaningful representations for multiple downstream tasks (Ericsson et al., 2022).
To further enhance the representation quality and alleviate the impact of limited training samples in particular settings where collecting sufficiently large data is prohibited (e.g., human-related services), several studies also employ data-related techniques, e.g., augmentation (Aboussalah et al., 2023) and transformation (Wu et al., 2023), on top of the existing learning methods. Accordingly, we can effectively increase the size and improve the quality of the training data. These techniques are also deemed essential in generating pretext tasks. Different from other data types, working with time-series data needs to consider their unique properties, such as temporal dependencies and multi-scale relationships (Wen et al., 2021).
Figure 2. Key design elements and downstream evaluation protocols.
1.2. Related Surveys
According to the background discussed above, there are three fundamental elements (also illustrated in Fig. 2) in designing a state-of-the-art universal representation learning method for time series: neural architectures, learning objectives, and training data. The neural architectures are designed to capture underlying temporal dependencies in time series and inter-relationships between variables of multivariate time series. Also, one or multiple learning objectives (i.e., loss functions) are newly defined to learn high-quality representations. These learning objectives are sometimes called pretext tasks if pseudo labels are generated. Last, to enhance the utility and quality of available training data, data-related techniques (e.g., augmentation) are employed or introduced.
Despite having the three key design elements, most existing surveys for time-series representation learning review the literature exclusively on either the neural architectural aspects (Längkvist et al., 2014) or the learning aspects (Zhang et al., 2023a; Ma et al., 2023; Meng et al., 2023). An early related survey article by Längkvist et al. (2014) reviews the deep learning for time-series data as unsupervised feature learning algorithms. The survey focuses particularly on neural architectures with a discussion on classical time-series applications. After about a decade, a few survey papers review time-series representation learning methods by focusing on learning objective aspects. For example, Zhang et al. (2023a) and Deldari et al. (2022) review self-supervised learning-based models, while, for a broader scope, Meng et al. (2023) review unsupervised learning-based methods. Similarly, Ma et al. (2023) present the survey for any learning objectives with a focus on analyzing reviewed articles from transfer learning and pre-training perspectives. For a smaller scope, a survey (Eldele et al., 2023) reviews the state-of-the-art studies that specifically tackle label scarcity in time-series data. Unlike existing survey papers, this paper comprehensively reviews the representation learning methods for time series by focusing on their universality with discussions on their intuitions and insights into how these methods enhance the quality of learned representations from all three design aspects. Specifically, we aim to review and identify research directions on how recent state-of-the-art studies design the neural architecture, devise corresponding learning objectives, and utilize training data to enhance the quality of the learned representations of time series for downstream tasks. Table 1 summarizes the differences between our survey and the related work.
Table 1. Comparison of the survey scope between this article and related papers.
Survey Article Novel Taxonomy Key Design Elements in Review Survey Coverage Future Research Direction
Neural
Architecture
Learning
Objective
Training
Data
Universality Irregularity
Längkvist et al. (2014) $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$
Deldari et al. (2022) $\checkmark$ $\times$ Only Self-Supervised Learning $\times$ $\checkmark$ $\times$ $\checkmark$
Eldele et al. (2023) $\checkmark$ $\times$ $\checkmark$ Only Augmentation $\checkmark$ $\times$ $\checkmark$
Ma et al. (2023) $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$
Zhang et al. (2023a) $\checkmark$ $\times$ Only Self-Supervised Learning $\times$ $\times$ $\times$ $\checkmark$
Meng et al. (2023) $\checkmark$ $\times$ Only Unsupervised Learning $\times$ $\checkmark$ $\times$ $\checkmark$
Our Survey $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
1.3. Survey Scope and Literature Collection
For the literature review, we search for papers using the following keywords and inclusion criteria.
1.3.1. Keywords
“time series” AND “representation”, “time series” AND “embedding”, “time series” AND “encoding”, “time series” AND “modeling”, “time series” AND “deep learning”, “temporal” AND “representation”, “sequential” AND “representation”, “audio” AND “representation”, “sequence” AND “representation”, and (“video” OR “action”) AND “representation”. We use these keywords to search well-known repositories, including ACM Digital Library, IEEE Xplore, Google Scholars, Semantic Scholars, and DBLP, for the relevant papers.
1.3.2. Inclusion Criteria
The initial set of papers found with the above search queries are further filtered by the following criteria. Only papers meeting the criteria are included for review.
  • Being written in the English language only
  • Being a deep learning or neural networks-based approach
  • Being published in or after 2018 at a top-tier conference or high-impact journal [^1] Top-tier venues are evaluated based on CORE (https://portal.core.edu.au), KIISE (https://www.kiise.or.kr), or Google Scholar (https://scholar.google.com). Only publications from venues rated at least A by CORE/KIISE or in the top 20 in at least one subcategory by Google Scholar metrics are included for review. Recent arXiv papers also included if their authors have publication records in the qualified venues.
  • Being evaluated on at least two downstream tasks using time-series, video, or audio datasets
(a) Focus Design Element.
(b) Publication Venue.
(c) Publication Year.
Figure 3. Quantitative summary of the selected papers.
1.3.3. Quantitative Summary of Selected Papers
Given the above keywords and inclusion criteria, we selected 105 papers in total. Fig. 3 shows the quantitative summary of the paper selected for review. We can notice from Fig. 2(a) that neural architectures and learning objectives are similarly considered important in designing state-of-the-art methods. Most papers were published at NeurIPS, followed by AAAI, CVPR, and ICLR (Fig. 2(b)). According to Fig. 2(c), we expect more papers on this topic will be published in the future.
1.4. Contributions
This paper aims to identify what are the essential elements in designing state-of-the-art representation learning methods for time series and how these elements affect the quality of the learned representations. To the best of our knowledge, this is the first survey on universal time-series representation learning. We propose a novel taxonomy for learning universal representations of time series from the novelty perspectives—whether the main contribution of a paper focuses on what part of the design elements discussed above—to summarize the selected studies. Table 2 summarizes and compares the reviewed articles based on the proposed taxonomy. From the perspective of neural architectures, we summarize the changes made on both the module level and network level of neural architectures used for representation learning. From the learning perspective, we classify what objectives are used to make the learned representations generalizable to various downstream tasks. Last, we also categorize data-centric methods for the papers that focus particularly on improving the quality of the training data. Overall, our main contributions are as follows.
  • We conduct an extensive literature review of universal time-series representation learning based on a novel and up-to-date taxonomy that categorizes the reviewed methods into three main categories: neural architectural, learning-focused, and data-centric approaches.
  • We provide a guideline on the experimental setup and benchmark datasets for assessing representation learning methods for time series.
  • We discuss several open research challenges and new insights to facilitate future work.
1.5. Article Organization
Section 2 introduces the definitions and specific background knowledge regarding time-series representation learning. In Section 3, we present a comprehensive review of the methods focusing on neural architecture aspects. We then discuss the methods focusing on deriving new learning objectives in Section 4. Section 5 gives a review on data-centric approaches. In addition, we discuss the evaluation protocol for time-series representation learning and promising future research directions in Section 6 and Section 7, respectively. Finally, Section 8 concludes this survey.
Table 2. Taxonomy of universal representation learning for time series.
Design Element Category Section Technique Key References
Neural Architectures (Neural Architectural Approaches) Basic Block Combination 3.1 Network-Level Combination (Chen et al., 2021b; Liu et al., 2020; Zerveas et al., 2021; Jhin et al., 2021; Abushaqra et al., 2022; Jhin et al., 2022; Kim et al., 2023a; Rubanova et al., 2019b; Sanchez et al., 2019; Xie et al., 2022; Han et al., 2020; Choi and Kang, 2023; Rahman et al., 2021; Franceschi et al., 2019; Wang et al., 2018; Yang and Hong, 2022; Li et al., 2023; Liang et al., 2023a; Fraikin et al., 2023)
Module-Level Combination (Anand and Nayak, 2021; Nguyen et al., 2023; Wu et al., 2023; Chen et al., 2019; Zhou et al., 2023; Luetto et al., 2023; Guo et al., 2021)
Innovative Redesign 3.2 Network-Level Design (Liang et al., 2023b; Zhang et al., 2023b; Bianchi et al., 2019; Schirmer et al., 2022; Biloš et al., 2023; Ansari et al., 2023; Wu et al., 2018; Sun et al., 2021; Liu et al., 2019; Fons et al., 2022; Wang et al., 2023a; Li et al., 2022; Cai et al., 2021; Lee et al., 2022a; Wang et al., 2023b; Chen et al., 2023; Zhang et al., 2023c; Zhong et al., 2023; Xu et al., 2023)
Module-Level Design (Ma et al., 2019; Wang et al., 2023c; Gorbett et al., 2023; Shukla and Marlin, 2021; Sener et al., 2020; Chowdhury et al., 2022; Liu et al., 2022; Naour et al., 2023; Lin et al., 2023)
Learning Objectives (Learning-Focused Approaches) Supervised Learning 4.1 Task-Specific Loss (Ma et al., 2019; Rahman et al., 2021; Hadji et al., 2021; Zhong et al., 2023)
Unsupervised Learning 4.2 Masked Prediction (Zerveas et al., 2021; Luetto et al., 2023; Chowdhury et al., 2022)
Input Reconstruction (Sanchez et al., 2019; Yuan et al., 2019; Tonekaboni et al., 2022; Nguyen et al., 2018; Dong et al., 2023; Li et al., 2023; Chen et al., 2019)
Self-Supervised Learning 4.3 Non-Contrasting Loss (Guo et al., 2022; Haresh et al., 2021; Liang et al., 2022; Wang et al., 2020, 2021; Duan et al., 2022; Fang et al., 2023; Fraikin et al., 2023)
Contrasting Loss (Liang et al., 2023b; Zhang et al., 2022a; Kong et al., 2020; Tonekaboni et al., 2022; Ding et al., 2022; Kim et al., 2023a; Chen et al., 2022; Zhang and Crandall, 2022; Haresh et al., 2021; Qing et al., 2022; Morgado et al., 2020; Nguyen et al., 2023; Han et al., 2020; Zhang et al., 2023d; Choi and Kang, 2023; Qian et al., 2022; Chowdhury et al., 2023; Chen et al., 2021c; Zhang et al., 2022b; Hajimoradlou et al., 2022; Qian et al., 2021; Liu et al., 2022; Dave et al., 2022; Yang et al., 2023a; Jenni and Jin, 2021; Jiao et al., 2020; Yang et al., 2022a; Somaiya et al., 2022; Yue et al., 2022; Tonekaboni et al., 2021; Franceschi et al., 2019; Lee et al., 2022a; Shin et al., 2023; Wang et al., 2023b; Liu et al., 2023; Zhang et al., 2023e; Sun et al., 2023; Eldele et al., 2021)
Training Data (Data-Centric Approaches) Data Augmentation 5.1 Random Augmentation (Yue et al., 2022; Zhang et al., 2022b, 2023e)
Policy-based Augmentation (Aboussalah et al., 2023; Kim et al., 2023b; Luo et al., 2023; Kim et al., 2023c; Zhang and Crandall, 2022; Yang and Hong, 2022; Chen et al., 2022; Shin et al., 2023; Yang et al., 2022a; Demirel and Holz, 2023)
Decomposition and Transformation 5.2 Time-Series Decomposition (Zeng et al., 2021; Behrmann et al., 2021; Wang et al., 2018; Yang et al., 2022b; Fang et al., 2023)
Input Space Transformation (Anand and Nayak, 2021; Biloš et al., 2022; Lee et al., 2022b; Wu et al., 2023; Chanpuriya et al., 2023; Zhong et al., 2023; Xu et al., 2023)
Methods for Sample Selection 5.3 Generative Methods (Nguyen et al., 2023; Zhao et al., 2023)
(Re-)Sampling Strategies (Chen et al., 2021b; Franceschi et al., 2019)
2. Preliminaries
This section presents definitions and notations used throughout this paper, descriptions of downstream tasks, unique properties in time series, and basic building blocks of neural architectures.
2.1. Definitions
Definition 2.1 (Time Series).
A time series $\mathbf{X}$ is an chronologically ordered sequence of $V$-variate data points recorded at specific time intervals. $\mathbf{X}=(\mathbf{x}_{1},\dots,\mathbf{x}_{t},\mathbf{x}_{T})$, where $\mathbf{x}_{t}\in\mathbb{R}^{V}$ is the observed value at $t$-th time step, $V$ is the number of variables, and $T$ is the length of the time series. When $V=1$, it becomes the univariate time series; otherwise, it is multivariate time series. Audio and video data can be considered special cases of time series with more dimensions. The time intervals are typically equally spaced, and the values can represent any measurable quantity, such as temperature, sales figures, or any phenomenon that changes over time.
Definition 2.2 (Irregularly-Sampled Time Series).
An irregularly-sampled time series is a time series that the intervals between observations are not consistent or regularly spaced. Thus, the time intervals between $(\mathbf{x}_{1},\mathbf{x}_{2})$ and $(\mathbf{x}_{2},\mathbf{x}_{3})$ is unequal, as illustrated in Fig. 4. It is often encountered in situations where data is collected opportunistically or when events occur irregularly and sporadically, e.g., sensor malfunctions, leading to varying time gaps between observations.
Definition 2.3 (Time-Series Representation Learning).
Given a raw time series $\mathbf{X}$, time-series representation learning aims to learn an encoder $f_{e}$, a nonlinear embedding function that maps $\mathbf{X}$ into a $R$-dimensional representation vector $\mathbf{Z}=(\mathbf{z}_{1},\dots,\mathbf{z}_{R-1},\mathbf{z}_{R})$ in the latent space, where $\mathbf{z}_{i}\in\mathbb{R}^{F}$. $\mathbf{Z}$ usually has either equal ($R=T$) or shorter ($R<T$) length of the original time series. When $R=T$, $\mathbf{Z}$ is timestamp-wise (or point-wise) representation that contains representation vectors $\mathbf{z}_{i}$ with feature size $F$ for each $t$. In contrast, when $R<T$, $\mathbf{Z}$ is the compressed version of $\mathbf{X}$ with reduced dimension, and $F$ is usually $1$, producing the series-wise (or instance-wise) representation.
A crucial measure to assess the quality of a representation learning method, i.e., the encoder $f_{e}$, is its ability to produce the representations $\mathbf{Z}$ that effectively facilitate downstream tasks, either with or without fine-tuning (see Section 6). Once we obtain the latent representation $\mathbf{Z}$, we can use it as input for downstream tasks to evaluate its actual performance. Here, we define the most common downstream tasks as follows.
Definition 2.4 (Forecasting).
Time-series forecasting (TSF) aims to predict the future values of a time series by explicitly modeling the dynamics and dependencies among historical observations. It can be short-term or long-term forecasting depending on the prediction horizon $H$. Formally, given a time series $\mathbf{X}$, TSF predicts the next $H$ values $(\mathbf{x}_{T+1},\dots,\mathbf{x}_{T+H})$ that are most likely to occur.
Definition 2.5 (Classification).
Time-series classification (TSC) aims to assign predefined class labels $\mathbf{C}=\{c_{1},\dots,c_{|\mathbf{C}|}\}$ to a set of time series. Let $\mathcal{D}=\{(\mathbf{X}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$ denote a time-series dataset with $N$ samples, where $\mathbf{X}_{i}\in\mathbb{R}^{T\times V}$ is a time series and $\mathbf{y}_{i}$ is the corresponding one-hot label vector of length $|\mathbf{C}|$. The $j$-th element of the one-hot vector $\mathbf{y}_{i}$ is equal to 1 if the class of $\mathbf{X}_{i}$ is $j$, otherwise 00. Formally, TSC trains a classifier on the given dataset $\mathcal{D}$ by learning the discriminative features to distinguish different classes from each other. Then, when an unseen dataset $\mathcal{D}^{\prime}$ is input to the trained classifier, it automatically determines to which class $c_{i}$ each time series belongs.
Definition 2.6 (Extrinsic Regression).
Time-series extrinsic regression (TSER) shares a similar goal to TSC with a key difference in label annotation. While TSC predicts a categorical value, TSER predicts a continuous value for a variable external to the input time series. That is, $y_{i}\in\mathbb{R}$. Formally, TSER trains a regression model to map a given time series $\mathbf{X}_{i}$ to a numerical value $y_{i}$.
Definition 2.7 (Clustering).
Time-series clustering (TSCL) is the process of finding natural groups, called clusters, in a set of time series $\mathcal{X}=\{\mathbf{X}_{i}\}_{i=1}^{N}$. It aims to partition $\mathcal{X}$ into a group of clusters $\mathbf{G}=\{g_{1},\dots,g_{i},g_{|\mathbf{G}|}\}$ by maximizing the similarities between time series within the same cluster and the dissimilarities between time series of different clusters. Formally, given a similarity measure $f_{s}(\cdot,\cdot)$, $\forall i_{1},i_{2},j~{}f_{s}(\mathbf{X}_{i_{1}},\mathbf{X}_{j})\gg f_{s}(\mathbf{X}_{i_{1}},\mathbf{X}_{i_{2}})~{}\text{for}~{}\mathbf{X}_{i_{1}},\mathbf{X}_{i_{2}}\in g_{i}$ and $\mathbf{X}_{j}\in g_{j}$.
(a) Regularly Sampled Time Series.
(b) Irregularly Sampled Time Series.
Figure 4. Illustrations of regularly and irregularly sampled multivariate time series ($V=3$).
Definition 2.8 (Segmentation).
Time-series segmentation (TSS) aims to assign a label to a subsequence $\mathbf{X}_{T_{s},T_{e}}$ of $\mathbf{X}$, where $T_{s}$ is the start offset and $T_{e}$ is the end offset, consisting of contiguous observations of $\mathbf{X}$ from time step $T_{s}$ to $T_{e}$. That is, $\mathbf{X}_{T_{s},T_{e}}=(\mathbf{x}_{T_{s}},\dots,\mathbf{x}_{T_{e}})$ and $1\leq T_{s}\leq T_{e}\leq T$. Let a change point (CP) be an offset $i\in[i,\dots,T]$ w.r.t. to a state transition in time series, TSS finds a set of segmentation of $\mathbf{X}$, having the ordered sequence of CPs in $\mathbf{X}$ (i.e., $\mathbf{x}_{i_{1}},\dots,\mathbf{x}_{i_{S}}$) with $1<i_{1}<\cdots<i_{S}<T$ at which the state of observations changed. After identifying the number and locations of all CPs, we can set the start offset $T_{e}$ and end offset $T_{e}$ for each segment in $\mathbf{X}$.
Definition 2.9 (Anomaly Detection).
Time-series anomaly detection (TSAD) aims to identify abnormal time points that significantly deviate from the other observations in a time series. Commonly, TSAD learns the representations of normal behavior from a time series $\mathbf{X}$. Then, the trained model computes anomaly scores $\mathbf{A}=(a_{i},\dots,a_{|\mathbf{X}^{\prime}|})$ of all values in an unseen time series $\mathbf{X}^{\prime}$ to determine which time point in $\mathbf{X}^{\prime}$ is anomalous. The final decisions are obtained by comparing each $a_{i}$ with a predefined threshold $\delta$: anomalous if $a_{i}>\delta$ and normal otherwise.
Definition 2.10 (Imputation of Missing Values).
Time-series imputation (TSI) aims to fill missing values in a time series with realistic values to facilitate subsequent analysis. Given a time series $\mathbf{X}$ and a binary $M=(m_{1},\dots,m_{t},m_{T})$, $\mathbf{x}_{t}$ is missing if $m_{t}=0$, and is observed otherwise. Let $\mathbf{\hat{X}}$ denote the predicted values generated by a TSI method, the imputed time series $\mathbf{X}_{\text{imputed}}=\mathbf{X}\odot M+\mathbf{\hat{X}}\odot(1-M)$.
Definition 2.11 (Retrieval).
Time-series retrieval (TSR) aims to obtain a set of time series that are most similar to a query provided. Given a query time series $\mathbf{X}_{q}$ and a similarity measure $f_{s}(\cdot,\cdot)$, find an ordered list $\mathcal{Q}=\{\mathbf{X}_{i}\}_{i=1}^{K}$ of time series in the given dataset or database, containing $K$ time series that are the most similar to $\mathbf{X}_{q}$.
Note that the above definitions are generally based on the raw time series $\mathbf{X}$. Following Definition 2.3, we can use the corresponding representation $\mathbf{Z}=f_{e}(\mathbf{X})$ to perform the above downstream tasks instead of the raw data.
2.2. Unique Properties of Time Series
In this subsection, we discuss unique properties in time-series data that have been explored by existing studies for time-series representation learning (Längkvist et al., 2014; Ma et al., 2023). Due to the following properties, techniques devised for image or text data are usually difficult to transfer directly to time series.
2.2.1. Temporal Dependency
Time series exhibits dependencies on the time variable, where a data point at a given time correlates with its previous values. Given an input $\mathbf{x}_{t}$ at time $t$, the model predicts $y_{t}$, but the same input at a later time could be a different prediction. Therefore, windows or subsequences of past observations are usually included as inputs to the model to learn such temporal dependency. The length of windows for capturing the time dependencies could also be unknown. In addition, there are local and global temporal dependencies. The former is usually associated with abrupt changes or noises, while the latter is associated with collective trends or recurrent patterns.
2.2.2. High Noise and Dimension
Time-series data, especially in real-world environments, often contain noises and have high dimensions. These noises usually arise from measurement errors or other sources of uncertainty. Dimensionality reduction techniques and wavelet transforms can address this issue by filtering some noises and reducing the dimension of the raw time series. Nevertheless, we may lose valuable information and need domain-specific knowledge to select the suitable dimensionality reduction and filtering techniques.
2.2.3. Inter-relationship across Variables
This characteristic is particularly notable in multivariate time series. It is often uncertain whether there is sufficient information to understand a phenomenon of a time series when only a limited number of variables are analyzed as there may exist relationships between variables underlying the process or state. For example, in electronic nose data, where an array of sensors with various selectivity for several gases are combined to determine a specific smell, there is no guarantee that the selection of sensors can identify the target odor. Similarly, in financial data, monitoring a single stock representing a fraction of a complex system may not provide enough information to forecast future values.
2.2.4. Variability and Nonstationarity
Time-series data also possess variability and nonstationarity properties, meaning that statistical characteristics, such as mean, variance, and frequency, change over time. These changes usually reveal seasonal patterns, trends, and fluctuations. Here, seasonality refers to repeating patterns that regularly appear, while trends describe long-term changes or shifts over time. In some cases, the change in frequency is so relevant to the task that it is more beneficial to work in the frequency domain than in the time domain.
2.2.5. Diverse Semantics
In contrast to image and text data, learning universal representations of time series is challenging due to the lack of a large-scale unified semantic time-series dataset. For instance, each word in a text dataset has similar semantics in different sentences with high probability. Accordingly, the word embeddings learned by the model can transfer across different scenarios. However, time-series datasets are challenging to obtain subsequences (corresponding to words in text sequences) that have consistent semantics across scenarios and applications, making it difficult to transfer the knowledge learned by the model. This property also makes time-series annotation tricky and challenging, even for domain experts.
2.3. Neural Architectures for Time Series
Choosing an appropriate neural architecture to model complex temporal and inter-relationship dependencies in time series is an essential part of the design elements. This subsection overviews basic neural network architectures used as building blocks in state-of-the-art representation learning methods for time series.
2.3.1. Multi-Layer Perceptrons (MLP)
The most basic neural network architecture is the MLP (Ismail Fawaz et al., 2019), i.e., fully connected (FC) network. In MLP models, the number of layers and neurons (or hidden units) are hyperparameters to be defined. Specifically, all neurons in a layer are connected to all neurons of the following layer. These connections contain weights in the neural network. The weights are later updated by applying a non-linearity to an input. As MLP-based models process input data in a single fixed-length representation without considering the temporal relationships between the data points, they are basically unsuitable for capturing the temporal dependencies and time-invariant features. Each time step is weighted individually, and time-series elements are learned independently from each other.
2.3.2. Recurrent Neural Networks (RNN)
RNN (Mao and Sejdić, 2022) is a neural architecture with internal memory (i.e., state) specifically designed to process sequential data, thus suitable for learning temporal features in time series. The memory component enables RNN-based models to refer to past observations when processing the current one, resulting in improved learning capability. RNNs can also process variable-length inputs and produce variable-length outputs. This capability is enabled by sharing parameters over time through direct connections between layers. However, they are ineffective in modeling long-term dependencies, besides being computationally expensive due to sequential processing. RNN-based models are usually trained iteratively using a technique called back-propagation through time. Unfolded RNNs are similar to deep networks with shared parameters connected to each RNN cell. Due to the depth and weight-sharing in RNNs, the gradients are summed up at each time step to train the model, undergoing continuous matrix multiplication due to the chain rule. Thus, the gradients often either shrink exponentially to small values (i.e., vanishing gradients) or blow up to large values (i.e., exploding gradients). These problems give rise to long short-term memory (LSTM) and gated recurrent unit (GRU).
Long Short-Term Memory (LSTM)
LSTM (Hochreiter and Schmidhuber, 1997) addresses the well-known vanishing and exploding gradient problems in the standard RNNs by integrating memory cells with a gating mechanism (i.e., cell state gate, input gate, output gate, and forget gate) into their state dynamics to control the information flow between cells. As the inherent nature of LSTM is the same as RNN, LSTM-based models are also suitable for learning sequence data like time series and video representation learning, with a better capability to capture long-term dependencies.
Gated Recurrent Unit (GRU)
GRU (Cho et al., 2014) is another popular RNN variant that can control information flow and memorize states across multiple time steps, similar to LSTM, but with a simpler cell architecture. Compared to LSTM, GRU cells have only two gates (reset and update gates), making it more computationally efficient and requiring less data to generalize.
2.3.3. Convolutional Neural Networks (CNN)
CNN (Ismail Fawaz et al., 2019) is a very successful neural architecture, proven in many domains, including computer vision, speech recognition, and natural language processing. Accordingly, researchers also adopt CNN for time series. To use CNN for time series, we need to encode the input data into an image-like format. The CNN receives embedding of the value at each time step and then aggregates local information from nearby time steps using convolution layers. The convolution layer, consisting of several convolution kernels (i.e., filters), aims to learn feature representations of the inputs by computing different feature maps. Each neuron of a feature map connects to a region of neighboring neurons in the previous layer called the receptive field. The feature maps can be created by convolving the inputs with learned kernels and applying an element-wise nonlinear activation to the convolved results. Here, all spatial locations of the input share the kernel for each feature map, and several kernels are used to obtain the entire feature map. Many improvements have been made to CNN, such as using deeper networks, applying smaller and more efficient convolutional filters, adding pooling layers to reduce the resolution of the feature maps, and utilizing batch normalization to improve the training stability. As standard CNNs are designed for processing images, widely used CNN architectures for time series are one-dimensional CNN and temporal convolutional networks (Eldele et al., 2023).
Temporal Convolutional Networks (TCN)
Different from the standard CNNs, TCN (Bai et al., 2018) uses the fully convolutional network to make all layers the same length and employ casual convolution operation to avoid information leakage from the future time step to the past. Compared to RNN-based models, TCN has recently shown to be more accurate, simpler, and more efficient across diverse downstream tasks (Ma et al., 2023).
2.3.4. Graph Neural Networks (GNN)
GNN (Jin et al., 2023a) aims to learn directly from graph representations of data. A graph consists of nodes and edges, with each edge connecting two nodes. Both nodes and edges can have associated features. Edges can be directional or un-directional and can be weighted. Graphs better represent data not easily represented in Euclidean space, such as spatio-temporal data like the electroencephalogram and traffic monitoring networks. GNNs receive the graph structure and any associated node and edge attributes as input. Typically, the core operation in GNNs is graph convolution, which involves exchanging information across neighboring nodes. This operation enables the GNN-based model to explicitly rely on the inter-variable dependencies represented by the graph edges for processing multivariate time-series data. While both RNNs and CNNs perform well on Euclidean data, time series are often more naturally represented as graphs in many scenarios. Consider a network of traffic sensors where the sensors are not uniformly spaced, deviating from a regular grid. Representing this data as a graph captures its irregularity more precisely than a Euclidean space. However, using standard deep learning algorithms to learn from graph structures is challenging as nodes may have varying numbers of neighboring nodes, making it difficult to apply the convolution operation. Thus, GNNs are more suitable to graph data.
2.3.5. Attention-based Networks
The attention mechanism was introduced by Bahdanau et al. (2015) to improve the performance of encoder-decoder models in machine translation. The encoder encodes a source sentence into a vector in latent space, and the decoder then decodes the latent vector into a target language sentence. The attention mechanism enables the decoder to pay attention to the segments of the source for each target through a context vector. Attention-based neural networks are designed to capture long-range dependencies with broader receptive fields, usually lacking in CNNs and RNNs. Thus, attention-based models provide more contextual information to enhance the models’ learning and representation capability. The underlying mechanism (i.e., attention mechanism) is proposed to make the model focus on essential features in the input while suppressing the unnecessary ones. For instance, it can be used to enhance LSTM performance in many applications by assigning attention scores for LSTM hidden states to determine the importance of each state in the final prediction. Moreover, the attention mechanism can improve the interpretability of the model. However, it can be more computationally expensive due to the large number of parameters, making it prone to overfitting when the training data is limited.
Self-Attention Module
Self-attention has been demonstrated to be effective in various natural language processing tasks due to its ability to capture long-term dependencies in text (Foumani et al., 2023). The self-attention module is usually embedded in encoder-decoder models to improve the model performance and leveraged in many studies to replace the RNN-based models to improve learning efficiency due to its fast parallel processing.
Transformers
The unprecedented performance of stacking multi-headed attention, called Transformers (Vaswani et al., 2017), has led to numerous endeavors to adapt multi-headed attention to time-series data. Transformers for time series usually contain a simple encoder structure consisting of multi-headed self-attention and feed-forward layers. They integrate information from data points in the time series by dynamically computing the associations between representations with self-attention. Thanks to the practical capability to model long-range dependencies, Transformers have shown remarkable performance in sequence data. Many recent studies use Transformers as the backbone architecture for time-series analysis. (Wen et al., 2023).
2.3.6. Neural Ordinary Differential Equations (Neural ODE)
Let $f_{\theta}$ be a function specifying continuous dynamics of a hidden state $\mathbf{h}(t)$ with paramters $\theta$. Neural ODEs are one of the continuous-time models that define the hidden state $\mathbf{h}(t)$ as a solution to ODE initial-value problem $\frac{d\mathbf{h}(t)}{dt}=f_{\theta}(\mathbf{h}(t),t)$ where $\mathbf{h}(t_{o})=\mathbf{h}_{0}$. The hidden state $\mathbf{h}(t)$ is defined at all time steps, and can be evaluated at any desired time steps using a numerical ODE solver. Formally, $\mathbf{h}_{0},\dots,\mathbf{h}_{T}=\text{ODESolver}(f_{\theta},\mathbf{h}_{0},(t_{0},\dots,t_{T}))$. For training ODE-based deep learning models using black-box ODE solvers, we can use the adjoint sensitivity method to compute memory-efficient gradients w.r.t. the neural network parameters $\theta$, as described by Rubanova et al. (2019a). Neural ODEs are usually combined with RNN or its variants to sequentially update the hidden state at observation times (Shukla and Marlin, 2020). These models provide an alternative recurrence-based solution with better properties than traditional RNNs in terms of their ability to handle irregularly sampled time series.
3. Neural Architectural Approaches
As neural architectures affect representation quality (Trirat et al., 2020), studies in this section focus on the novel design of neural architectures by combining basic building blocks or redesigning a neural architecture from scratch to improve the capability of capturing temporal dependencies and inter-relationships between variables of multivariate time series. We can further categorize the studies into the basic block combination and innovative redesign categories based on the degree of adjustment.
3.1. Basic Block Combination
Studies in this group are based on a combination of basic building blocks without or with trivial additional modifications on particular elements. The combination can be done at the network level or module level of neural architectures.
3.1.1. Network-Level Combination
Wang et al. (2018) propose a wavelet-based neural architecture, called mWDN, by integrating multi-level discrete wavelet decomposition into existing neural networks for building frequency-aware deep models. It enables the fine-tuning of all parameters within the deep learning framework, while preserving the benefits of multi-layer discrete wavelet decomposition in frequency learning. Based on mWDN, two task-specific neural networks are introduced for classification (i.e., residual classification flow) and forecasting (i.e., multi-freqeuncy LSTM). To improve efficiency and scalability for learning long time series, Franceschi et al. (2019) propose a representation learning framework that adopts an encoder-only network based on dilated causal 1D-CNN in their neural architecture design. Sanchez et al. (2019) focus on leveraging a combination of variational autoencoder and generative adversarial networks to create disentangled representations of satellite image time series, which separate common information shared across the time series images from exclusive information unique to each image. Similar to mWDN, Chen et al. (2021b) propose multi-task representation learning method called MTRL by exploiting supervised learning for classification and unsupervised learning for retrieval. MTRL jointly optimizes the two downstream tasks via a combination of deep wavelet decomposition networks to extract multi-scale subseries and 1D-CNN residual networks to learn time-domain features, thus improving the performance of downstream tasks. Unlike earlier studies using CNNs, TST (Zerveas et al., 2021) is the first Transformer-based framework for representation learning of multivariate time series. TST combines a base Transformer network with an additional input encoder and a learnable positional encoding to make the multivariate time series work seamlessly with the Transformer encoder. Specifically, the paper argues that the multi-headed attention mechanism makes the Transformer models suitable for time-series data by concurrently representing each input sequence element using its future-past contexts, while multiple attention heads can consider different representation subspaces, i.e., multiple aspects of relevance between input elements.
More recently, MARINA (Xie et al., 2022) combines an MLP network with residual connections and a graph attention network (an attention-based variant of GNNs) to form a temporal module and a spatial module, respectively. The temporal module efficiently captures temporal correlations in the time series; at the same time, the spatial module learns the similarities among time series to enhance the downstream performance. To enhance the representation learning process even further, Yang and Hong (2022) propose an unsupervised representation learning framework for time series, named BTSF. It enhances the representation quality through the more reasonable construction of contrastive pairs and the adequate integration of temporal and spectral information. BTSF constitutes an iterative application of a novel bi-linear temporal-spectral fusion, explicitly encoding affinities between time and frequency pairs. To adequately use the informative affinities, BTSF further uses a cross-domain interaction with spectrum-to-time and time-to-spectrum aggregation modules to iteratively refine temporal and spectral features for cycle update, proven effective by empirical and theoretical analysis. Kim et al. (2023a) introduce another MLP-based feature-wise encoder together with a element-wise gating layer built on top of TS2Vec (Yue et al., 2022) i.e., feature-agnostic temporal representation using TCN. Choi and Kang (2023) also extend the TS2Vec encoder (Yue et al., 2022) with a multi-task self-supervised learning framework by combining contextual, temporal, and transformation consistencies into a single networks. Based on the existing TCN architecture, Fraikin et al. (2023) attach a time-embedding module to explicitly learn time-related features, such as trend, periodicity, and distribution shifts. Liang et al. (2023a) design a self-supervised pre-training universal framework, UniTS, that uses a pre-training module consisting of templates from various self-supervised learning methods. Subsequently, the pre-trained representations are fused, and the results of feature fusion are applied to task-specific output models.
Another line of studies use Neural ODEs to handle irregular time series. ODE-RNN (Rubanova et al., 2019b) is an early attempt applying Neural ODEs for time series. ODE-RNN serves as an encoder in the latent ODE model, facilitating interpolation, extrapolation, and classification tasks for irregular time series. In recent work, ANCDEs (Jhin et al., 2021) and EXIT (Jhin et al., 2022) propose neural controlled differential equation (Neural CDE)-based approaches for time-series classification and forecasting with irregular time series. ANCDEs leverages two Neural CDEs, one for learning attention from the input time series and another for creating representations for downstream tasks. Likewise, EXIT utilizes Neural CDEs as part of the encoder-decoder network, enabling interpolation and extrapolation of irregular time-series data. Additionally, CrossPyramid (Abushaqra et al., 2022) addresses the limitation of ODE-RNN which is the high dependence on the initial observations by using pyramid attention and cross-level ODE-RNN.
For video learning, T-C3D (Liu et al., 2020) incorporates a residual 3D-CNN that captures both appearance information from single frames and motion information between consecutive frames. This multi-granularity learning allows for a more comprehensive understanding of video actions. MemDPC (Han et al., 2020) is a framework of video representation using memory-augmented dense predictive coding. It is trained on a predictive attention mechanism over the set of compressed memories. Rahman et al. (2021) introduce a tri-modal VilBERT-inspired model by integrating separate encoders for vision modality, pose modality, and audio modality into a single network.
3.1.2. Module-Level Combination
Audio Word2vec (Chen et al., 2019) extends vector representations to consider the sequential phonetic structures of the audio segments trained with speaker-content disentanglement based segmental sequence-to-sequence autoencoder. Guo et al. (2021) design SSAN with a separable self-attention module to capture spatial and temporal correlations in videos separately. Spatial self-attention is applied independently to input frames. The spatial contextual information is then aggregated over the temporal dimension and sent to a temporal attention module.
Instead learning the representations from raw time series, DelTa (Anand and Nayak, 2021) uses 2D images of time series such that existing pre-trained models on large image datasets can be easily used. DelTa proposes two versions of using pre-trianed vision models: layout aligned version and layout independent version. UniTTab (Luetto et al., 2023) is a Transformer-based framework for time-dependent heterogeneous tabular data. UniTTab uses row-type dependent encoders and different feature representation methods for categorical and numerical data, respectively. Wu et al. (2023) introduce TimesNet, which reshapes the 1D time series into a 2D tensor using a newly proposed neural network module (called TimesBlock). TimesBlock combines a Fast Fourier Transform layer with a parameter-efficient Inception block. In particular, it initially extracts periods using the Fourier transform and reshapes the data into 2D. Subsequently, TimesNet utilizes the parameter-efficient Inception network to extract representations from the 2D image-like data, which are later restored to 1D time series. Instead of training the Transformer models from scratch, One Fits All (Zhou et al., 2023) uses a pre-trained language model (e.g., GPT-2) by freezing self-attention and feed-forward layers and fine-tuning the remaining layers. Input embedding and normalization layers are also modified for time-series data. By doing so, it can benefit from the universality of the Transformer models on time-series data. Nguyen et al. (2023) propose CoInception by integrating the dilated CNN into the Inception block to build a scalable and robust neural architecture with a wide receptive field. CoInception uses multi-scale filters to capture the temporal dependency at multiple scales and resolutions. Specifically, it incorporates a novel convolution-based aggregator and extra skip connections in the Inception block to enhance the capability for capturing long-term dependencies in the input time series.
Remark. The evolving landscape of neural architecture design for time-series data showcases a blend of creative combinations at both the network and module levels. Researchers have integrated various techniques, from wavelet decomposition to attention-based networks, towards enhancing efficiency, scalability, and performance. These endeavors expand the horizons of time-series representation and underscore the importance of adaptability in deep neural networks.
3.2. Innovative Redesign
Unlike methods in Section 3.1, methods in this group focus on a novel design or significant modification of existing building blocks. Similarly, we can re-categorize each method into the network-level or module-level design.
3.2.1. Network-Level Design
Wu et al. (2018) introduce random warping series, a novel time-series representation model with a positive-definite kernel based on the dynamic time warping algorithm to capture similarity between time series. For graph time series, STANE (Liu et al., 2019) guides the context sampling process to focus on the crucial part of the data in the graph attention networks. Moreover, to interpret the network embedding results, STANE enables users to investigate the graph context distributions along nodes, training window length, and time. For time series domains that exhibit a common causal structure but possess varying time lags, SASA (Cai et al., 2021) aligns the source and target representation spaces by establishing alignment on intra-variable and inter-variable sparse associative graphs. Lee et al. (2022a) present a novel BMA-Memory framework for bimodal representation learning, focusing on sound and image data. This memory system allows for the association of features between different modalities, even when data pairs are weakly paired or unpaired. DTS (Li et al., 2022) is a disentangled representation learning framework for semantic meanings and interpretability through two disentangled components. The individual factor disentanglement extracts different semantic independent factors. The group segment disentanglement makes a batch of segments to enhance group-level semantics.
Unlike others, HyperTime (Fons et al., 2022) introduces implicit neural representation for times series used for imputation and reconstruction by taking timestamps as input and outputting the original time series. It consists of two networks Set Encoder and HyperNet Decoder. From another different perspective, Zhang et al. (2023b) redesign standard multi-layer encoder-decoder sequence models to learn time-series processes as state-space models via the companion matrix and a new closed-loop view of the state-space models, resulting in a new module named SpaceTime layer, which consists of 1D-CNN and feed-forward networks. Multiple SpaceTime layers are stacked to form the final encoder-decoder architecture. HierCorrPool (Wang et al., 2023a) is a multivariate time-series representation learning framework that captures both hierarchical correlations and dynamic properties by using a novel hierarchical correlation pooling scheme and sequential graphs. Liang et al. (2023b) propose CSL, a representation learning framework based on a unified shapelet-based encoder with multi-scale alignment to transform raw multivariate time series into a set of shapelets and learning the representations using them. COMET (Wang et al., 2023b) consists of observation-level, sample-level, trial-level, and patient-level contrastive blocks to represent multiple levels of medical time series. These blocks help to capture cross-sample features, making the learned representations more robust. To better model multi-scale temporal patterns and inter-channel dependencies, MSD-Mixer (Zhong et al., 2023) employs MLPs along different dimensions to mix intra- and inter-patch variations (generated from a novel multi-scale temporal patching approach) and channel-wise correlations. It learns to explicitly decompose the input time series into different components by generating the component representations in different layers, and accomplishes the analysis task based on such representations.
To handle irregular time series, the temporal kernelized autoencoder (Bianchi et al., 2019) is proposed to learn representations aligned to a kernel function designed for handling missing values. Continuous recurrent units (Schirmer et al., 2022) update hidden states based on linear stochastic differential equations, which are solved by the continuous-discrete Kalman filter. Based on continuous-discrete filtering theory, Ansari et al. (2023) introduce a neural continuous-discrete state space model to model continuous-time modeling of irregularly-sampled time series. Biloš et al. (2023) suggest a representation learning approach based on denoising diffusion models adapted for irregularly-sampled time series with complex dynamics. By gradually adding noise to the entire function, the model can effectively capture underlying continuous processes.
3.2.2. Module-Level Design
LIME-RNN (Ma et al., 2019) introduces a weighted linear memory vector into RNNs for time-series imputation and prediction. Following TST (Zerveas et al., 2021), Chowdhury et al. (2022) propose TARNet, a new representation learning model based on Transformers. TARNet aims to reconstruct important timestamps using a newly designed masking layer to improve downstream task performance. It decouples data reconstruction from the downstream task and uses a data-driven masking strategy instead of random masking via self-attention score distribution generated by the transformer encoder during the downstream task training to determine a set of important timestamps to mask. As a result, the reconstruction process becomes task-aware. Another recent method using the wavelet transform, called WHEN (Wang et al., 2023c), newly designs two types of attention modules, i.e., WaveAtt and DTWAtt modules. In the WaveAtt module, the study proposes a novel data-dependent wavelet function to analyze dynamic frequency components in non-stationary time series. In the DTWAtt module, WHEN transforms the dynamic time warping (DTW) technique into the form of the DTW attention. Here, all input sequences are synchronized with a universal parameter sequence to overcome the time distortion problem in multiple time series. Then, the outputs from the new modules are further combined with task-dependent neural networks to perform the downstream tasks, such as classification and forecasting. Due to the proliferation of edge devices, another recent study (Gorbett et al., 2023) proposes a novel model compression technique to make lightweight Transformers for multivariate time-series problems using network pruning, weight binarization, and task-specific modification of attention modules that can substantially reduce both model size (i.e., # of parameters) and computational complexity (i.e., # of multiply-adds or FLOPs). These compressed neural networks have the potential to enable DL-based models across new applications and smaller computational environments. The paper also demonstrates that the compressed Transformers using the proposed technique achieve comparable accuracy to their original counterparts (i.e., Dense Transformers (Zerveas et al., 2021)) despite the substantial reduction in model size and computational complexity. In addition, it claims that this is the first work applying sparse and binary-weighted Transformers to multivariate time-series analysis, including forecasting, classification, and anomaly detection. NuTime (Lin et al., 2023) is a window-wise data embedding for one window to deal with time-series pre-training. This embedding module consists of normalized shape embedding and numerically multi-scaled embedding based on window mean and window standard deviation.
A few studies in this group also proposed methods to handle irregularly-sampled time series. mTAN (Shukla and Marlin, 2021) learns the representation of continuous time values by applying the attention mechanism to irregularly observed time series. TE-ESN (Sun et al., 2021) employs a requisite time encoding mechanism to acquire knowledge from irregular time series data, with the representation being learned within echo state networks. Naour et al. (2023) introduce TimeFlow to deal with modeling issues such as irregular time steps. With a hyper-network, the framework modulates implicit neural representation that is a parameterized continuous function on multiple time series.
Unlike learning standard time-series data, Sener et al. (2020) suggest a flexible multi-granular temporal aggregation framework with non-local blocks, coupling block, and temporal aggregation block to solve reasoning from current and past observations for long-range videos. TCGL (Liu et al., 2022) is a self-supervised learning framework to consider multi-scale temporal dependencies within videos. TCGL is trained on a temporal contrastive graph by jointly modeling the inter-snippet and intra-snippet temporal dependencies through a hybrid graph contrastive learning strategy.
Remark. The methods discussed in this subsection represent a diverse range of innovative approaches to designing neural architectures for time-series representation learning, with a focus on either network-level or module-level design. These approaches encompass novel techniques such as random warping series, hierarchical correlation pooling, disentangled representation learning, and continuous-discrete state space models. Additionally, the application of Transformers in representation learning showcases the adaptability of attention mechanisms to time-series analysis. The exploration of model compression techniques for lightweight Transformers addresses the challenges posed by edge devices. Overall, these advancements contribute significantly to enhancing the interpretability, efficiency, and performance of representation learning for diverse time series.
4. Learning-focused Approaches
Studies in this category focus on devising novel objective functions or pretext tasks used for the representation learning process, i.e., model training. The learning objectives can be categorized into supervised, unsupervised, or self-supervised learning, depending on the use of labeled instances. In our survey, the difference between unsupervised and self-supervised learning is the presence of pseudo labels. Specifically, unsupervised learning is based on the reconstruction of its input, while self-supervised learning uses pseudo labels from pretext tasks as self-supervision signals.
4.1. Supervised Learning
Due to the requirement of labeled instances in supervised learning, a supervised representation learning algorithm trains the parameterized feature encoder $f_{e}$ by solving a supervised task, such as classification, forecasting, and regression, on $\mathcal{D}=\{(\mathbf{X}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$. It aims to learn generic feature extractor $f_{e}:\mathbb{R}^{T\times V}\to\mathbb{R}^{R\times F}$ by optimizing a supervised loss $\mathcal{L}$ on a labeled dataset $\mathcal{D}$ automatically without manual feature engineering. Ma et al. (2019) develop a set of task-specific loss functions to train the LIME-RNN with incomplete time series in an end-to-end way, simultaneously achieving imputation and prediction. TriBERT (Rahman et al., 2021) designs separate losses for a specific modality. Weakly supervised classification for vision-based modalities and classification loss for audio modality. Hadji et al. (2021) use a soft version of dynamic time warping (soft-DTW) to compute the loss between two videos with the same class in the weakly-supervised setting. Cycle-consistency is forced by matching two DTW results ($X\to Y$ and $Y\to X$) as the same. Last, Zhong et al. (2023) propose MSD-Mixer with a novel loss function to constrain both the magnitude and auto-correlation of the decomposition residual used together with the supervised loss of the target downstream task during training. This loss function facilitates MSD-Mixer to extract more temporal patterns into the components to be used for the downstream task.
Remark. Supervised learning methods leverage labeled instances to train models on specific objective functions tailored for downstream tasks like classification. Notably, recent advancements include task-specific loss functions for handling incomplete time series and distinct losses for different modalities such as vision and audio. However, supervised learning cannot attract much attention from the universal representation learning community due to the insufficient labels in time-series datasets (as opposed to large datasets like ImageNet for vision tasks) and the limited generalizability of the trained models.
4.2. Unsupervised Learning
Unsupervised representation learning does not require labeled samples to learn the encoder $f_{e}$, hence $\mathcal{D}=\{\mathbf{X}_{i}\}_{i=1}^{N}$. The difference from supervised representation learning is that unsupervised representation learning trains $f_{e}$ by solving an unsupervised task, such as reconstruction and masked input prediction (self-prediction), on $\mathcal{D}$ using an unsupervised loss function. Compared to supervised learning, unsupervised representation learning methods are more practical since they do not require labeled instances.
4.2.1. Reconstruction Loss
The reconstruction loss computes the difference between the input and output, i.e., reconstructed input. Autoencoders are representative models that leverage the reconstruction loss comprising encoder and decoder networks. The encoder extracts latent representations, and the decoder subsequently reconstructs the inputs from the extracted representations.
Sqn2Vec (Nguyen et al., 2018) is an unsupervised method that learns low-dimensional continuous feature vectors for sequence data by predicting singleton symbols and sequential patterns to satisfy a gap constraint. Chen et al. (2019) introduce unsupervised training of Audio Word2vec using a sequence-to-sequence autoencoder (SA) on unannotated audio time series. It presents a new segmental audio word2vec operating at the utterance level, combining unsupervised spoken word boundary segmentation and audio word2vec through a segmental SA. The encoder includes a segmentation gate trained with reinforcement learning to represent utterances as vectors carrying phonetic information. Yuan et al. (2019) propose Wave2Vec to jointly model inherent and temporal representations of biosignals and provide clinically meaningful interpretation with enhanced interpretability. Sanchez et al. (2019) propose a method for satellite image time series by combining unsupervised learning techniques (variational autoencoder and generative adversarial networks) to learn image-to-image translation, where one image in a time series is translated into another.
More recently, Tonekaboni et al. (2022) present a generative method by employing variational approximation to decouple local and global representations of time-series data with counterfactual regularization that minimizes the mutual information between the local and global variables. Ti-MAE (Li et al., 2023), a masked autoencoder framework, addresses the distribution shift problem by learning strong representations with less inductive bias or hierarchical trick. Its masking strategy creates different views for the encoder in each iteration, fully leveraging the whole input time series during training. Similarly, Dong et al. (2023) propose SimMTM, another masked autoencoder. SimMTM reconstructs the original time series from multiple masked time series with series-wise similarity learning and point-wise aggregation to implicitly reveal the local structure of the manifold. To do so, it introduces a neighborhood aggregation design for reconstruction by aggregating the point-wise representations of time series based on the similarities learned in the series-wise representation space. Thus, the masked time points are recovered by weighted aggregation of multiple neighbors outside the manifold, allowing SimMTM to assemble complementary temporal variations from multiple masked time series and improve the quality of the reconstructed time series. In addition to the reconstruction loss, a constraint loss is proposed to guide the series-wise representation learning based on the neighborhood assumption of the time series manifold.
4.2.2. Masked Prediction
Masked prediction primarily focuses on predicting the masked segment of the input rather than reconstructing the original one. Therefore, encoder/decoder-only architectures are effective in addressing masked prediction.
To achieve unsupervised learning, TST (Zerveas et al., 2021) adopts an encoder-only Transformer network with a masked prediction (denoising) objective by computing. The losses of TST are computed only from the masked parts or timestamps. Specifically, TST leverages unlabeled time series by training the Transformer encoder to extract dense vector representations of multivariate time series using the denoising objective on randomly masked input time series. Similar to TST, TARNet (Chowdhury et al., 2022) improves downstream task performance by learning to reconstruct important timestamps using a data-driven masking strategy. The masked timestemps are determined by self-attention scores during the downstream task training. This reconstruction process is trained alternately with the downstream task at every epoch by sharing parameters in a single network. It also enables the model to learn the representations via task-specific reconstruction, which results in improved downstream task performance. For a masked token pretext task, UniTTab (Luetto et al., 2023) uses row and timestamp masking with neighborhood label smoothing on the time-dependent heterogeneous tabular data during the pre-training stage to address the heterogeneity problems.
Remark. Unsupervised representation learning, unlike its supervised counterpart, trains an encoder without relying on labeled data, making it more practical for various datasets and applications. A range of methods employ techniques like reconstruction loss and masked prediction to extract meaningful features from data. These approaches underscore the versatility and potential of unsupervised learning in capturing intricate patterns and structures from diverse datasets.
4.3. Self-Supervised Learning
Self-supervised representation learning is similar to the unsupervised learning in the sense that it does not require labeled data. However, instead of learning by using the direct input, the loss function of self-supervised learning (i.e., pretext task) is trained with self-supervised signals, called pseudo labels, on the dataset $\mathcal{D}=\{(\mathbf{X}_{i},\hat{\mathbf{y}})\}_{i=1}^{N}$ where $\hat{\mathbf{y}}$ indicate the pseudo labels. Due to the high labeling cost, there is a surge in the popularity of self-supervised learning algorithms. Following the impressive performance of contrastive learning in computer vision, contrasting loss has gained attention in time-series analysis, demonstrating excellent performance. Here, we review both non-contrastive and contrastive self-supervised learning methods.
4.3.1. Non-Contrasting Loss
Non-contrasting loss exploits the inherent structures, relationships, or patterns in the data as the effective supervision signals for training. For example, making the model to predict whether a given segment is the past or future data based on a reference time-series segment. Haresh et al. (2021) leverage soft dynamic time warping loss between two videos to learn representation in the unsupervised manner for temporal alignment. Wang et al. (2021) propose a novel pretext task to address the self-supervised video representation learning using motion-aware curriculum learning. Several spatial partitioning patterns are employed to encode rough spatial locations instead of exact spatial Cartesian coordinates. In recent work, Liang et al. (2022) introduce three novel pretext tasks, including (1) clip continuity prediction, (2) discontinuity localization, and (3) missing section approximation, based on video continuity by cutting a clip into three consecutive short clips. CACL (Guo et al., 2022) aims to capture temporal details of a video by learning to predict an approximate Edit distance between a video and its temporal shuffle. In addition, contrastive learning is performed by generating four positive samples from two different encoders, 3D CNN and video transformer, as well as two different augmentations. Duan et al. (2022) introduce TransRank with a pretext task aiming at assessing the relative magnitude of transformations. It allows the model to capture inherent characteristics of the video, such as speed, even when the challenge of matching the transformations across videos varies. Fang et al. (2023) propose a multivariate time-series modeling method to capture the spatial relation shared across all instances. A prior graph is learned to minimize the distance between dynamic graph. On the other hand, to capture the instance-specific spatial relation, a dynamic graph is learned to maximize the distance between prior graph. Another recent work T-Rep (Fraikin et al., 2023) is claimed to be the first self-supervised framework for time series to leverage the time-embedding module in its pretext tasks which enable the model to learn fine-grained temporal dependencies, giving the latent space a more coherent temporal structure than existing methods.
4.3.2. Contrasting Loss
Contrastive learning involves learning to differentiate between positive and negative samples, assuming that the positive samples should have similar representations among each other while being different from the negative ones. Therefore, the design of the contrastive loss and the selection of positive and negative samples are crucial for contrasting-based methods.
Franceschi et al. (2019) employ the triplet loss (i.e., T-Loss) inspired by Word2Vec (Goldberg and Levy, 2014) for learning scalable representations of multivariate time series. It considers a sub-segment belonging to the input time segment as a positive sample to explore sub-series consistency. From another perspective, Eldele et al. (2021) introduce a self-supervised time-series representation learning framework via temporal and contextual contrasting (TS-TCC). Time series are first transformed into two different yet correlated views by weak and strong augmentations. Then, TS-TCC uses a novel temporal contrasting module to learn temporal dependencies by designing a hard cross-view prediction task that utilizes the past latent features of one augmentation to predict the future of another augmentation for a certain time step. This operation forces the model to learn robust representation by a harder prediction task against any perturbations introduced by different time steps and augmentations. Last, it proposes a contextual contrasting module built upon the contexts from the temporal contrasting module to further learn discriminative representations by maximizing the similarity among different contexts of the same sample while minimizing similarity among contexts of different samples. TS-TCC shows high efficiency in few-labeled data and transfer learning scenarios. TNC (Tonekaboni et al., 2021) introduces a contrastive learning framework for complex multivariate non-stationary time series. This approach aims to capture the progression of the underlying temporal dynamics. It takes advantage of the local smoothness of a signal’s generative process to define neighborhoods in time with stationary properties. The neighborhood boundaries are determined automatically using the properties of the signal and augmented Dickey-Fuller statistical testing. Additionally, TNC learns the representations by ensuring that the distribution of signals from within a neighborhood is distinguishable from the distribution of non-neighboring signals in the latent space using a debiased contrastive loss. This loss incorporates positive unlabeled learning for sample weight adjustment to account for potential bias introduced in sampling negative examples.
Yue et al. (2022) propose a universal framework for learning representations of time series in an arbitrary semantic level, called TS2Vec. TS2Vec is based on novel contextual consistency learning using two contrastive policies over two augmented time segments with different contexts from randomly overlapped subsequences. Unlike other methods, TS2Vec performs contrastive learning in a hierarchical way over the augmented context views, making a robust contextual representation for each timestamp. Its overall representation can be obtained by max pooling over the corresponding timestamps. By using multi-scale contextual information with granularity, this approach can capture multi-resolution/multi-scale contextual information with both temporal and instance-wise contrastive loss for the given time series and generate fine-grained representations for any granularity. Similar to BTSF (Yang and Hong, 2022), Zhang et al. (2022b) introduce a strategy for self-supervised pre-training in time series by modeling time-frequency consistency (TF-C). This study argues that time-based and frequency-based representations, learned from the same time series, should be closer to each other in the time-frequency latent space than representations of different time series. TF-C aims to minimize the distance between time-based and frequency-based embeddings using a novel consistency loss. TimeCLR (Yang et al., 2022a) is a self-supervised contrastive learning framework, enabling the feature extractor to learn invariant representations by minimizing the similarity between two augmented views of the same sample. Hajimoradlou et al. (2022) introduce a self-supervised framework with similarity distillation along the temporal and instance dimensions for pre-training universal representations. For robot sensor data, Somaiya et al. (2022) propose a simple representation learning framework, TS-Rep, based on T-Loss (Franceschi et al., 2019). TS-Rep considers real-world scenarios of robotics by making the model available on varying length time series without any padding techniques.
CSL (Liang et al., 2023b) proposes to learn time-series representations shapelets with multi-grained contrasting and multi-scale alignment for capturing information in various time ranges. Nguyen et al. (2023) design a novel contrastive loss function by combining ideas from hierarchical loss (Yue et al., 2022) and triplet loss. FEAT (Kim et al., 2023a) uses a combination of contrastive losses to jointly learn both feature-based consistency and temporal consistency by using hierarchical temporal contrasting loss, feature contrasting loss, and reconstruction loss. Choi and Kang (2023) introduce uncertainty weighting approach to weigh multiple contrastive loss by considering homoscedastic uncertainty of multiple tasks, including contextual, temporal, and transformation consistencies. FOCAL (Liu et al., 2023) enforces the modality consistency to learn the features shared across modalities and the transformation consistency to learn the modality-specific feature. Additionally to accommodate sporadic deviations from locality due to periodic patterns, temporally close sample pairs and distant sample pairs are constrained by a loose ranking loss. COMET (Wang et al., 2023b) composes four-level contrastive losses with respect to medical time series. To compromise these multiple losses, the overall loss has hyper-coefficients about each loss. TS-CoT (Zhang et al., 2023e) integrates contrastive learning into a multi-view time series representation learning framework, emphasizing the importance of global consistency and complementary information to enhance robustness against noise. By utilizing the ability of large language models (LLM), TEST (Sun et al., 2023), an embedding method grounded in contrastive learning, enables LLMs to handle time-series analysis tasks effectively by leveraging intrinsic information and avoiding the need for extensive pre-defined knowledge or annotations.
Many studies also employ contrastive learning to learn representations of irregular time-series data. For instance, TimeAutoML (Jiao et al., 2020) adopts the AutoML framework, enabling automated configuration and hyperparameter optimization. This approach utilizes a contrastive learning framework where negative samples are created by introducing random noise within the range defined by the minimum and maximum values of the given instances. Another work named PrimeNet (Chowdhury et al., 2023) adopts a contrastive learning approach by generating augmented samples based on the observation density and employs a reconstruction task to facilitate the learning of irregular patterns.
For video representations, Morgado et al. (2020) leverages 360° video data with strong spatial cues and conducts audio-visual spatial alignment as a pretext task. This pretext task involves spatially misaligned audio and video clips, treated as negative examples for contrastive learning. MemDPC (Han et al., 2020) constitutes a self-supervised learning approach, specifically emphasizing representations for action recognition. Within this framework, training MemDPC involves a predictive attention mechanism applied to a collection of compressed memories. This training paradigm ensures that any subsequent states can be synthesized consistently through a convex combination of the condensed representations. Wang et al. (2020) suggest a pretext task that classifies the pace of a clip into five categories: super slow, slow, normal, fast, and super fast. Contrastive learning is adopted by selecting positive pairs from a pair of clips in the same video and negative pair from a pair of clips in different videos. CCL (Kong et al., 2020) uses the inclusive relation of a video and frames for a contrastive learning strategy. The video and frames of the inclusive relation are learned to be close to each other in the video and frame embedding spaces. In addition to the non-contrastive loss, Haresh et al. (2021) jointly use a temporal regularization term (i.e., Contrastive-IDM) to encourage different frames to be mapped to different points in the embedding space. Jenni and Jin (2021) propose a time-equivariant contrastive learning model. In this framework, a pair of clips is the unit of contrastive learning. If two pairs have same temporal transformation within each pair, then the output of each clip is concatenated for each pair and the concatenated outputs become similar. Auxiliary tasks are also suggested such as classifying each temporal difference between two clips: predicting clip A as 2x speed and clip B as slow speed. CVRL (Qian et al., 2021) is a contrastive learning framework that uses the (temporally consistent) spatial augmentation and clip selection strategy, where each frames are spatially augmented. Here, two clips in the same video are positive, while two clips in the different videos are negative. RSPNet (Chen et al., 2021c) solves the relative speed perception task and the appearance-focused video instance discrimination task with self-supervised representation learning by using a triplet loss and the InfoNCE loss.
Concerning other modalities, Lee et al. (2022a) employ contrastive learning to effectively associate sound and image modalities with limited paired data to enhance representation learning by encouraging similarity in positive pairs and dissimilarity in negative pairs. CSTP (Zhang et al., 2022a) introduces a pretext task for video representation called spatio-temporal overlap rate prediction that considers the intermediate of contrastive learning. With a joint optimization combining pretext tasks with contrastive learning, CSTP enhances the spatio-temporal representation learning for downstream tasks. DCLR (Ding et al., 2022) presents a dual contrastive formulation for video representation by decoupling the input RGB video sequence into static scene and dynamic motion. TCLR (Dave et al., 2022) is a contrastive learning framework for video understanding tasks. It is trained with novel local–local and global–local temporal contrastive losses. Chen et al. (2022) propose sequence contrastive loss to sample two subsequences with an overlap for each video. The overlapped timestamps are considered positives, while the clips from other videos are negatives. Two timestamps neighboring each other also become positive pairs with the Gaussian weight proportional to the temporal distance. Qian et al. (2022) propose a windowing based contrastive learning by sampling a long clip from a video and a short clip that lies inside the duration of the long clip. These long and short clips becomes a positive pair for contrastive learning and other long clips become negative instances. When conducting contrastive learning, final vectors are made in two different embedding spaces. The first one is a fine-grained space and each embedding is made at each timestamp. The second embedding space is a persistent embedding space and each timestamp embedding is global average pooled for contrastive learning. Qing et al. (2022) conduct representation learning on untrimmed video to reduce the amount of labor required for manual trimming and utilize the rich semantics of untrimmed video. Hierarchical contrastive learning teaches clips that are near in time and topic to be similar. Zhang and Crandall (2022) introduce a self-supervised video representation learning trained with decoupled learning objectives into two contrastive sub-tasks, which are hierarchically spatial and temporal contrast. With graph learning, TCGL (Liu et al., 2022) uses a spatial-temporal knowledge discovering module for motion-enhanced spatial-temporal representations. TCGL introduces intra- and inter-snippet temporal contrastive graphs to explicitly model multi-scale temporal dependencies, employing a hybrid graph contrastive learning strategy. A recent study (Yang et al., 2023a) proposes TempCLR, a contrastive learning framework for exploring temporal dynamics in video-paragraph alignment, leveraging a novel negative sampling strategy based on temporal granularity. By focusing on sequence-level comparison and using dynamic time warping, TempCLR captures temporal dynamics more effectively. Zhang et al. (2023d) model videos as stochastic processes via a novel process-based contrastive learning framework by enforcing an arbitrary frame to agree with a time-variant Gaussian distribution conditioned on the start and end frames.
Remark. Self-supervised representation learning has emerged as a powerful approach in the realm of time-series analysis, particularly due to its ability to mitigate the high labeling cost. The contrastive and non-contrastive self-supervised learning methods discussed showcase a diverse range of strategies, such as temporal contrasting, combination of contrastive consistencies, and dynamic graph modeling, among others, for capturing intricate temporal dependencies and spatial relations in various applications. Whether addressing multivariate time series, irregular patterns, or video representations, the contrastive-based approaches emphasize the importance of selecting meaningful positive and negative samples. The variety of techniques presented underscores the versatility of self-supervised learning in enhancing the efficiency and robustness of representation learning, paving the way for advancements in time-series analysis across different domains.
5. Data-Centric Approaches
In this group, we categorize the methods that focus on finding a new way to enhance the usefulness of the training data at hand. To capture the underlying patterns, trends, and relevant features within the time series, data-centric approaches prioritize engineering the data itself rather than focusing on model architecture and loss function design. We categorize data-centric approaches into three techniques: data augmentation, decomposition and transformation, and sample selection.
5.1. Data Augmentation Strategies
The studies in this category aim to increase the size and diversity of the training data. Especially in representation learning, most contrastive learning methods generate positive and negative samples by data augmentation. Also, different from other data types, time-series data augmentation should consider the unique properties in the time series, such as temporal dependencies, multi-scale dependencies, and inter-variable dependencies. We further categorize each method into random or policy-based augmentation.
5.1.1. Random Augmentation
In addition to hierarchical contrasting loss, TS2Vec (Yue et al., 2022) randomly treats two overlapping time segments as positive pairs for contrastive learning. Specifically, timestamp masking and random cropping are applied to the input time series to generate a context. Here, the contrasting objective in TS2Vec is based on the augmented context views, meaning that representations of the same sub-series in two augmented contexts should be consistent. As a result, we can obtain a robust contextual representation for each sub-series without introducing unappreciated inductive biases like transformation- and cropping-invariance, which proven effective for a wide range of downstream tasks. By considering frequency information, TF-C (Zhang et al., 2022b) aims to use both time and frequency features of time series to improve the representation quality. TF-C is claimed to be the first study to develop frequency-based contrastive augmentation exploiting spectral information and to explore time-frequency consistency in time series. It consists of a set of novel augmentations based on the characteristic of the frequency spectrum and produces a frequency-based embedding through contrastive instance discrimination. More specifically, TF-C introduces frequency-domain augmentation by randomly adding or removing frequency components, thereby exposing the model to a range of frequency variations. Recently, TS-CoT (Zhang et al., 2023e) facilitates the creation of diverse views for contrastive learning by enhancing robustness to noisy time series that contributes to the overall effectiveness of the representation learning.
5.1.2. Policy-based Augmentation
Instead of relying on randomness, several studies devised a specific criterion for data augmentation. TimeCLR (Yang et al., 2022a) addresses the challenges of applying self-supervised learning to time series by introducing dynamic time warping-based augmentation, which is designed to induce phase shifts and amplitude changes in time-series data while preserving its structural and feature information. BTSF (Yang and Hong, 2022) uses the entire time series as input and applies a standard dropout as an instance-level augmentation to generate different views for representation learning of time series. This paper argues that using dropout as the augmentation method can preserve the global temporal information and capture long-term dependencies of time series more effectively. The construction of contrastive pairs also ensures that the augmented time series do not change their raw properties, effectively reducing the potential false negatives and positives.
Luo et al. (2023) propose InfoTS, a meta-learning-based data augmentation method. InfoTS is an adaptive data augmentation method that uses information-aware criteria for selecting the optimal data augmentations to generate feasible positive samples for contrastive learning. The positive samples generated by InfoTS exhibit both high fidelity and diversity, which are essential attributes for contrastive learning. RIM (Aboussalah et al., 2023) is another new time-series augmentation method that generates more samples by using a recursive interpolation function from the original time series to be used in training. This method can control how much the augmented time series trajectory deviates from the original trajectory. Based on consistency regularization, Shin et al. (2023) propose a novel method of data augmentation called context-attached augmentation, which adds preceding and succeeding instances to a target instance to form its augmented instance, to fully exploit the sequential characteristic of time series in consistency regularization. This context-attached augmentation generates instances augmented with varying contexts while maintaining the target instance, thus avoiding direct perturbations on the target instance’s attributes. Another recent method that uses frequency-domain information is proposed by Demirel and Holz (2023). This method is based on a mixup technique for non-stationary quasi-periodic time series. This method aims to connect intra-class samples together to find order in the latent space. With a mixture of the magnitude and phase of frequency components together with tailored operations, it can prevent destructive interference of mixup.
For video data, Chen et al. (2022) use overlap augmentation techniques by sampling two subsequences with an overlap for each video. Overlapped timestamps are positives, and clips from other videos are negatives. Two timestamps neighboring each other also become positive pairs with Gaussian weight proportional to temporal distance. Zhang and Crandall (2022) introduce a novel self-supervised video representation learning technique emphasizing spatial and temporal features through a contrastive approach. Augmentation is used as a regularization tool, guiding the network to learn desired semantics in contrastive learning. FreqAug (Kim et al., 2023c) pushes the model to focus more on dynamic features via dropping spatial or temporal low-frequency components. For various downstream tasks, FreqAug stochastically eliminates particular frequency components from the video to help the latent representation better capture meaningful features from the remaining data. Kim et al. (2023b) propose DynaAugment by dynamically changing the augmentation magnitude over time in order to learn the temporal variations of a video. By Fourier sampling the magnitudes, it regulates the augmentations diversely while maintaining the temporal consistency.
Remark. Recent advancements in time-series representation learning have emphasized the importance of augmentation strategies tailored to capture both temporal and frequency characteristics effectively. While random augmentation methods leverage augmented context views for robust representation, policy-based approaches employ specific criteria to generate diverse and high-fidelity positive samples for contrastive learning. These augmentation techniques not only enhance the model’s ability to handle noisy and diverse time-series data but also ensure the preservation of essential temporal and spectral attributes.
5.2. Decomposition and Transformation
Unlike the above technique that increases the number of samples, methods in this subcategory aim to extract more meaningful information from the training data.
5.2.1. Time-Series Decomposition
mWDN (Wang et al., 2018) is an early attempt that integrates multi-level discrete wavelet decomposition into a deep neural network framework. This framework allows for the preservation of frequency learning advantages while facilitating the fine-tuning of parameters within a deep learning context. Fang et al. (2023) decompose the spatial relation within multivariate time series into prior and dynamic graphs to model both common relation shared across all instances and distinct correlation that varies across instances.
To enhance video-level tasks, including video classification, and more detailed tasks like action segmentation, Behrmann et al. (2021) separate the representation space into stationary and non-stationary characteristics through contrastive learning from long and short views. Zeng et al. (2021) improve generalization across various downstream tasks by learning spatially-local/temporally-global and spatially-global/temporally-local features from audio-visual modalities to capture global and local information inside a video. This method enables the capturing of both slowly changing patch-level information and fast changing frame-level information. More recently, Yang et al. (2022b) propose a unified framework for joint audio-visual speech recognition and synthesis. Each modality is decomposed into modality-specific and modality-invariant features. Modality-invariant codebook enhances the alignment between the linguistic representation space of visual and audio modalities.
5.2.2. Input Space Transformation
To make effectively exploit vision models, Anand and Nayak (2021) transform 1D time series into 2D images and use the transformed images as features for the subsequent learning phase with models pretrained on large image datasets, such as ResNet50. Similarly, TimesNet (Wu et al., 2023) learns the temporal variations in the 2D space to enhance the representation capability by capturing variations both within individual periods and across multiple periods in the time series. It analyzes the time series from a new dimension of multi-periodicity by extending the analysis of temporal variations from 1D into 2D space. Besides, by transforming the 1D time series into a set of 2D tensors, TimesNet breaks the bottleneck of representation capability in the original 1D space, enabling well-performing vision backbones applicable to the transformed time series.
To address tasks involving irregular time series data, some input space transformation techniques have been proposed. SplineNet (Biloš et al., 2022) generates splines from the input time series and directly utilizes them as input for a neural network. This approach introduces a learnable spline kernel to effectively process the input spline. MIAM (Lee et al., 2022b) considers multiple views of the input data, including time intervals, missing data indicators, and observation values. These transformed input data are then processed by the multi-view integration attention module to solve downstream tasks. Recently, Zhong et al. (2023) propose a novel multi-scale temporal patching approach to divide the input time series into non-overlapping patches along the temporal dimension in each layer. It treats the time series as patches. Different layers have different patch sizes such that they can focus on different time scales, making the representations more expressive.
For temporal graph networks, Chanpuriya et al. (2023) construct time-decayed line graph to train edge representation by creating a sparse matrix of proximities between temporal edges.
Remark. Time-series decomposition and input space transformation have significantly enhanced the capabilities of deep models for analyzing and processing time-series and video data. Notably, innovations in transforming 1D time series into 2D images or tensors, as well as methods addressing irregular time series, have expanded the applicability and performance of vision backbones.
5.3. Methods for Sample Selection
Methods in this group aim to generate or select the most useful samples based on the available training data for enhancing the informativeness of the learned representations.
5.3.1. Generative Methods
Sample generation is a popular technique that increases the size and diversity of the training data when the data is scarce. This line of work explicitly generate new samples by transformation or generative models. For enhancing the noise resilience, Nguyen et al. (2023) propose a novel noise-resilient sampling strategy by exploiting a parameter-free discrete wavelet transform low-pass filter to generate a perturbed version of the original time series. By leveraging the large language model (LLM), LAVILA (Zhao et al., 2023) learns better video-language embedding when only few text annotations are available. Using the available video-text data, it first fine-tunes LLM to generate text narration given visual input. Then densely annotated videos via a fine-tuned narrator are used for video-text contrastive learning.
5.3.2. (Re-)Sampling Strategies
Sampling strategies aim to effectively select the best samples from available training data for a particular learning scenario. An early study adopting sample selection for time-series representation learning is proposed by Franceschi et al. (2019). This work uses time-based negative sampling with the T-Loss mentioned above. It determine several negative samples by independently choosing sub-series from other time series at random, whereas a sub-series within the referenced time series is consider a positive sample. This technique encourages representations of the input time segment and its sampled sub-series to be close to each other. MTRL (Chen et al., 2021b) utilizes discriminative samples to design a distance-weighted sampling strategy for achieving high convergence speed and accuracy.
Remark. Both generative methods and sampling strategies play pivotal roles in enhancing the efficacy and diversity of training data. Generative techniques offer methods to augment data and improve resilience against noise, especially when data is limited. On the other hand, sampling strategies underscore the importance of judiciously selecting samples to optimize representation learning, emphasizing both convergence speed and accuracy in training. As most of the time-series representation learning methods focus on extracting features useful for the downstream tasks, sample generation may be less popular. Furthermore, compared to other data types such as image and natural language, there is a lack of generative foundation models that can be utilized off-the-shelf for sample generation.
6. Experimental Design
This section describes the typically used experimental design for comparing universal representation learning methods for time series. We describe widely-used protocol and introduce publicly available benchmark datasets with evaluation metrics according to the downstream tasks.
Given a set of $N$ time series $\{(\mathbf{X}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$ and $J$ pre-trained representation learning models $\{f_{e,j}\}_{j=1}^{J}$, this section describes how we evaluate each model to determine the best one. As discussed in Section 1, representations of time series play a vital role in solving time-series analysis tasks. We expect that the learned representations by $f_{e}$ generalize to unseen downstream tasks. Accordingly, the most common evaluation method is how learned representations help solve downstream tasks.
Additionally, we need a function $g_{d}$ that maps a representation (feature) space to a label space. For example, $g_{d}(f_{e}(\mathbf{X})):\mathbb{R}^{R\times F}\to\mathbb{R}^{|C|}$ for classification or $g_{d}(f_{e}(\mathbf{X})):\mathbb{R}^{R\times F}\to\mathbb{R}^{H}$ for forecasting. This is because $f_{e}$ is designed to extract feature representations, not to solve the downstream task. Commonly, $g_{d}$ is implemented as a simple function, such as linear regression, support vector machines, or shallow neural networks because it is enough to solve the downstream task if the learned representations already capture meaningful and discriminative features.
6.1. Evaluation Procedure
Let $\mathcal{D}=\{\mathbf{X}_{i},\mathbf{y}_{i}\}_{i=1}^{N}$ the downstream data. We then compare the encoders $f_{e}$ by using the task-specific evaluation metrics of the downstream task. The evaluation procedures are as follows.
  1. (1)
    Train $f_{e}$ and $g_{d}$ on the downstream dataset $\mathcal{D}$ with the pre-trained encoder $f_{e,j}$ for each $j$.
  2. (2)
    Compare task-specific evaluation metric values, e.g., accuracy for classification or mean square error for regression.
In the first step, there are two common protocols to evaluate the encoders: frozen and fine-tuning.
6.1.1. Frozen Protocol
As we expect $f_{e}$ to learn meaningful representations for downstream tasks, we do not update the pre-trained $f_{e}$, i.e., freezing its weights. Training $g_{d}$ uses less computation budget and converges faster than the training of $f_{e}$ and $g_{d}$ from scratch on the downstream dataset. The standard choice of $g_{d}$ is a linear model (e.g., linear regression and logistic regression) or non-parametric method (e.g., k-nearest neighbors). This evaluation approach is usually referred as linear probing. To further improve performance with additional computing cost, we can also implement $g_{d}$ as a nonlinear model, such as shallow neural networks with a nonlinear activation function.
6.1.2. Fine-Tuning Protocol
In fine-tuning protocol, we train both pre-trained $f_{e}$ and $g_{d}$ as a single model to obtain more performance gain of the downstream task or fill the gap between representation learning and downstream tasks. $g_{d}$ is also known as a task-specific projection head in the representation learning network. This protocol usually uses small learning rate to optimize the model to preserve the original representation quality.
The combination of linear probing and fine-tuning protocol is also possible, especially for out-of-distribution as the encoder’s performance on data sampled from out-of-distribution degrades after fine-tuning. Even though the fine-tuning protocol requires more computing budget than the frozen protocol, it empirically performs better than the frozen protocol (Nozawa and Sato, 2022).
6.2. Benchmark Datasets and Metrics for Downstream Tasks
We summarize widely-used benchmark datasets and evaluation metrics according to the downstream tasks. Some of these datasets are single-purpose datasets for a particular downstream task, and some are general-purpose time-series datasets that we can use for model evaluation across different tasks. Table 3 presents useful information of the reviewed datasets, including dataset names, dimensions, sizes, application domains, and reference sources.
6.2.1. Forecasting and Imputation
Since the output of forecasting and imputation tasks are numerical sequences, most studies uses the same benchmark datasets for both tasks. Commonly used datasets are from several application domains and services, including electricity (e.g., ETT (Zhou et al., 2021)), transportation (e.g., Traffic (Wu et al., 2023)), meteorology (e.g., Weather (Wu et al., 2023)), finance (e.g., Exchange (Lai et al., 2018)) and control systems (e.g., MoJoCo (Jhin et al., 2022)). To facilitate the evaluation of forecasting models, Godahewa et al. (2021) also introduce a publicly accessible archive for time-series forecasting. Given the numerical nature of the predicted results, the most commonly used metrics are mean squared error (MSE) and mean absolute error (MAE).
6.2.2. Classification and Clustering
As the classification and clustering tasks both aim to identify the real category to which a time-series sample belongs, existing studies usually use the same set of benchmark datasets to evaluate the model performance. Curated benchmarks comprising heterogeneous time series from various application domains, such as UCR (Dau et al., 2019) and UEA (Bagnall et al., 2018), are the most widely used because they can provide a comprehensive evaluation regarding the generalization of the model being evaluated. Besides, many researchers also use human activity (e.g., HAR (Anguita et al., 2013)) and health (e.g., PhysioNet Sepsis (Reyna et al., 2020)) related datasets due to their practicality for real-world applications. We recommend referring to Table 3 for the exhaustive list (including audio and video modalities) of datasets for the classification and clustering tasks.
Regarding the evaluation metrics, while we can evaluate the classification task with accuracy, precision, recall, and F1 score, we usually assess the clustering task with Silhouette score, adjusted rand index (ARI), and normalized mutual information (NMI) to assess the inherent clusterability due to the absence of label instances. For classification, we may also use the area under the precision-recall curve (AUPRC) to handle the class imbalance cases.
6.2.3. Regression
Compared to the forecasting and classification tasks, time-series regression, particularly with deep learning, remains relatively underexplored. Only a handful of public benchmark datasets (e.g., heart rate monitoring data (Demirel and Holz, 2023) and air quality (Zhang et al., 2023c)) exist. The TSER archive, introduced by Tan et al. (2021b), is the most comprehensive benchmark for time-series regression. Like forecasting and imputation tasks, the metrics for regression are MSE and MAE. Additional metrics, such as root mean squared error (RMSE) and R-squared ($R^{2}$), are also commonly used.
6.2.4. Segmentation
Likewise, time-series segmentation with deep learning is also relatively underexplored. There are two standard curated benchmarks: UTSA (Gharghabi et al., 2017) and TSSB (Ermshaus et al., 2023). To assess the segmentation performance, F1 and covering scores are typically used. The F1 score emphasizes the importance of detecting the correct timestamps at which the underlying process changes. In contrast, the covering score focuses on splitting a time series into homogeneous segments and reports a measure for the overlaps of predicted versus labeled segments.
6.2.5. Anomaly Detection
Anomaly detection is one of the most popular research topics in time series. There are several benchmarks publicly available, as listed in Table 3. However, as argued by recent studies (Wu and Keogh, 2021; Wagner et al., 2023), most existing benchmarks are deeply flawed and cannot provide a meaningful assessment of the anomaly detection models. Therefore, we recommend using newly proposed datasets, such as ASD (Li et al., 2021), TimeSeAD (Wagner et al., 2023), and TSB-UAD (Paparrizos et al., 2022a).
Concerning the evaluation metrics, point-adjust F1 score (Xu et al., 2018) is the most widely used metric for time-series anomaly detection. Nevertheless, this metric is also found to have an overestimation problem, which cannot give a reliable performance evaluation. Accordingly, recent studies (Yang et al., 2023b; Nam et al., 2023; Sylligardos et al., 2023) have started to adopt more robust evaluation metrics, e.g., VUS (Paparrizos et al., 2022b), PA%K (Kim et al., 2022), and eTaPR (Hwang et al., 2022).
6.2.6. Retrieval
Although we find that a few studies of the reviewed articles use particular datasets for the retrieval task (e.g., EK-100 (Damen et al., 2022), Howto100M (Miech et al., 2019), and MUSIC (Zhao et al., 2018)), the task itself can be evaluated with any benchmark dataset because it is basically based on arbitrary query time series. For time-series retrieval (Chen et al., 2021b), benchmark datasets for classification (e.g., UCR) are commonly used. For the evaluation, the top-$k$ recall rate (higher is better) is the standard metric to examine the overlap percentage of the top-$k$ results and the ground truth. $k$ is usually set to 5, 10, and 20.
Table 3. Summary of public datasets widely used for time-series representation learning. $T$ and $V$ indicate the varying time-series length (or # frames) and number of variables (or video resolution) per sample, respectively.
Downstream Tasks Dataset Name Size Dimension Domain Modality Reference Source
Forecasting & Imputation ETTh 14,307 7 Electric Power time series (Zhou et al., 2021)
ETTm 57,507 7 Electric Power time series (Zhou et al., 2021)
Electricity 26,304 321 Electricity Consumption time series (Wu et al., 2023)
Traffic 17,451 862 Transportation time series (Wu et al., 2023)
PEMS-BAY 16,937,179 325 Transportation spatiotemporal (Li et al., 2018)
METR-LA 6,519,002 207 Transportation spatiotemporal (Li et al., 2018)
Weather 52,603 21 Climatological Data time series (Wu et al., 2023)
Exchange 7,588 8 Daily Exchange Rate time series (Lai et al., 2018)
ILI 861 7 Illness time series (Wu et al., 2023)
Google Stock 3,773 6 Stock Prices time series (Yoon et al., 2019)
Monash TSF $30\times T$ $V$ Multiple time series (Godahewa et al., 2021)
Solar 52,560 137 Solar Power Production time series (Lai et al., 2018)
MoJoCo 10,000 x 100 14 Control Tasks time series (Jhin et al., 2022)
USHCN-Climate 386,068 5 Climatological Data time series (Schirmer et al., 2022)
Classification & Clustering UCR $128\times T$ 1 Multiple time series (Dau et al., 2019)
UEA $30\times T$ $V$ Multiple time series (Bagnall et al., 2018)
PhysioNet Sepsis $40,336\times T$ 34 Medical Data time series (Reyna et al., 2020)
PhysioNet ICU $12,000\times T$ 36 Medical Data time series (Silva et al., 2012)
PhysioNet ECG $12,186\times T$ 1 Medical Data time series (Clifford et al., 2017)
HAR 10,299 9 Human Activity time series (Anguita et al., 2013)
EMG 163 1 Medical Data time series (PhysioBank, 2000)
Epilepsy 11,500 1 Brain Activity time series (Andrzejak et al., 2001)
Waveform 76,567 2 Medical Data time series (Zhang et al., 2023e)
Gesture 440 3 Hand Gestures time series (Liu et al., 2009)
MOD 39,609 2 Moving Object time series (Liu et al., 2023)
PAMAP2 9,611 10 Human Activity time series (Liu et al., 2023)
Sleep-EEG 371,005 1 Sleep Stages time series (Dong et al., 2023)
RealWorld-HAR 12,887 9 Human Activity time series (Liu et al., 2023)
Speech Commands 5,630,975 20 Spoken Words audio (Jhin et al., 2022)
LRW 13,050,000 $64\times 64$ Lib Reading video (Chung and Zisserman, 2017)
ESC50 10,000 1 Environmental Sound audio (Piczak, 2015)
UCF101 333,000 $320\times 240$ Human Activity video (Soomro et al., 2012)
HMDB51 $6,849\times T$ $V_{width}\times V_{height}$ Human Activity video (Kuehne et al., 2011)
Kinetics-400 7,656,125 $V_{width}\times V_{height}$ Human Activity video (Kay et al., 2017)
AD 1,527,552 16 Medical Data time series (Wang et al., 2023b)
PTB 18,711,000 15 Medical Data time series (Wang et al., 2023b)
TDBrain 3,035,136 33 Brain Activity time series (Wang et al., 2023b)
MMAct
36,764
(only number of instances)
N/A Human Activity multi-modality (Kong et al., 2019)
PennAction $2,326\times T$ $640\times 480$ Human Activity video (Chen et al., 2022)
FineGym $4,883\times T$ $V_{width}\times V_{height}$ Human Activity video (Chen et al., 2022)
Pouring $84\times T$ $V_{width}\times V_{height}$ Human Activity video (Chen et al., 2022)
Something-Something 2,650,164 $84\times 84$ Human Activity video (Kim et al., 2023b)
Regression TSER Archive $19\times T$ $V$ Multiple time series (Tan et al., 2021b)
Neonate $79\times T$ 18 Neonatal EEG Recordings time series (Stevenson et al., 2019)
IEEE SPC $22\times T$ 5 Heart Rate Monitoring time series (Demirel and Holz, 2023)
DaLia $15\times T$ 11 Heart Rate Monitoring time series (Demirel and Holz, 2023)
IHEPC 2,075,259 1 Electricity Consumption time series (Franceschi et al., 2019)
AEPD 19,735 29 Appliances Energy time series (Zhang et al., 2023c)
BMAD 420,768 6 Air Quality time series (Zhang et al., 2023c)
SML2010 4,137 18 Smart home time series (Zhang et al., 2023c)
Segmentation TSSB $75\times T$ 1 Multiple time series (Ermshaus et al., 2023)
UTSA $32\times T$ 1 Multiple time series (Gharghabi et al., 2017)
Anomaly Detection FD-A 8,184 1 Mechanical System time series (Lessmeier et al., 2016)
FD-B 13,640 1 Mechanical System time series (Lessmeier et al., 2016)
KPI 5,922,913 1 Server Machine time series (Yue et al., 2022)
TODS $T$ $V$ Synthetic Data time series (Lai et al., 2021)
SMD 1,416,825 38 Server Machine time series (Su et al., 2019)
ASD 154,171 19 Server Machine time series (Li et al., 2021)
PSM 220,322 26 Server Machine time series (Abdulaal et al., 2021)
MSL 130,046 55 Spacecraft time series (Hundman et al., 2018)
SMAP 562,800 25 Spacecraft time series (Hundman et al., 2018)
SWaT 944,919 51 Infrastructure time series (Mathur and Tippenhauer, 2016)
WADI 1,221,372 103 Infrastructure time series (Ahmed et al., 2017)
Yahoo 572,966 1 Multiple time series (Yue et al., 2022)
TimeSeAD $21\times T$ $V$ Multiple time series (Wagner et al., 2023)
TSB-UAD $1,980\times T$ 1 Multiple time series (Paparrizos et al., 2022a)
UCR-TSAD $250\times T$ 1 Multiple time series (Wu and Keogh, 2021)
UCFCrime $1,900\times T$ $V_{width}\times V_{height}$ Surveillance video (Sultani et al., 2018)
Oops! $20,338\times T$ $V_{width}\times V_{height}$ Human Activity video (Epstein et al., 2020)
DFDC $128,154\times T$ $256\times 256$ Deepfake video (Dolhansky et al., 2020)
Retrieval EK-100 89,977 $V_{width}\times V_{height}$ Human Activity video (Damen et al., 2022)
HowTo100M 136,600,000 $V_{width}\times V_{height}$ Human Activity video (Miech et al., 2019)
MUSIC 714 $V_{width}\times V_{height}$ Musical Instrument multi-modality (Zhao et al., 2018)
6.3. Additional Metric for Inherent Representation Quality
Besides the task-specific metrics, recent studies (Wu et al., 2023; Dong et al., 2023) also evaluate the inherent quality of the learned representation by calculating the centered kernel alignment (CKA) similarity between representations from the first and the last layers. A higher CKA similarity indicates more similar representations. As the bottom layer representations usually contain low-level or detailed information, a smaller similarity means the top layer contains different information from the bottom layer and indicates the model tends to learn high-level representations or more abstract information.
In general, better forecasting and anomaly detection accuracy is related to higher CKA similarity, while better imputation and classification results corresponds to the lower CKA similarity. Specifically, the lower CKA similarity means that the representations are distinguishing among different layers, thus being hierarchical representations. These results also indicate the property of representations that each task requires. Thus, we can use this metric to evalaute whether an encoder learn appropriate representations for different tasks.
7. Open Challenges and Future Research Directions
In this section, we discuss open challenges around time-series analysis and outline promising future research directions that have the potential to enhance the existing universal time-series representation learning methods.
7.1. Time-Series Annotation and Active Learning
The complexity and length of time-series data significantly increase the cost of manual labeling. Obtaining annotated time series is challenging due to the domain-specific nature and lack of publicly accessible sources. Even domain experts may struggle to provide consistent annotations because time-series shapes can be perceived differently among annotators. For instance, datasets like Sleep-EDF (Goldberger et al., 2000), containing lengthy EEG recordings, require doctors to identify anomalies with a consensus among multiple experts for each annotation. This level of expertise is also demanded by applications such as anomaly detection in smart factory sensors, making the process time-consuming and labor-intensive.
Efforts within the deep learning community have aimed to address label sparsity, notably through active learning (Shin et al., 2021; Rana and Rawat, 2022). This approach minimizes labeling costs by selecting the most informative unlabeled instances and presenting them to an expert (oracle) for annotation. Active learning involves developing a query strategy that prioritizes which data points to label next based on predefined criteria, e.g., informativeness and diversity. The process iterates by training a model with the initial labels, selecting the most informative or diverse unlabeled data points for labeling, retraining the model, and repeating this cycle until a certain performance threshold (e.g., annotation budget) is reached. Therefore, developing an effective annotation process with active learning for time-series representation learning, enabling a more robust supervision signal, is expected to be a promising research direction.
7.2. Distribution Shifts and Adaptation
Distribution shifts between training and testing data are one of the factors causing the trained model to exhibit low performance during the test phase. When a time series model undergoes continuous testing, especially during the model deployment, various forms of concept drift, such as sudden, gradual, incremental, and recurrent drifts, may occur due to the accumulation of unseen time-series data (Agrahari and Singh, 2022). A domain shift problem may also arise due to differences between the source domain used in the training phase and the target domain used in the test phase.
Given that distribution shifts resulting from concept drift and domain shift are the factors that degrade model performance, previous studies have focused on concept drift adaptation and domain adaptation to address these shifts in specific downstream tasks, such as classification, anomaly detection, and forecasting (Yuan et al., 2022; Ragab et al., 2023; Jin et al., 2022). Addressing distribution shifts in the test phase is also crucial for learning representations for various downstream environments. Therefore, researchers should consider future directions to develop distribution shift adaptations for universal representation learning with discrepancy-based and adversarial methods, for example.
7.3. Reliable Data Augmentation
Determining the appropriate type and degree of data augmentation is essential, as minor changes can significantly alter the original characteristics of time series. Various techniques, including jittering, shifting, scaling, and warping, have been used for time-series representation learning. However, the reliability of these methods has not been fully explored. With the advent of new learning paradigms, such as contrastive learning, the importance of data augmentation increases. Data augmentation serves not only to expand the size of datasets but also to provide diverse class-invariant features. Unfortunately, most current approaches focus more on the evolving role of data augmentation, often overlooking its fundamental task to maintain the integrity of the original data characteristics. Recent studies (Luo et al., 2023; Demirel and Holz, 2023) have started to address the reliability of data augmentation and develop new techniques or adaptive strategies based on more reliable criteria. Despite these advancements, many studies still rely on empirical methods for determining data augmentation. Therefore, a promising direction is devising an evaluation method for time-series data augmentation to estimate the reliability and effectiveness of selecting the potential optimal augmentation strategy.
7.4. Neural Architecture Search
Making efficient deep learning models for universal representation learning is challenging because it laboriously relies on tedious manual trial-and-error to design the network architectures and select corresponding hyperparameters relevant to the three design elements we discussed. The inefficient hand-crafted design will inevitably involve human bias, leading to sub-optimal architectures (Zhang et al., 2023f). Accordingly, recent endeavors employ NAS to discover an optimal architecture by automatically designing neural architectures and network hyperparameters, e.g., the number of layers, network type, embedding dimension, and hidden dimension. These configurations significantly affect the learned representation quality and the downstream task performance. Even though NAS has demonstrated its successes in diverse tasks for computer vision and natural language processing (White et al., 2023), NAS for time series is still underexplored. There are only NAS methods for a specific time-series analysis task, such as forecasting (Shah et al., 2021; Jie and Gao, 2021; Wang et al., 2022; Deng et al., 2022; Lai et al., 2023) and classification (Rakhshani et al., 2020; Xiao et al., 2021; Ren et al., 2022), thus having very limited generalizability. Automatically discovering neural architectures for universal time-series representation learning that perform well across downstream tasks is an important future direction, especially for industry-scale time series having high dimensions and large volumes newly generated every day.
7.5. Large Language and Foundation Models
Nowadays, many large language models (LLMs) across various domains have significantly transformed the landscape of natural language processing and computer vision. Integrating LLMs into time-series representation learning will enable the model to capture rich meanings embedded in these time-dependent patterns. The synergy between linguistic context and temporal dependencies additionally allows the model to identify complex relationships within the data, enabling more robust and fine-grained representations. The applications of LLMs, such as zero-shot learning, few-shot learning, and fine-tuning, have demonstrated performance comparable to existing deep learning methods. We note ongoing attempts to leverage LLMs for time-series analysis, yet mostly limited to forecasting tasks (Jin et al., 2023b). Thus, using LLMs in time-series representation learning is expected to enhance the embedding quality by accurately capturing time-dependent patterns in the time series. We also expect future research on aligning time-series representations with language embeddings to provide valuable insights, not only in the context of single-modal time series but also in the realm of multi-modal or multivariate time series.
7.6. Representations of Irregularly-Sampled Time Series
Irregular time series naturally appear in real-world scenarios across various fields, including finance, healthcare, and environmental observation. This irregularity usually stems from sporadic observations and system failures. However, managing irregularly-sampled time series effectively has proven challenging because existing deep learning models are devised initially for regularly-spaced natural language data. Early attempts like GRU-D (Che et al., 2018) modify existing deep models to handle irregular time series. Another line of work uses Neural ODEs (Chen et al., 2018) that learn continuous vector space to represent irregular time series. While the concept of neural differential equations (NDEs) aligns well with irregular time series, these methods still have limitations. First, the training time of NDEs may be instable due to the use of the Runge-Kutta method for integration calculations. Second, the requirement that the hidden state of NDEs should be differentiable limits their ability to represent high-frequency time series. Compared to NDEs, approaches modifying attention (Shukla and Marlin, 2021) or employing interpolation (Shukla and Marlin, 2019) to address irregular time series also show promising results.
In addition to the model-centric approaches, we expect a potential research direction for irregular time series to be a data-centric approach. Unlike regular time-series datasets, irregular time series can contain diverse causes of irregularities often overlooked by previous studies. For instance, medical data may exhibit irregularities from sensor malfunctions, while financial data (e.g., credit card transactions) may contain irregularities due to its sporadic event-based nature. By integrating the understanding of the causes of irregularities into the learning process, we can obtain more precise representations of irregular time series.
7.7. Multi-Modal and Multi-View Representation Learning
Vision-language models such as CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021) have recently shown remarkable performance in zero-shot learning and fine-tuning for various vision-related downstream tasks by utilizing the semantics of human languages. If time-series data entail semantics understandable to humans, we can make multi-view representations of time series by using human languages as an additional modality to annotate them. As a result, the learned representations from time series-language representation learning will become more expressive and fine-grained in semantics. For example, human activity data (Chen et al., 2021a) or time-series anomalies can be described in human languages based on human movement and the context of abnormal situations. However, unlike video representation learning, which can leverage a vision-language pre-trained model for the annotation, time-series data annotation with human language can be laborious and highly domain-specific. Thus, building a large multi-modal time series-text dataset will be a promising future direction.
8. Conclusions
This article introduces a universal time-series representation learning research and its importance for downstream time-series analysis. We present a comprehensive and up-to-date literature review of universal representation learning for time series by categorizing the recent advancements from design perspectives. Our main goal is to answer how each fundamental design element—neural architectures, learning objectives, and training data—of state-of-the-art time-series representation learning methods contributes to the improvement of the learned representation quality, resulting in a novel structured taxonomy with fifteen subcategories. Although most state-of-the-art studies consider all design elements in their methods, only one or two elements are newly proposed. Given the current review of the selected studies, we find that decomposition and transformation methods and sample selection techniques in the data-centric approaches are still underexplored. In addition, we provide a practical guideline about standard experimental setups and widely used time-series datasets for particular downstream tasks, together with discussions on various open challenges and future research directions related to time-series representation learning. Ultimately, we expect this survey to be a valuable resource for practitioners and researchers interested in a multi-faceted understanding of the universal representation learning methods for time series.
Acknowledgements.
This work was supported by Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. 2022-0-00157, Robust, Fair, Extensible Data-Centric Continual Learning, 50% and No. 2020-0-00862, DB4DL: High-Usability and Performance In-Memory Distributed DBMS for Deep Learning, 50%).