### v3

# Universal Time-Series Representation Learning: A Survey

Abstract.
Time-series data exists in every corner of real-world systems and services, ranging from satellites in the sky to wearable devices on human bodies. Learning representations by extracting and inferring valuable information from these time series is crucial for understanding the complex dynamics of particular phenomena and enabling informed decisions. With the learned representations, we can perform numerous downstream analyses more effectively. Among several approaches, deep learning has demonstrated remarkable performance in extracting hidden patterns and features from time-series data without manual feature engineering. This survey first presents a novel taxonomy based on three fundamental elements in designing state-of-the-art universal representation learning methods for time series. According to the proposed taxonomy, we comprehensively review existing studies and discuss their intuitions and insights into how these methods enhance the quality of learned representations. Finally, as a guideline for future studies, we summarize commonly used experimental setups and datasets and discuss several promising research directions. An up-to-date corresponding resource is available at https://github.com/itouchz/awesome-deep-time-series-representations.
time series, representation learning, neural networks, temporal modeling
ccs: Computing methodologies Neural networksccs: Computing methodologies Learning latent representationsccs: Mathematics of computing Time series analysisccs: Information systems Data mining
1. Introduction
A time series is a sequence of data points recorded in chronological order, reflecting the complex dynamics of particular variables or phenomena over time. Time-series data can represent various meaningful information across application domains at different time points, enabling informed decision-making and predictions, such as sensor readings in the Internet of Things (Fathy et al., 2018; Shin et al., 2019; Trirat et al., 2023), measurements in cyber-physical systems (Giraldo et al., 2018; Luo et al., 2021), fluctuation in stock markets (Ge et al., 2022; Cao, 2022), and human activity on wearable devices (Chen et al., 2021c; Gu et al., 2021). However, to extract and understand meaningful information from such complicated observations, we need a mechanism to represent these time series, which leads to the emergence of time-series representation research. Based on the new representations, we can effectively perform various downstream time-series analyses (Esling and Agon, 2012), e.g., forecasting (Lim and Zohren, 2021), classification (Ismail Fawaz et al., 2019), regression (Tan et al., 2021), and anomaly detection (Choi et al., 2021). Fig. 1 depicts the basic concept of representation methods for time-series data.
Early attempts (Sharma et al., 2020) represent time series using piecewise linear methods (e.g., piecewise aggregate approximation), symbolic-based methods (e.g., symbolic aggregate approximation), feature-based methods (e.g., shapelets), or transformation-based methods (e.g., discrete wavelet transform). These traditional time-series representation methods are known to be time-consuming and less effective because of their dependence on domain knowledge and the poor generality of predefined priors. Since the quality of representations significantly affects the downstream task performance, many studies propose to learn the meaningful time-series representations automatically (Deldari et al., 2022). The main goal of these studies is to obtain high-quality learned representations of time series that capture valuable information within the data and unveil the underlying dynamics of the corresponding systems or phenomena. Among several approaches, neural networks or deep learning (DL) have demonstrated unprecedented performance in extracting hidden patterns and features from a wide range of data, including time series, without the requirement of manual feature engineering. Note that representation learning can be categorized into task-specific (Cheng et al., 2023) and task-agnostic(Yue et al., 2022; Liu and Chen, 2023), i.e., universal, approaches. This survey focuses exclusively on universal representation learning, which refers to methods that are designed for and evaluated on at least two downstream tasks.
Given the sequential nature of time series, recurrent neural networks (RNNs) and their variants, such as long short-term memory and gated recurrent unit, are considered a popular choice for capturing temporal dependencies in time series (Lalapura et al., 2021). Nevertheless, RNNs are complex and computationally expensive. Another line of work adopts one-dimensional convolutional neural networks (CNN) to improve the computational efficiency with the parallel processing of convolutional operations (Iwana and Uchida, 2021). Even though RNN- and CNN-based models are shown to be good at capturing temporal dependencies, they cannot explicitly model the relationship between different variables within a multivariate time series. Many studies propose to use attention-based networks or graph neural networks to jointly learn the temporal dependencies in each variable and the correlations between different variables using attention mechanisms or graph structures (Wen et al., 2023; Jin et al., 2023a). Despite the significant progress in architectural design, time series can be collected irregularly or have missing values caused by sensor malfunctions in real-world scenarios, making the commonly used neural networks inefficient due to the adversarial side-effect during the imputation process. Consequently, recent research integrates neural differential equations into existing networks such that the models can produce continuous hidden states, thereby being robust to irregular time series (Rubanova et al., 2019a; Sun et al., 2020).
Figure 1. Basic concept of time-series representation methods.
[…]
Basic concept of time-series representation methods.
The reliability and efficacy of DL-based methods are generally contingent upon the availability of sufficiently well-annotated data, commonly known as supervised learning. Time-series data, however, is naturally continuous-valued, contains high levels of noise, and has less intuitively discernible visual patterns. In contrast to human-recognizable patterns in images or texts, time series can have inconsistent semantic meanings in real-world settings across application domains. As a result, obtaining a well-annotated time series is inevitably inefficient and considerably more challenging even for domain experts due to the convoluted dynamics of the time-evolving observations collected from diverse sensors or wearable devices with different frequencies. For example, we can collect a large set of sensor signals from a smart factory, while only a few of them can be annotated by domain experts. To circumvent the laborious annotation process and reduce the reliance on labeled instances, there has been a growing interest in unsupervised and self-supervised learning using self-generated labels from various pretext tasks without relying on human annotation (Eldele et al., 2023).
While unsupervised and self-supervised representation learning share the same objective to extract latent representations from intricate raw time series without the human-annotated labels, their underlying mechanism is different. Unsupervised learning methods (Meng et al., 2023) usually adopt autoencoders and sequence-to-sequence models to learn meaningful representations using reconstruction-based learning objectives. However, accurately reconstructing the complex time-series data remains challenging, especially with high-frequency signals. On the contrary, self-supervised learning methods (Zhang et al., 2023e) leverage pretext tasks to autonomously generate labels by utilizing intrinsic information derived from the unlabeled data. Lately, pretext tasks with contrasting loss (also known as contrastive learning) have been proposed to enhance learning efficiency through discriminative pre-training with self-generated supervised signals. Contrastive learning aims to bring similar samples closer while pushing dissimilar samples apart in the feature space. These pretext tasks are self-generated challenges the model learns to solve from the unlabeled data, thereby being able to produce meaningful representations for multiple downstream tasks (Ericsson et al., 2022).
To further enhance the representation quality and alleviate the impact of limited training samples in particular settings where collecting sufficiently large data is prohibited (e.g., human-related services), several studies also employ data-related techniques, e.g., augmentation (Aboussalah et al., 2023) and transformation (Wu et al., 2023), on top of the existing learning methods. Accordingly, we can effectively increase the size and improve the quality of the training data. These techniques are also deemed essential in generating pretext tasks. Different from other data types, working with time-series data needs to consider their unique properties, such as temporal dependencies and multi-scale relationships (Wen et al., 2021).
[…]
Key design elements and evaluation protocols of a time-series representation learning framework.
Figure 2. Key design elements and evaluation protocols of a time-series representation learning framework.
1.1. Related Surveys
According to the background discussed above, there are three fundamental elements (also illustrated in Fig. 2) in designing a state-of-the-art universal representation learning method for time series: training data, network architectures, and learning objectives. To enhance the utility and quality of available training data, data-related techniques (e.g., augmentation) are employed or introduced. The neural architectures are then designed to capture underlying temporal dependencies in time series and inter-relationships between variables of multivariate time series. Last, one or multiple learning objectives (i.e., loss functions) are newly defined to learn high-quality representations. These learning objectives are sometimes called pretext tasks if pseudo labels are generated.
Despite having the three key design elements, most existing surveys for time-series representation learning review the literature exclusively on either the neural architectural aspects (Längkvist et al., 2014; Wang et al., 2024a) or the learning aspects (Zhang et al., 2023e; Ma et al., 2023; Meng et al., 2023). An early related survey article by Längkvist et al. (2014) review the DL for time series as unsupervised feature learning algorithms. The survey focuses particularly on neural architectures with a discussion on classical time-series applications. After about a decade, a few survey papers review time-series representation learning methods by focusing on learning objective aspects. For example, Zhang et al. (2023e) and Deldari et al. (2022) review self-supervised learning-based models, while, for a broader scope, Meng et al. (2023) review unsupervised learning-based methods. Similarly, Ma et al. (2023) present the survey for any learning objectives with a focus on analyzing reviewed articles from transfer learning and pre-training perspectives. For a smaller scope, a survey (Eldele et al., 2023) reviews the state-of-the-art studies that specifically tackle label scarcity in time-series data. With the advent of foundation and large language models (LLMs), recent articles (Ye et al., 2024; Liang et al., 2024) review the adaptation of these models to time series with the main focus on learning aspects. Unlike these papers, we comprehensively review the representation learning methods for time series by focusing on their universality—effectiveness across diverse downstream tasks—with discussions on their intuitions and insights into how these methods enhance the quality of learned representations from all three design aspects. Specifically, we aim to review and identify research directions on how recent state-of-the-art studies design the neural architecture, devise corresponding learning objectives, and utilize training data to enhance the quality of the learned representations of time series for downstream tasks. Table 1 summarizes the differences between our survey and the related work.
Table 1. Comparison of the survey scope between this article and related papers.
Survey Article Key Design Elements in Review Survey Coverage
Training Data Neural Architecture Learning Objective Universality Irregularity Experimental Design
Generation Augmentation Supervised Unsupervised Self-Supervised
Längkvist et al. (2014) $\checkmark$
Deldari et al. (2022) $\checkmark$ $\checkmark$
Eldele et al. (2023) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Ma et al. (2023) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Zhang et al. (2023e) $\checkmark$ $\checkmark$
Meng et al. (2023) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Ye et al. (2024) $\checkmark$ Only LLMs $\checkmark$ $\checkmark$ $\checkmark$
Liang et al. (2024) $\checkmark$ $\checkmark$ $\checkmark$
Wang et al. (2024a) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Our Survey $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
1.2. Survey Scope and Literature Collection
For the literature review, we search for papers using the following keywords and inclusion criteria.
Keywords. “time series” AND “representation”, “time series” AND “embedding”, “time series” AND “encoding”, “time series” AND “modeling”, “time series” AND “deep learning”, “temporal” AND “representation”, “sequential” AND “representation”, “audio” AND “representation”, “sequence” AND “representation”, and (“video” OR “action”) AND “representation”. We use these keywords to search well-known repositories, including ACM Digital Library, IEEE Xplore, Google Scholars, Semantic Scholars, and DBLP, for the relevant papers.
Inclusion Criteria. The initial set of papers found with the above search queries are further filtered by the following criteria. Only papers meeting the criteria are included for review.
  • Being written in the English language only
  • Being a deep learning or neural networks-based approach
  • Being published in or after 2018 at a top-tier conference or high-impact journal [^1] Top-tier venues are evaluated based on CORE (https://portal.core.edu.au), KIISE (https://www.kiise.or.kr), or Google Scholar (https://scholar.google.com). Only publications from venues rated at least A by CORE/KIISE or in the top 20 in at least one subcategory by Google Scholar metrics are included for review. Recent arXiv papers also included if their authors have publication records in the qualified venues.
  • Being evaluated on at least two downstream tasks using time-series, video, or audio datasets
Table 2. The proposed taxonomy of universal representation learning for time series.
Design Element Coarse-to-Fine Taxonomy References
Data-Centric Approaches (Training Data) Improving Quality (§$\S$3.1) Sample Selection (Chen et al., 2021a; Franceschi et al., 2019)
Time-Series Decomposition (Zeng et al., 2021; Behrmann et al., 2021; Wang et al., 2018; Yang et al., 2022a; Fang et al., 2023)
Input Space Transformation (Anand and Nayak, 2021; Biloš et al., 2022; Lee et al., 2022a; Wu et al., 2023; Zhong et al., 2023; Xu et al., 2024)
Increasing Quantity (§$\S$3.2) Augmentation Random Augmentation (Yue et al., 2022; Zhang et al., 2022b, 2023f; Liu and Chen, 2023)
Policy-Based Augmentation (Kim et al., 2023c; Chen et al., 2022; Kim et al., 2023b; Zhang and Crandall, 2022; Aboussalah et al., 2023; Luo et al., 2023; Yang et al., 2022b; Yang and Hong, 2022; Shin et al., 2023; Demirel and Holz, 2023; Zheng et al., 2024; Duan et al., 2024; Li et al., 2024)
Generation and Curation (Nguyen et al., 2023; Zhao et al., 2023; Liu et al., 2024; Goswami et al., 2024)
Neural Architectural Approaches (Network Architecture) Task-Adaptive Submodules (§$\S$4.1) (Chen et al., 2021a; Wang et al., 2023c; Gorbett et al., 2023; Liang et al., 2023a; Gao et al., 2024; Zhang et al., 2024; Goswami et al., 2024)
General Temporal Modeling (§$\S$4.2) Intra-Variable Modeling Long-Term Modeling (Zerveas et al., 2021; Tonekaboni et al., 2022; Zhang et al., 2023c; Nguyen et al., 2023; Luo and Wang, 2024; Han et al., 2020; Franceschi et al., 2019)
Multi-Resolution Modeling (Liu et al., 2022; Zhong et al., 2023; Fraikin et al., 2023; Nguyen et al., 2023; Liu et al., 2020; Sanchez et al., 2019; Wang et al., 2023a; Sener et al., 2020; Wang et al., 2024c; Eldele et al., 2024)
Inter-Variable Modeling (Zhong et al., 2023; Xie et al., 2022; Guo et al., 2021; Cai et al., 2021; Xiao et al., 2023; Wang et al., 2023b, 2024c; Luo and Wang, 2024; Zhang et al., 2024; Wang et al., 2024b)
Frequency-Aware Modeling (Yang and Hong, 2022; Wang et al., 2023c; Wu et al., 2023; Xu et al., 2024; Wang et al., 2018; Zhou et al., 2023b; Eldele et al., 2024; Duan et al., 2024)
Auxiliary Feature Extraction (§$\S$4.3) Shapelet and Motif Feature Modeling (Liang et al., 2023b; Qu et al., 2024)
Contextual Modeling and LLM Alignment (Chen et al., 2019; Anand and Nayak, 2021; Kim et al., 2023a; Choi and Kang, 2023; Rahman et al., 2021; Zhou et al., 2023a; Lin et al., 2024; Zhou et al., 2023b; Liu et al., 2019; Li et al., 2022; Lee et al., 2022b; Li et al., 2024; Luetto et al., 2023; Chowdhury et al., 2022)
Continuous Temporal and Irregular Time-Series Modeling (§$\S$4.4) Neural Differential Equations (Jhin et al., 2021; Chen et al., 2023; Abushaqra et al., 2022; Jhin et al., 2022; Rubanova et al., 2019b; Oh et al., 2024)
Implicit Neural Representations (Naour et al., 2023; Fons et al., 2022)
Auxiliary Networks (Ma et al., 2019; Zhang et al., 2023a; Bianchi et al., 2019; Schirmer et al., 2022; Shukla and Marlin, 2021; Ansari et al., 2023; Sun et al., 2021; Senane et al., 2024)
Learning-Focused Approaches (Learning Objective) Task-Adaptive Losses (§$\S$5.1) (Ma et al., 2019; Rahman et al., 2021; Hadji et al., 2021; Zhong et al., 2023; Wang et al., 2024c; Gao et al., 2024)
Non-Contrasting Losses for Temporal Dependency Learning (§$\S$5.2) Input Reconstruction Learning (Chen et al., 2019; Sanchez et al., 2019; Nguyen et al., 2018; Li et al., 2023; Yuan et al., 2019; Chen et al., 2024; Lee et al., 2024a)
Masked Prediction Learning (Zerveas et al., 2021; Biloš et al., 2023; Dong et al., 2023; Chowdhury et al., 2022; Zhang et al., 2024; Senane et al., 2024)
Customized Pretext Tasks (Zhang et al., 2022a; Guo et al., 2022; Haresh et al., 2021; Wu et al., 2018; Liang et al., 2022; Wang et al., 2020, 2021; Nguyen et al., 2018; Duan et al., 2022; Lee et al., 2022b; Fang et al., 2023; Fraikin et al., 2023; Sun et al., 2023; Dong et al., 2024; Liu et al., 2024; Bian et al., 2024)
Contrasting Losses for Consistency Modeling (§$\S$5.3) Subseries Consistency (Behrmann et al., 2021; Qian et al., 2022; Wang et al., 2020; Qian et al., 2021; Somaiya et al., 2022; Franceschi et al., 2019)
Temporal Consistency (Morgado et al., 2020; Wang et al., 2024b; Chen et al., 2022; Zhang et al., 2023b; Haresh et al., 2021; Eldele et al., 2021; Yang et al., 2022b; Tonekaboni et al., 2021; Hajimoradlou et al., 2022; Yang et al., 2023a)
Contextual Consistency (Yue et al., 2022; Zhang et al., 2023f; Zhong et al., 2023; Xu et al., 2023; Chen et al., 2022; Shin et al., 2023; Jiao et al., 2020; Chowdhury et al., 2023)
Transformation Consistency (Jenni and Jin, 2021; Eldele et al., 2021; Chen et al., 2021b)
Hierarchical and Cross-Scale Consistency (Liang et al., 2023b; Nguyen et al., 2023; Kong et al., 2020; Qing et al., 2022; Wang et al., 2023a; Zhang and Crandall, 2022; Liu et al., 2022; Dave et al., 2022; Duan et al., 2024)
Cross-Domain and Multi-Task Consistency (Zeng et al., 2021; Ding et al., 2022; Kim et al., 2023a; Han et al., 2020; Choi and Kang, 2023; Zhang et al., 2022b; Liu et al., 2023; Liu and Chen, 2023; Lee et al., 2024b; Li et al., 2024)
(a) Focus Design Element.
(b) Publication Venue.
(c) Publication Year.
Figure 3. Quantitative summary of the selected papers.
[…]
Quantitative summary of the selected papers.
Quantitative Summary. Given the above keywords and inclusion criteria, we selected 127 papers in total. Fig. 3 shows the quantitative summary of the paper selected for review. We can notice from Fig. 3(a) that neural architectures and learning objectives are similarly considered important in designing state-of-the-art methods. Most papers were published at ICLR, NeurIPS, followed by AAAI and ICML (Fig. 3(b)). According to Fig. 3(c), we expect more papers on this topic will be published in the future.
1.3. Contributions
This paper aims to identify what are the essential elements in designing state-of-the-art representation learning methods for time series and how these elements affect the quality of the learned representations. To the best of our knowledge, this is the first survey on universal time-series representation learning. We propose a novel taxonomy for learning universal representations of time series from the novelty perspectives—whether the main contribution of a paper focuses on what part of the design elements discussed above—to summarize the selected studies. Table 2 outlines and compares the reviewed articles based on the proposed taxonomy. We begin by exploring the essence of data-driven methods, categorizing data-centric approaches from the papers that primarily aim to enhance the usefulness of the training data. From the perspective of network architectures, we review the evolution of neural architectures by focusing on how these advancements enhance the quality of learned representations. From the learning perspective, we classify how different objectives enhance the generalizability of learned representations across diverse downstream tasks. Overall, our main contributions are as follows.
  • We conduct an extensive literature review of universal time-series representation learning based on a novel and up-to-date taxonomy that categorizes the reviewed methods into three main categories: data-centric, neural architectural, and learning-focused approaches.
  • We provide a guideline on the experimental setup and benchmark datasets for assessing representation learning methods for time series.
  • We discuss several open research challenges and new insights to facilitate future work.
Article Organization. Section 2 introduces the definitions and specific background knowledge regarding time-series representation learning. Section 3 gives a review on data-centric approaches. In Section 4, we present an extensive review of the methods focusing on neural architecture aspects. We then discuss the methods focusing on deriving new learning objectives in Section 5. In addition, we discuss the evaluation protocol for time-series representation learning and promising future research directions in Section 6 and Section 7, respectively. Finally, Section 8 concludes this survey.
Figure 4. Illustrations of regularly and irregularly sampled multivariate time series ($V=3$).
[…]
Illustrations of regularly and irregularly sampled multivariate time series with V = 3.
2. Preliminaries
This section presents definitions and notations used throughout this paper, descriptions of downstream time-series analysis tasks, and unique properties of time series.
2.1. Definitions
Definition 2.1 (Time Series).
A time series $\mathbf{X}$ is an chronologically ordered sequence of $V$-variate data points recorded at specific time intervals. $\mathbf{X}=(\mathbf{x}_{1},\dots,\mathbf{x}_{t},\mathbf{x}_{T})$, where $\mathbf{x}_{t}\in\mathbb{R}^{V}$ is the observed value at $t$-th time step, $V$ is the number of variables, and $T$ is the length of the time series. When $V=1$, it becomes the univariate time series; otherwise, it is multivariate time series. Audio and video data can be considered special cases of time series with more dimensions. The time intervals are typically equally spaced, and the values can represent any measurable quantity, such as temperature, sales figures, or any phenomenon that changes over time.
Definition 2.2 (Irregularly-Sampled Time Series).
An irregularly-sampled time series (ISTS) is a time series where the intervals between observations are not consistent or regularly spaced. Thus, the time intervals between $(\mathbf{x}_{1},\mathbf{x}_{2})$ and $(\mathbf{x}_{2},\mathbf{x}_{3})$ is unequal, as illustrated in Fig. 4. It is often encountered in situations where data is collected opportunistically or when events occur irregularly and sporadically, e.g., sensor malfunctions, leading to varying time gaps between observations.
Definition 2.3 (Time-Series Representation Learning).
Given a raw time series $\mathbf{X}$, time-series representation learning aims to learn an encoder $f_{e}$, a nonlinear embedding function that maps $\mathbf{X}$ into a $R$-dimensional representation vector $\mathbf{Z}=(\mathbf{z}_{1},\dots,\mathbf{z}_{R-1},\mathbf{z}_{R})$ in the latent space, where $\mathbf{z}_{i}\in\mathbb{R}^{F}$. $\mathbf{Z}$ usually has either equal ($R=T$) or shorter ($R<T$) length of the original time series. When $R=T$, $\mathbf{Z}$ is timestamp-wise (or point-wise) representation that contains representation vectors $\mathbf{z}_{i}$ with feature size $F$ for each $t$. In contrast, when $R<T$, $\mathbf{Z}$ is the compressed version of $\mathbf{X}$ with reduced dimension, and $F$ is usually $1$, producing the series-wise (or instance-wise) representation.
A crucial measure to assess the quality of a representation learning method, i.e., the encoder $f_{e}$, is its ability to produce the representations $\mathbf{Z}$ that effectively facilitate downstream tasks, either with or without fine-tuning (see Section 6). Once we obtain $\mathbf{Z}$, we can use it as input for downstream tasks to evaluate its actual performance. Here, we define the common downstream tasks as follows.
Definition 2.4 (Forecasting).
Time-series forecasting (TSF) aims to predict the future values of a time series by explicitly modeling the dynamics and dependencies among historical observations. It can be short-term or long-term forecasting depending on the prediction horizon $H$. Formally, given a time series $\mathbf{X}$, TSF predicts the next $H$ values $(\mathbf{x}_{T+1},\dots,\mathbf{x}_{T+H})$ that are most likely to occur.
Definition 2.5 (Classification).
Time-series classification (TSC) aims to assign predefined class labels $\mathbf{C}=\{c_{1},\dots,c_{|\mathbf{C}|}\}$ to a set of time series. Let $\mathcal{D}=\{(\mathbf{X}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$ denote a time-series dataset with $N$ samples, where $\mathbf{X}_{i}\in\mathbb{R}^{T\times V}$ is a time series and $\mathbf{y}_{i}$ is the corresponding one-hot label vector of length $|\mathbf{C}|$. The $j$-th element of the one-hot vector $\mathbf{y}_{i}$ is equal to 1 if the class of $\mathbf{X}_{i}$ is $j$, otherwise 00. Formally, TSC trains a classifier on the given dataset $\mathcal{D}$ by learning the discriminative features to distinguish different classes from each other. Then, when an unseen dataset $\mathcal{D}^{\prime}$ is input to the trained classifier, it automatically determines to which class $c_{i}$ each time series belongs.
Definition 2.6 (Extrinsic Regression).
Time-series extrinsic regression (TSER) shares a similar goal to TSC with a key difference in label annotation. While TSC predicts a categorical value, TSER predicts a continuous value for a variable external to the input time series. That is, $y_{i}\in\mathbb{R}$. Formally, TSER trains a regression model to map a given time series $\mathbf{X}_{i}$ to a numerical value $y_{i}$.
Definition 2.7 (Clustering).
Time-series clustering (TSCL) is the process of finding natural groups, called clusters, in a set of time series $\mathcal{X}=\{\mathbf{X}_{i}\}_{i=1}^{N}$. It aims to partition $\mathcal{X}$ into a group of clusters $\mathbf{G}=\{g_{1},\dots,g_{i},g_{|\mathbf{G}|}\}$ by maximizing the similarities between time series within the same cluster and the dissimilarities between time series of different clusters. Formally, given a similarity measure $f_{s}(\cdot,\cdot)$, $\forall i_{1},i_{2},j~{}f_{s}(\mathbf{X}_{i_{1}},\mathbf{X}_{j})\gg f_{s}(\mathbf{X}_{i_{1}},\mathbf{X}_{i_{2}})~{}\text{for}~{}\mathbf{X}_{i_{1}},\mathbf{X}_{i_{2}}\in g_{i}$ and $\mathbf{X}_{j}\in g_{j}$.
Definition 2.8 (Segmentation).
Time-series segmentation (TSS) aims to assign a label to a subsequence $\mathbf{X}_{T_{s},T_{e}}$ of $\mathbf{X}$, where $T_{s}$ is the start offset and $T_{e}$ is the end offset, consisting of contiguous observations of $\mathbf{X}$ from time step $T_{s}$ to $T_{e}$. That is, $\mathbf{X}_{T_{s},T_{e}}=(\mathbf{x}_{T_{s}},\dots,\mathbf{x}_{T_{e}})$ and $1\leq T_{s}\leq T_{e}\leq T$. Let a change point (CP) be an offset $i\in[i,\dots,T]$ w.r.t. to a state transition in time series, TSS finds a set of segmentation of $\mathbf{X}$, having the ordered sequence of CPs in $\mathbf{X}$ (i.e., $\mathbf{x}_{i_{1}},\dots,\mathbf{x}_{i_{S}}$) with $1<i_{1}<\cdots<i_{S}<T$ at which the state of observations changed. After identifying the number and locations of all CPs, we can set the start offset $T_{e}$ and end offset $T_{e}$ for each segment in $\mathbf{X}$.
Definition 2.9 (Anomaly Detection).
Time-series anomaly detection (TSAD) aims to identify abnormal time points that significantly deviate from the other observations in a time series. Commonly, TSAD learns the representations of normal behavior from a time series $\mathbf{X}$. Then, the trained model computes anomaly scores $\mathbf{A}=(a_{i},\dots,a_{|\mathbf{X}^{\prime}|})$ of all values in an unseen time series $\mathbf{X}^{\prime}$ to determine which time point in $\mathbf{X}^{\prime}$ is anomalous. The final decisions are obtained by comparing each $a_{i}$ with a predefined threshold $\delta$: anomalous if $a_{i}>\delta$ and normal otherwise.
Definition 2.10 (Imputation of Missing Values).
Time-series imputation (TSI) aims to fill missing values in a time series with realistic values to facilitate subsequent analysis. Given a time series $\mathbf{X}$ and a binary $M=(m_{1},\dots,m_{t},m_{T})$, $\mathbf{x}_{t}$ is missing if $m_{t}=0$, and is observed otherwise. Let $\mathbf{\hat{X}}$ denote the predicted values generated by a TSI method, the imputed time series $\mathbf{X}_{\text{imputed}}=\mathbf{X}\odot M+\mathbf{\hat{X}}\odot(1-M)$.
Definition 2.11 (Retrieval).
Time-series retrieval (TSR) aims to obtain a set of time series that are most similar to a query provided. Given a query time series $\mathbf{X}_{q}$ and a similarity measure $f_{s}(\cdot,\cdot)$, find an ordered list $\mathcal{Q}=\{\mathbf{X}_{i}\}_{i=1}^{K}$ of time series in the given dataset or database, containing $K$ time series that are the most similar to $\mathbf{X}_{q}$.
Following Definition 2.3, we can use the corresponding representation $\mathbf{Z}=f_{e}(\mathbf{X})$ to perform the above downstream tasks instead of the raw time series $\mathbf{X}$.
2.2. Unique Properties of Time Series
In this subsection, we discuss unique properties in time-series data that have been explored by existing studies for time-series representation learning (Längkvist et al., 2014; Ma et al., 2023). Due to the following properties, techniques devised for image or text data are usually difficult to transfer directly to time series.
2.2.1. Temporal Dependency
Time series exhibits dependencies on the time variable, where a data point at a given time correlates with its previous values. Given an input $\mathbf{x}_{t}$ at time $t$, the model predicts $y_{t}$, but the same input at a later time could be a different prediction. Therefore, windows or subsequences of past observations are usually included as inputs to the model to learn such temporal dependency. The length of windows for capturing the time dependencies could also be unknown. There are local and global temporal dependencies. The former is usually associated with abrupt changes or noises, while the latter is associated with collective trends or recurrent patterns.
2.2.2. High Noise and Dimension
Time-series data, especially in real-world environments, often contain noises and have high dimensions. These noises usually arise from measurement errors or other sources of uncertainty. Dimensionality reduction techniques and wavelet transforms can address this issue by filtering some noises and reducing the dimension of the raw time series. Nevertheless, we may lose valuable information and need domain-specific knowledge to select the suitable dimensionality reduction and filtering techniques.
2.2.3. Inter-Relationship across Variables
This characteristic is particularly notable in multivariate time series. It is often uncertain whether there is sufficient information to understand a phenomenon of a time series when only a limited number of variables are analyzed as there may exist relationships between variables underlying the process or state. For example, in electronic nose data, where an array of sensors with various selectivity for several gases are combined to determine a specific smell, there is no guarantee that the selection of sensors can identify the target odor. Similarly, in financial data, monitoring a single stock representing a fraction of a complex system may not provide enough information to forecast future values.
2.2.4. Variability and Nonstationarity
Time-series data also possess variability and nonstationarity properties, meaning that statistical characteristics, such as mean, variance, and frequency, change over time. These changes usually reveal seasonal patterns, trends, and fluctuations. Here, seasonality refers to repeating patterns that regularly appear, while trends describe long-term changes or shifts over time. In some cases, the change in frequency is so relevant to the task that it is more beneficial to work in the frequency domain than in the time domain.
2.2.5. Diverse Semantics
In contrast to image and text data, learning universal representations of time series is challenging due to the lack of a large-scale unified semantic time-series dataset. For instance, each word in a text dataset has similar semantics in different sentences with high probability. Accordingly, the word embeddings learned by the model can transfer across different scenarios. However, time-series datasets are challenging to obtain subsequences (corresponding to words in text sequences) that have consistent semantics across scenarios and applications, making it difficult to transfer the knowledge learned by the model. This property also makes time-series annotation tricky and challenging, even for domain experts.
2.3. Neural Architectures for Time Series
Choosing an appropriate neural architecture to model complex temporal and inter-relationship dependencies in time series is an essential part of the design elements. This subsection overviews basic neural network architectures used as building blocks in state-of-the-art representation learning methods for time series.
2.3.1. Multi-Layer Perceptrons (MLP)
The most basic neural network architecture is the MLP (Ismail Fawaz et al., 2019), i.e., fully connected (FC) network. In MLP models, the number of layers and neurons (or hidden units) are hyperparameters to be defined. Specifically, all neurons in a layer are connected to all neurons of the following layer. These connections contain weights in the neural network. The weights are later updated by applying a non-linearity to an input. As MLP-based models process input data in a single fixed-length representation without considering the temporal relationships between the data points, they are basically unsuitable for capturing the temporal dependencies and time-invariant features. Each time step is weighted individually, and time-series elements are learned independently from each other.
2.3.2. Recurrent Neural Networks (RNN)
RNN (Mao and Sejdić, 2022) is a neural architecture with internal memory (i.e., state) specifically designed to process sequential data, thus suitable for learning temporal features in time series. The memory component enables RNN-based models to refer to past observations when processing the current one, resulting in improved learning capability. RNNs can also process variable-length inputs and produce variable-length outputs. This capability is enabled by sharing parameters over time through direct connections between layers. However, they are ineffective in modeling long-term dependencies, besides being computationally expensive due to sequential processing. RNN-based models are usually trained iteratively using a technique called back-propagation through time. Unfolded RNNs are similar to deep networks with shared parameters connected to each RNN cell. Due to the depth and weight-sharing in RNNs, the gradients are summed up at each time step to train the model, undergoing continuous matrix multiplication due to the chain rule. Thus, the gradients often either shrink exponentially to small values (i.e., vanishing gradients) or blow up to large values (i.e., exploding gradients). These problems give rise to long short-term memory (LSTM) and gated recurrent unit (GRU).
Long Short-Term Memory (LSTM)
LSTM (Hochreiter and Schmidhuber, 1997) addresses the well-known vanishing and exploding gradient problems in the standard RNNs by integrating memory cells with a gating mechanism (i.e., cell state gate, input gate, output gate, and forget gate) into their state dynamics to control the information flow between cells. As the inherent nature of LSTM is the same as RNN, LSTM-based models are also suitable for learning sequence data like time series and video representation learning, with a better capability to capture long-term dependencies.
Gated Recurrent Unit (GRU)
GRU (Cho et al., 2014) is another popular RNN variant that can control information flow and memorize states across multiple time steps, similar to LSTM, but with a simpler cell architecture. Compared to LSTM, GRU cells have only two gates (reset and update gates), making it more computationally efficient and requiring less data to generalize.
2.3.3. Convolutional Neural Networks (CNN)
CNN (Ismail Fawaz et al., 2019) is a very successful neural architecture, proven in many domains, including computer vision, speech recognition, and natural language processing. Accordingly, researchers also adopt CNN for time series. To use CNN for time series, we need to encode the input data into an image-like format. The CNN receives embedding of the value at each time step and then aggregates local information from nearby time steps using convolution layers. The convolution layer, consisting of several convolution kernels (i.e., filters), aims to learn feature representations of the inputs by computing different feature maps. Each neuron of a feature map connects to a region of neighboring neurons in the previous layer called the receptive field. The feature maps can be created by convolving the inputs with learned kernels and applying an element-wise nonlinear activation to the convolved results. Here, all spatial locations of the input share the kernel for each feature map, and several kernels are used to obtain the entire feature map. Many improvements have been made to CNN, such as using deeper networks, applying smaller and more efficient convolutional filters, adding pooling layers to reduce the resolution of the feature maps, and utilizing batch normalization to improve the training stability. As standard CNNs are designed for processing images, widely used CNN architectures for time series are one-dimensional CNN and temporal convolutional networks (Eldele et al., 2023).
Temporal Convolutional Networks (TCN)
Different from the standard CNNs, TCN (Bai et al., 2018) uses the fully convolutional network to make all layers the same length and employ casual convolution operation to avoid information leakage from the future time step to the past. Compared to RNN-based models, TCN has recently shown to be more accurate, simpler, and more efficient across diverse downstream tasks (Ma et al., 2023).
2.3.4. Graph Neural Networks (GNN)
GNN (Jin et al., 2023a) aims to learn directly from graph representations of data. A graph consists of nodes and edges, with each edge connecting two nodes. Both nodes and edges can have associated features. Edges can be directional or un-directional and can be weighted. Graphs better represent data not easily represented in Euclidean space, such as spatio-temporal data like the electroencephalogram and traffic monitoring networks. GNNs receive the graph structure and any associated node and edge attributes as input. Typically, the core operation in GNNs is graph convolution, which involves exchanging information across neighboring nodes. This operation enables the GNN-based model to explicitly rely on the inter-variable dependencies represented by the graph edges for processing multivariate time-series data. While both RNNs and CNNs perform well on Euclidean data, time series are often more naturally represented as graphs in many scenarios. Consider a network of traffic sensors where the sensors are not uniformly spaced, deviating from a regular grid. Representing this data as a graph captures its irregularity more precisely than a Euclidean space. However, using standard deep learning algorithms to learn from graph structures is challenging as nodes may have varying numbers of neighboring nodes, making it difficult to apply the convolution operation. Thus, GNNs are more suitable to graph data.
2.3.5. Attention-based Networks
The attention mechanism was introduced by Bahdanau et al. (2015) to improve the performance of encoder-decoder models in machine translation. The encoder encodes a source sentence into a vector in latent space, and the decoder then decodes the latent vector into a target language sentence. The attention mechanism enables the decoder to pay attention to the segments of the source for each target through a context vector. Attention-based neural networks are designed to capture long-range dependencies with broader receptive fields, usually lacking in CNNs and RNNs. Thus, attention-based models provide more contextual information to enhance the models’ learning and representation capability. The underlying mechanism (i.e., attention mechanism) is proposed to make the model focus on essential features in the input while suppressing the unnecessary ones. For instance, it can be used to enhance LSTM performance in many applications by assigning attention scores for LSTM hidden states to determine the importance of each state in the final prediction. Moreover, the attention mechanism can improve the interpretability of the model. However, it can be more computationally expensive due to the large number of parameters, making it prone to overfitting when the training data is limited.
Self-Attention Module
Self-attention has been demonstrated to be effective in various natural language processing tasks due to its ability to capture long-term dependencies in text (Foumani et al., 2023). The self-attention module is usually embedded in encoder-decoder models to improve the model performance and leveraged in many studies to replace the RNN-based models to improve learning efficiency due to its fast parallel processing.
Transformers
The unprecedented performance of stacking multi-headed attention, called Transformers (Vaswani et al., 2017), has led to numerous endeavors to adapt multi-headed attention to time-series data. Transformers for time series usually contain a simple encoder structure consisting of multi-headed self-attention and feed-forward layers. They integrate information from data points in the time series by dynamically computing the associations between representations with self-attention. Thanks to the practical capability to model long-range dependencies, Transformers have shown remarkable performance in sequence data. Many recent studies use Transformers as the backbone architecture for time-series analysis. (Wen et al., 2023).
2.3.6. Neural Differential Equations (NDE)
Let $f_{\theta}$ be a function specifying continuous dynamics of a hidden state $\mathbf{h}(t)$ with paramters $\theta$. Neural ordinary differential equations (Neural ODE) are one of the continuous-time models that define the hidden state $\mathbf{h}(t)$ as a solution to ODE initial-value problem $\frac{d\mathbf{h}(t)}{dt}=f_{\theta}(\mathbf{h}(t),t)$ where $\mathbf{h}(t_{o})=\mathbf{h}_{0}$. The hidden state $\mathbf{h}(t)$ is defined at all time steps, and can be evaluated at any desired time steps using a numerical ODE solver. Formally, $\mathbf{h}_{0},\dots,\mathbf{h}_{T}=\text{ODESolver}(f_{\theta},\mathbf{h}_{0},(t_{0},\dots,t_{T}))$. For training ODE-based deep learning models using black-box ODE solvers, we can use the adjoint sensitivity method to compute memory-efficient gradients w.r.t. the neural network parameters $\theta$, as described by Rubanova et al. (2019a). Neural ODEs are usually combined with RNN or its variants to sequentially update the hidden state at observation times (Shukla and Marlin, 2020). These models provide an alternative recurrence-based solution with better properties than traditional RNNs in terms of their ability to handle irregularly sampled time series.
Figure 5. Illustrative examples of data-centric approaches.
[…]
Illustrative examples of data-centric approaches. (a) Improving quality focuses on selecting the most useful samples or extracting underlying properties from available training data, while (b) increasing data quantity aim to increase the size and diversity of the data.
3. Data-Centric Approaches
This section presents the methods that focus on finding a new way to enhance the usefulness of the training data at hand. These approaches prioritize engineering the data itself rather than focusing on model architecture and loss function design to capture the underlying patterns, trends, and relevant features within the time series. As in Fig. 5, we categorize these data-centric approaches into two groups based on their objectives: improving data quality or increasing data quantity.
3.1. Improving Data Quality
3.1.1. Sample Selection
Sample selection aims to effectively choose the best samples from available training data for a particular learning scenario. An early study adopting sample selection for time-series representation learning is proposed by Franceschi et al. (2019). This work uses time-based negative sampling which determines several negative samples by independently choosing sub-series from other time series at random, whereas a sub-series within the referenced time series is consider a positive sample. This technique encourages representations of the input time segment and its sampled sub-series to be close to each other. MTRL (Chen et al., 2021a) utilizes discriminative samples to design a distance-weighted sampling strategy for achieving high convergence speed and accuracy.
3.1.2. Time-Series Decomposition
mWDN (Wang et al., 2018) is an early attempt that integrates multi-level discrete wavelet decomposition into a deep neural network framework. This framework allows for the preservation of frequency learning advantages while facilitating the fine-tuning of parameters within a deep learning context. Fang et al. (2023) decompose the spatial relation within multivariate time series into prior and dynamic graphs to model both common relation shared across all instances and distinct correlation that varies across instances.
To enhance video-level tasks, including video classification, and more detailed tasks like action segmentation, Behrmann et al. (2021) separate the representation space into stationary and non-stationary characteristics through contrastive learning from long and short views. Zeng et al. (2021) improve generalization across various downstream tasks by learning spatially-local/temporally-global and spatially-global/temporally-local features from audio-visual modalities to capture global and local information inside a video. This method enables the capturing of both slowly changing patch-level information and fast changing frame-level information. More recently, Yang et al. (2022a) propose a unified framework for joint audio-visual speech recognition and synthesis. Each modality is decomposed into modality-specific and modality-invariant features. Modality-invariant codebook enhances the alignment between the linguistic representation space of visual and audio modalities.
3.1.3. Input Space Transformation
To effectively exploit vision models, DeLTa (Anand and Nayak, 2021) transforms 1D time series into 2D images and use the transformed images as features for the subsequent learning phase with models pretrained on large image datasets, such as ResNet50. Similarly, TimesNet (Wu et al., 2023) learns the temporal variations in the 2D space to enhance the representation capability by capturing variations both within individual periods and across multiple periods in the time series. It analyzes the time series from a new dimension of multi-periodicity by extending the analysis of temporal variations from 1D into 2D space. By transforming the 1D time series into a set of 2D tensors, TimesNet breaks the bottleneck of representation capability in the original 1D space, enabling well-performing vision backbones applicable to the transformed time series. Recently, Zhong et al. (2023) propose a novel multi-scale temporal patching approach to divide the input time series into non-overlapping patches along the temporal dimension in each layer. It treats the time series as patches. FITS (Xu et al., 2024) conducts interpolation in the frequency domain to extend time series and incorporates a low-pass filter to ensure a compact representation while preserving essential information.
To address tasks involving irregular time series, some input space transformation techniques have been proposed. SplineNet (Biloš et al., 2022) generates splines from the input time series and directly utilizes them as input for a neural network. This approach introduces a learnable spline kernel to effectively process the input spline. MIAM (Lee et al., 2022a) considers multiple views of the input data, including time intervals, missing data indicators, and observation values. These transformed input data are then processed by the multi-view integration attention module to solve downstream tasks.
3.2. Increasing Data Quantity
3.2.1. Data Augmentation
Unlike other data types, augmenting time-series data requires special attention to its unique characteristics, such as temporal dependencies, multi-scale patterns, and inter-variable relationships. We categorize each method as either random or policy-based augmentation.
Random Augmentation
TS2Vec (Yue et al., 2022) randomly treats two overlapping time segments as positive pairs for contrastive learning. Specifically, timestamp masking and random cropping are applied to the input time series to generate a context. Here, the contrasting objective in TS2Vec is based on the augmented context views, meaning that representations of the same sub-series in two augmented contexts should be consistent. As a result, we can obtain a robust contextual representation for each sub-series without introducing unappreciated inductive biases like transformation- and cropping-invariance, which proven effective for a wide range of downstream tasks. By considering frequency information, TF-C (Zhang et al., 2022b) aims to use both time and frequency features of time series to improve the representation quality. It exploits spectral information and explores time-frequency consistency in time series with a set of novel augmentations based on the characteristic of the frequency spectrum. Specifically, TF-C introduces frequency-domain augmentation by randomly adding or removing frequency components, thereby exposing the model to a range of frequency variations. TimesURL (Liu and Chen, 2023) uses a frequency-temporal-based augmentation that generates augmented contexts by frequency mixing and random cropping to maintain the temporal property unchanged. Recently, TS-CoT (Zhang et al., 2023f) creates diverse views for contrastive learning by enhancing robustness to noisy time series that contributes to the overall effectiveness of the representation learning.
Policy-based Augmentation
Instead of relying on randomness, several studies devise a specific criterion for data augmentation. TimeCLR (Yang et al., 2022b) introduces a new data augmentation method based on dynamic time warping (DTW) that induces phase shifts and amplitude changes in time-series data while preserving its structural and feature information. Similarly, Yang et al. (2023a) introduce TempCLR, a contrastive learning framework for exploring temporal dynamics in video-paragraph alignment, presenting a new negative sampling strategy based on temporal granularity. TempCLR focuses on sequence-level comparison, using DTW to model the temporal order and dynamics of the data. BTSF (Yang and Hong, 2022) uses the entire time series as input and applies a standard dropout as an instance-level augmentation to generate different views for representation learning of time series. This paper argues that using dropout as the augmentation method can preserve the global temporal information and capture long-term dependencies of time series more effectively. The construction of contrastive pairs also ensures that the augmented time series do not change their raw properties, effectively reducing the potential false negatives and positives. Based on consistency regularization, Shin et al. (2023) propose a novel method of data augmentation called context-attached augmentation, which adds preceding and succeeding instances to a target instance to form its augmented instance, to fully exploit the sequential characteristic of time series in consistency regularization. This context-attached augmentation generates instances augmented with varying contexts while maintaining the target instance, thus avoiding direct perturbations on the target instance’s attributes. RIM (Aboussalah et al., 2023) is another new time-series augmentation method that generates more samples by using a recursive interpolation function from the original time series to be used in training. This method can control how much the augmented time series trajectory deviates from the original trajectory.
Using frequency-domain information, Demirel and Holz (2023) propose a novel data augmentation technique. This method is based on a mixup technique for non-stationary quasi-periodic time series that aims to connect intra-class samples together to find order in the latent space. With a mixture of the magnitude and phase of frequency components together with tailored operations, it can prevent destructive interference of mixup. MF-CLR (Duan et al., 2024) is a multi-frequency based method designed for self-supervised learning on multi-dimensional time series with varying frequencies. It uses a new data augmentation method called Dual Twister to enhance learning across different frequency levels by adding noise along both dimensions of an input time series to generate positive views with similar semantic meanings while preserving the original frequency structure.
Given the challenge of selecting meaningful augmentations on the fly, Luo et al. (2023) propose an adaptive data augmentation method, InfoTS, that uses information-aware criteria for selecting the optimal data augmentations to generate feasible positive samples for contrastive learning. The positive samples generated by InfoTS exhibit both high fidelity and diversity, which are essential for contrastive learning. Likewise, AutoTCL (Zheng et al., 2024) adaptively augments data based on a proposed input factorization technique to avoid ad-hoc decisions and reduce trial-and-error tuning by using a neural network to separate an instance into informative and task-irrelevant parts. The informative part is then transformed to preserve semantics, creating positive samples for learning. The method increases diversity in the augmented data and optimizes the augmentation process alongside contrastive learning. Li et al. (2024) introduce a novel approach to address the limitations of existing time-series models, which often suffer from high bias and low generalization due to predefined and rigid augmentation methods. The authors propose a unified and trainable augmentation operation that preserves patterns and minimizes bias by leveraging spectral information, which is designed to handle cross-domain datasets, making it highly scalable and generalizable.
For video data, Chen et al. (2022) use overlap augmentation techniques by sampling two subsequences with an overlap for each video. Overlapped timestamps are positives, and clips from other videos are negatives. Two timestamps neighboring each other also become positive pairs with Gaussian weight proportional to temporal distance. Zhang and Crandall (2022) present a novel technique that enhances the understanding of both spatial and temporal features. They achieve this by using temporal and spatial augmentations as a form of regularization, guiding the network to learn the desired semantics within a contrastive learning framework. Kim et al. (2023c) propose DynaAugment by dynamically changing the augmentation magnitude over time in order to learn the temporal variations of a video. By Fourier sampling the magnitudes, it regulates the augmentations diversely while maintaining temporal consistency. FreqAug (Kim et al., 2023b) pushes the model to focus more on dynamic features via dropping spatial or temporal low-frequency components. For various downstream tasks, FreqAug stochastically eliminates particular frequency components from the video to help the latent representation better capture meaningful features from the remaining data.
3.2.2. Sample Generation and Curation
Sample generation is a popular technique that increases the size and diversity of the training data when the data is scarce. It explicitly creates new samples by transformation or generative models. For enhancing the noise resilience, Nguyen et al. (2023) propose a novel noise-resilient sampling strategy by exploiting a parameter-free discrete wavelet transform low-pass filter to generate a perturbed version of the original time series. By leveraging the LLM, LAVILA (Zhao et al., 2023) learns better video-language embedding when only few text annotations are available. Using the available video-text data, it first fine-tunes LLM to generate text narration given visual input. Then, densely annotated videos via a fine-tuned narrator are used for video-text contrastive learning. Regarding curation, Liu et al. (2024) introduce a large-scale unified time-series dataset, encompassing seven domains with up to 1 billion time points. Similarly, another study (Goswami et al., 2024) present a comprehensive benchmark called Time Series Pile, which aggregates a diverse collection of publicly available datasets across 13 unique domains, comprising 13 million unique time series and 1.23 billion timestamps. These extensive datasets are expected to to play a crucial role in training large-scale time-series models that can be transferred to various data-scarce scenarios.
Figure 6. Illustrative examples of neural architectural approaches.
[…]
Illustrative examples of neural architectural approaches.
4. Neural Architectural Approaches
As neural architectures play a crucial role in the quality of representations (Trirat et al., 2020), this section examines novel network architecture designs aimed at enhancing representation learning. These improvements (depicted in Fig. 6) include, for example, better temporal modeling, handling missing values and irregularities, and extracting inter-variable dependencies in multivariate time series.
4.1. Task-Adaptive Submodules
Chen et al. (2021a) propose multi-task representation learning method called MTRL by exploiting supervised learning for classification and unsupervised learning for retrieval. MTRL jointly optimizes the two downstream tasks via a combination of deep wavelet decomposition networks to extract multi-scale subseries and 1D-CNN residual networks to learn time-domain features, thus improving the performance of downstream tasks. Another method using the wavelet transform, called WHEN (Wang et al., 2023c), newly designs two types of attention modules: WaveAtt and DTWAtt. In the WaveAtt module, the study proposes a novel data-dependent wavelet function to analyze dynamic frequency components in non-stationary time series. In the DTWAtt module, WHEN transforms the DTW technique into the form of the DTW attention. Here, all input sequences are synchronized with a universal parameter sequence to overcome the time distortion problem in multiple time series. Then, the outputs from the new modules are further combined with task-dependent neural networks to perform the downstream tasks, such as forecasting. Liang et al. (2023a) introduce UniTS, which uses a pre-training module that consists of templates from various self-supervised learning methods. Subsequently, the pre-trained representations are fused. Then, the results of feature fusion are applied to task-specific output models. Due to the proliferation of edge devices, a study (Gorbett et al., 2023) proposes a novel model compression technique to make lightweight Transformers for multivariate time-series problems using network pruning, weight binarization, and task-specific modification of attention modules that can substantially reduce both model size (i.e., # parameters) and computational complexity (i.e., FLOPs). The paper demonstrates that compressed Transformers using the proposed technique achieve comparable accuracy to their original counterparts (Zerveas et al., 2021) despite the substantial reduction. These compressed neural networks have the potential to enable DL-based models across new applications and smaller computational environments.
Recently, Gao et al. (2024) present an innovative approach to time-series analysis that unifies various tasks within a single, adaptable model. This model leverages a token-based architecture inspired by LLMs, enabling it to handle diverse time series tasks without requiring task-specific modules. The design’s strength lies in its use of sequence tokens, prompt tokens, and task tokens, which provide context and instructions, allowing the model to quickly adapt to new tasks across different domains and datasets without the need for fine-tuning. Similarly, Zhang et al. (2024) propose a pretraining-based encoder-decoder network with sparse dependency graph construction and temporal-channel layers. A sparse dependency graph is constructed to capture the dependencies between different channels in the multivariate data. The temporal-channel layers sit between the frozen pre-trained encoder and the decoder. These layers are composed of a standard Transformer layer combined with a Graph Transformer layer, which takes the sparse dependency graph as input. This allows the model to capture both temporal and cross-channel dependencies more effectively during fine-tuning for different downstram tasks. MOMENT (Goswami et al., 2024) leverages transformer architectures and adapts them for time series tasks by using techniques like masking and patching to manage varying time-series lengths and complexities. MOMENT models are versatile and can be fine-tuned for various tasks by using a different projection head, e.g., reconstruction or forecasting.
4.2. General Temporal Modeling
4.2.1. Intra-Variable Modeling
Studies in this group focus on capturing the patterns and dependencies within individual time series variables (or features). We classify these studies into long-term and multi-scale modeling based on their emphasis on understanding the temporal dynamics.
Long-Term Modeling
To improve efficiency and scalability for learning long time series, Franceschi et al. (2019) adopt an encoder-only network based on dilated causal 1D-CNN in their neural architecture design. MemDPC (Han et al., 2020) is a novel framework with memory-augmented networks for video representation using dense predictive coding that incorporates an external memory module to store and retrieve relevant information over long periods, while allowing for better temporal coherence. The model is trained on a predictive attention mechanism over the set of compressed memories. TST (Zerveas et al., 2021) is the first Transformer-based framework for representation learning of multivariate time series. TST combines a base Transformer network with an additional input encoder and a learnable positional encoding to make time series work seamlessly with the Transformer encoder. This paper argues that the multi-headed attention mechanism makes the Transformer models suitable for time-series data by concurrently representing each input sequence element using its future-past contexts, while multiple attention heads can consider different representation subspaces. Tonekaboni et al. (2022) propose a decoupled global and local encoders to learn representations for global and local factors in time series by encouraging decoupling between them through counterfactual regularization minimizing mutual information. From another different perspective, Zhang et al. (2023c) redesign standard multi-layer encoder-decoder sequence models to learn time-series processes as state-space models via the companion matrix and a new closed-loop view of the state-space models, resulting in a new module named SpaceTime layer, which consists of 1D-CNN and feed-forward networks. Multiple SpaceTime layers are stacked to form the final encoder-decoder architecture. Nguyen et al. (2023) propose CoInception by integrating the dilated CNN into the Inception block to build a scalable and robust neural architecture with a wide receptive field. Specifically, it incorporates a novel convolution-based aggregator and extra skip connections in the Inception block to enhance the capability for capturing long-term dependencies in the input time series. Lately, ModernTCN (Luo and Wang, 2024), a modernized version of the traditional temporal convolutional network (TCN), is specifically designed to achieve a significantly larger effective receptive field (ERF). This larger ERF helps the model to better capture dependencies across time steps.
Multi-Scale and Multi-Resolution Modeling
To better model multi-scale temporal patterns, MSD-Mixer (Zhong et al., 2023) employs MLPs along different dimensions to mix intra- and inter-patch variations generated from a novel multi-scale temporal patching approach. It learns to explicitly decompose the input time series into different components by generating the component representations in different layers. Based on the existing TCN architecture, Fraikin et al. (2023) attach a time-embedding module to explicitly learn time-related features, such as trend, periodicity, and distribution shifts. It jointly learns these time-embeddings alongside feature extraction, which allows for a more nuanced understanding of different temporal scales. CoInception (Nguyen et al., 2023) also uses multi-scale filters to capture the temporal dependency at multiple scales and resolutions. COMET (Wang et al., 2023a) consists of observation-level, sample-level, trial-level, and patient-level contrastive blocks to represent multiple levels of medical time series. These blocks help to capture cross-sample features and get robust representations. Wang et al. (2024c) introduce a token blend module, allowing a model to process information at multiple scales, which is critical for effectively capturing long-term dependencies in time series. To enhance the model’s ability to capture both short-term and long-term dependencies, TSLANet (Eldele et al., 2024) incorporates a novel adaptive spectral block that leverages Fourier analysis, which also mitigates noise through adaptive thresholding.
T-C3D (Liu et al., 2020) integrates a residual 3D-CNN and temporal encoding to capture static and dynamic features of actions in videos across various temporal scales, achieving universal time-series representation learning that encompasses both short-term and long-term actions. Sener et al. (2020) suggest a flexible multi-granular temporal aggregation framework with non-local blocks, coupling block, and temporal aggregation block to solve reasoning from current and past observations for long-range videos. TCGL (Liu et al., 2022) considers multi-scale temporal dependencies within videos based on spatial-temporal knowledge discovering modules. It jointly models the inter-snippet and intra-snippet temporal dependencies through a hybrid graph contrastive learning strategy. Sanchez et al. (2019) combine VAE and GAN to effectively model the characteristics of satellite image time series across various resolutions and temporal scales. By doing so, it enables the creation of effective disentangled image time series representations that separate shared common information across the time-series images from exclusive information unique to each image.
4.2.2. Inter-Variable Modeling
Guo et al. (2021) introduce a separable self-attention (SSA) module designed to capture spatial and temporal correlations in videos separately. The proposed SSA improves the understanding of actions in videos and demonstrating superior performance in action recognition and video retrieval tasks. For time series domains that exhibit a common causal structure but possess varying time lags, SASA (Cai et al., 2021) aligns the source and target representation spaces by establishing alignment on intra-variable and inter-variable sparse associative graphs. MARINA (Xie et al., 2022) comprises a temporal module learning temporal correlations using MLP and residual connections alongside a spatial module capturing spatial correlations between time-series data using GAT. It comprehends relationships among various variables and grasps intricate patterns within time series data, thereby achieving universal and flexible time-series representation learning. Wang et al. (2024b) propose a fully-connected spatial-temporal graph neural network (FC-STGNN) to model the spatio-temporal dependencies of multivariate time-series data. FC-STGNN consists of a fully-connected graph to model correlations between various sensors and a moving-pooling GNN layer to capture local temporal patterns. MSD-Mixer (Zhong et al., 2023) introduces a novel temporal patching technique that breaks down the time series into multi-scale patches, helping the model to better capture both intra- and inter-patch variations and correlations between different channels.
Xiao et al. (2023) propose a novel positional embedding, group embedding, which assigns input instances to a set of learnable group tokens to embed instance-specific inter-channel relationships and temporal structures. Grouping occurs in two sequential transformers from channel-wise and temporal perspectives. HierCorrPool (Wang et al., 2023b) is a new framework that captures both hierarchical correlations and dynamic properties by using a novel hierarchical correlation pooling scheme and sequential graphs. A recent work, CARD (Wang et al., 2024c), is proposed to effectively capture dependencies across multiple channels (variables) by incorporating channel alignment that allows the model to share information among different channels, improving the ability to capture inter-dependencies. Luo and Wang (2024) propose a modernized TCN model by separating the processing of temporal and feature information, which is a departure from traditional CNNs that typically mix these aspects together. This separation is achieved through depth-wise convolution and convolutional feed-forward networks. Zhang et al. (2024) introduce a new pre-training framework, UP2ME, that constructs a dependency graph among channels to capture cross-channel relationships when the pre-trained model is fine-tuned on multivariate time series. UP2ME incorporates learnable temporal-channel layers that adjust both temporal and cross-channel dependencies.
4.2.3. Frequency-Aware Aggregation
Wang et al. (2018) propose a wavelet-based neural architecture, called mWDN, by integrating multi-level discrete wavelet decomposition into existing neural networks for building frequency-aware deep models. This integration enables the fine-tuning of all parameters within the framework while preserving the benefits of multi-layer discrete wavelet decomposition in frequency learning. Yang and Hong (2022) propose an unsupervised representation learning framework for time series, named BTSF. BTFS enhances the representation quality through the more reasonable construction of contrastive pairs and the adequate integration of temporal and spectral information. BTSF constitutes an iterative application of a novel bi-linear temporal-spectral fusion, explicitly encoding affinities between time and frequency pairs. To adequately use the informative affinities, BTSF further uses a cross-domain interaction with spectrum-to-time and time-to-spectrum aggregation modules to iteratively refine temporal and spectral features for cycle update, proven effective by empirical and theoretical analysis.
Another recent method (Wang et al., 2023c) introduces a data-dependent wavelet function within a BiLSTM network to analyze dynamic frequency components of non-stationary time series. Zhou et al. (2023b) design a frequency adapter using fast Fourier transform to capture the frequency domain based on a pre-trained language model. TimesNet (Wu et al., 2023) introduces TimesBlocks, using Fourier transform to extract periods and Inception blocks for efficient parameter extraction, followed by adaptive aggregation using amplitude values. Xu et al. (2024) propose FITS, a lightweight yet powerful model for time-series analysis, which extends time series segments by interpolating in the complex frequency domain. Eldele et al. (2024) integrate an adaptive spectral block, leveraging Fourier analysis to enhance feature representation and effectively handle noise through adaptive thresholding. The authors also employ an interactive convolution block to improve its ability to decode complex temporal patterns. MF-CLR (Duan et al., 2024) introduces a novel approach, especially focusing on datasets where the data is collected at multiple frequencies, such as in financial markets. MF-CLR has a hierarchical mechanism that processes different frequency components of time-series data separately. It creates embeddings by contrasting subseries with adjacent frequencies, ensuring that the model can capture the relationships between different frequency bands effectively.
4.3. Auxiliary Feature Extraction
4.3.1. Shapelet and Motif Modeling
Liang et al. (2023b) propose CSL, a unified shapelet-based encoder with multi-scale alignment, to transform raw multivariate time series into a set of shapelets and learn the representations using the shapelets. Qu et al. (2024) demonstrate that shapelets, traditionally used for time-series modeling, are equivalent to specific CNN kernels that involves a squared norm and pooling operation. The authors propose a novel CNN layer called ShapeConv, where the kernels act as shapelets, enabling high interpretability with shaping regularization.
4.3.2. Contextual Modeling and LLM Alignment
Audio word2vec (Chen et al., 2019) extends vector representations to consider the sequential phonetic structures of the audio segments trained with speaker-content disentanglement based segmental sequence-to-sequence autoencoder. For graph time series, STANE (Liu et al., 2019) guides the context sampling process to focus on the crucial part of the data in the graph attention networks. Rahman et al. (2021) introduce a tri-modal VilBERT-inspired model by integrating separate encoders for vision, pose, and audio modalities into a single network. DelTa (Anand and Nayak, 2021) uses 2D images of time series such that pre-trained models on large image datasets can be used. DelTa proposes two versions of using pre-trianed vision models: layout aligned version and layout independent version. Lee et al. (2022b) present a novel BMA-Memory framework for bimodal representation learning, focusing on sound and image data. This memory system allows for the association of features between different modalities, even when data pairs are weakly paired or unpaired. Following TST (Zerveas et al., 2021), Chowdhury et al. (2022) propose TARNet to reconstruct important timestamps using a newly designed masking layer to improve downstream task performance. It decouples data reconstruction from the downstream task and uses a data-driven masking strategy instead of random masking via self-attention score distribution generated by the transformer encoder during the downstream task training to determine a set of important timestamps to mask. Kim et al. (2023a) introduce MLP-based feature-wise encoder together with a element-wise gating layer built on top of TS2Vec (Yue et al., 2022), i.e., feature-agnostic temporal representation using TCN, to flexibly learn the influence of feature-specific patterns per timestamp in a data-driven manner. Choi and Kang (2023) also extend the TS2Vec encoder (Yue et al., 2022) with multi-task self-supervised learning by combining contextual, temporal, and transformation consistencies into a single networks.
One Fits All (Zhou et al., 2023a) uses a pre-trained language model (e.g., GPT-2) by freezing self-attention and feed-forward layers and fine-tuning the remaining layers. Input embedding and normalization layers are modified for time-series data. By doing so, it can benefit from the universality of the Transformer models on time-series data. Based on a pre-trained language model, Zhou et al. (2023b) additionally design task-specific gates and adapters. These adapters allow the model to effectively leverage the generalization capabilities of the pre-trained LM while adapting to the specific demands of various time series tasks. For example, temporal adapters focus on modeling time-based correlations, channel adapters on handling multi-dimensional data, and frequency adapters on capturing global patterns through Fourier transforms. DTS (Li et al., 2022) is a disentangled representation learning framework for semantic meanings and interpretability through two disentangled components. The individual factors disentanglement extracts different semantic independent factors. The group segment disentanglement makes a batch of segments to enhance group-level semantics.
NuTime (Lin et al., 2024) introduces a novel approach to pre-training models on large-scale time series by leveraging a numerically multi-scaled embedding (NME) technique. The model starts by partitioning the data into non-overlapping windows, each represented by three components: its normalized shape, mean, and standard deviation. These components are then concatenated and transformed into tokens suitable for a Transformer encoder. By considering all possible numerical scales, NME enables the model to effectively handle scalar values of arbitrary magnitudes within the data. This technique ensures the smooth flow of gradients during training, making the model highly effective for large-scale time series data. UniTTab (Luetto et al., 2023) is a Transformer-based framework for time-dependent heterogeneous tabular data. It uses row-type dependent embedding and different feature representation methods for categorical and numerical data, respectively. Li et al. (2024) employ an LLM-based encoder initialized with pre-trained weights from the text encoder of CLIP (Radford et al., 2021) designed to maintain variable independence to address the challenge of embedding time-series data from multiple domains without introducing domain-specific biases. This approach enables the pre-trained models to be highly adaptable and effective across various time series tasks.
4.4. Continuous Temporal and Irregular Time-Series Modeling
4.4.1. Neural Differential Equations (NDE)
ODE-RNN (Rubanova et al., 2019b) is an early attempt applying neural ordinary differential equations (Neural ODE) for time series. It serves as an encoder in a latent ODE model, facilitating interpolation, extrapolation, and classification tasks for ISTS. ANCDE (Jhin et al., 2021) and EXIT (Jhin et al., 2022) propose neural controlled differential equation (Neural CDE)-based approaches for classification and forecasting with ISTS. ANCDE leverages two Neural CDEs, one for learning attention from the input time series and another for creating representations for downstream tasks. Likewise, EXIT uses Neural CDEs as part of the encoder-decoder network, enabling interpolation and extrapolation of ISTS. Also, CrossPyramid (Abushaqra et al., 2022) addresses the limitation of ODE-RNN which is the high dependence on the initial observations by using pyramid attention and cross-level ODE-RNN. Contiformer (Chen et al., 2023) merges NDEs into the attention mechanism of a transformer by modeling key and value with Neural ODEs. Oh et al. (2024) enhance the stability and performance of neural stochastic differential equations (Neural SDEs) when applied to ISTS with three distinct classes of Neural SDEs. Each class features unique drift and diffusion functions designed to address the challenges of stability and robustness in the stochastic modeling of ISTS. The authors also emphasize the importance of well-defined diffusion functions to prevent issues like stochastic destabilization.
4.4.2. Implicit Neural Representations (INR)
HyperTime (Fons et al., 2022) introduces INR for time series used for imputation and reconstruction by taking timestamps as input and outputting the original time series. It consists of two networks Set Encoder and HyperNet Decoder. Naour et al. (2023) introduce TimeFlow to deal with modeling issues such as irregular time steps. With a hyper-network, the framework modulates INR that is a parameterized continuous function on multiple time series.
4.4.3. Auxiliary (Sub-)Networks
LIME-RNN (Ma et al., 2019) introduces a weighted linear memory vector into RNNs for time-series imputation and prediction. Another work by Bianchi et al. (2019) proposes a temporal kernelized autoencoder to learn representations aligned to a kernel function which are designed for handling missing values. mTAN (Shukla and Marlin, 2021) learns the representation of continuous time values by applying the attention mechanism to ISTS. The key innovation of mTANs lies in its continuous-time attention mechanism, which generalizes positional encoding typically used in transformers to operate in continuous time rather than discrete steps. This mechanism leverages multiple time embeddings to flexibly capture both periodic and non-periodic patterns in the data, allowing the network to produce fixed-length representations of time series regardless of the number or irregularity of observations. TE-ESN (Sun et al., 2021) employs a requisite time encoding mechanism to acquire knowledge from ISTS, with the representation being learned within echo state networks.
Continuous recurrent units (Schirmer et al., 2022) update hidden states based on linear stochastic differential equations, which are solved by the continuous-discrete Kalman filter. Based on continuous-discrete filtering theory, Ansari et al. (2023) introduce a neural continuous-discrete state space model to model continuous-time modeling of ISTS. Biloš et al. (2023) suggest a representation learning approach based on denoising diffusion models adapted for ISTS with complex dynamics. By gradually adding noise to the entire function, the model can effectively capture underlying continuous processes. TriD-MAE (Zhang et al., 2023a) is a pre-trained model based on TCN blocks with attention scale fusion to handle ISTS. Senane et al. (2024) propose a novel architecture called TSDE specifically designed to handle ISTS with dual-orthogonal Transformer encoders that are integrated with a crossover mechanism. This structure processes the observed segments of the time series, which are divided by an imputation-interpolation-forecasting mask. The architecture then conditions a reverse diffusion process on these embeddings to predict and correct noise in the masked parts of the series, allowing TSDE to generate robust representations, making it particularly effective for irregular and noisy time series.
Figure 7. Illustrative examples of learning-focused approaches.
[…]
Illustrative examples of learning-focused approaches.
5. Learning-Focused Approaches
Studies in this category center on devising novel learning objective functions for the representation learning process, i.e., model (pre-)training. As in Fig. 7, these studies can be classified into three groups based on the learning objectives: task-adaptive, non-contrasting, and contrasting losses.
5.1. Task-Adaptive Objective Functions
Ma et al. (2019) develop a set of task-specific loss functions to train the LIME-RNN with incomplete time series in an end-to-end way, simultaneously achieving imputation and prediction. TriBERT (Rahman et al., 2021) designs separate losses for a specific modality. Weakly supervised classification for vision-based modalities and classification loss for audio modality. Hadji et al. (2021) use a soft version of DTW (soft-DTW) to compute the loss between two videos with the same class in the weakly-supervised setting. Cycle-consistency is forced by matching two DTW results ($X\to Y$ and $Y\to X$) as the same. Zhong et al. (2023) propose MSD-Mixer with a novel loss function to constrain both the magnitude and auto-correlation of the decomposition residual used together with the supervised loss of the target downstream task during training. This loss function facilitates MSD-Mixer to extract more temporal patterns into the components to be used for the downstream task. Wang et al. (2024c) propose a robust loss function designed to mitigate over-fitting by weighting forecasting errors based on prediction uncertainties, which contributes to more accurate and robust time-series forecasting and anomaly detection. UniTS (Gao et al., 2024) unifies multiple tasks within a single model framework by using a universal task specification that allows a single model to handle diverse downstream tasks without the need for task-specific modules. It employs a prompting-based framework to convert different tasks into a unified token representation. This unified approach with a single set of shared weights can generalize better and perform multiple tasks simultaneously.
5.2. Non-Contrasting Losses for Temporal Dependency Learning
5.2.1. Input Reconstruction
Sqn2Vec (Nguyen et al., 2018) learns low-dimensional continuous feature vectors for sequence data by predicting singleton symbols and sequential patterns to satisfy a gap constraint. Chen et al. (2019) introduce unsupervised training of audio word2vec using a sequence-to-sequence autoencoder on unannotated audio time series. The encoder includes a segmentation gate trained with reinforcement learning to represent utterances as vectors carrying phonetic information. Wave2Vec (Yuan et al., 2019) jointly models inherent and temporal representations of biosignals and provides clinically meaningful interpretation with enhanced interpretability. Sanchez et al. (2019) propose a method for satellite image time series by combining VAE and GAN to learn image-to-image translation, where one image in a time series is translated into another.
Ti-MAE (Li et al., 2023), a masked autoencoder framework, addresses the distribution shift problem by learning strong representations with less inductive bias or hierarchical trick. Its masking strategy creates different views for the encoder in each iteration, fully leveraging the whole input time series during training. Chen et al. (2024) use probabilistic masked autoencoding where segment-wise masking schemes and rate-aware positional encodings are devised to enable the characterization of multi-scale temporal dynamics. In the pre-training phase, the encoders generate rich and holistic representations of multi-rate time series. A temporal alignment mechanism is devised to refine synthesized features for dynamic predictive modeling through feature block division and block-wise convolution. Lee et al. (2024a) argue that embedding time series patches independently rather than capturing dependencies between them can lead to better representation learning and introduce a simple patch reconstruction task, where each time series patch is autoencoded without considering other patches. This task helps in learning meaningful representations for each patch independently.
5.2.2. Masked Prediction
TST (Zerveas et al., 2021) adopts an encoder-only Transformer network with a masked prediction (denoising) objective. The losses of TST are computed only from the masked parts or timestamps. Specifically, TST trains the Transformer encoder to extract dense vector representations of multivariate time series using the denoising objective on randomly masked time series. Similar to TST, TARNet (Chowdhury et al., 2022) improves downstream task performance by learning to reconstruct important timestamps using a data-driven masking strategy. The masked timestamps are determined by self-attention scores during the downstream task training. This reconstruction process is trained alternately with the downstream task at every epoch by sharing parameters in a single network. It also enables the model to learn the representations via task-specific reconstruction, which results in improved downstream task performance. Biloš et al. (2023) deal with ISTS by treating time-series data as discretization of continuous functions and using denoising diffusion models. By masking some parts of the continuous function and predicting the rest, it captures the temporal dependencies of time-series data and enables generalization across various types of time series.
SimMTM (Dong et al., 2023) reconstructs the original series from multiple masked series with series-wise similarity learning and point-wise aggregation to reveal the local structure of the manifold implicitly. It introduces a neighborhood aggregation design for reconstruction by aggregating the point-wise representations of time series based on the similarities learned in the series-wise representation space. Thus, the masked time points are recovered by weighted aggregation of multiple neighbors outside the manifold, allowing SimMTM to assemble complementary temporal variations from multiple masked time series and improve the quality of the reconstruction. In addition, a constraint loss is proposed to guide the series-wise representation learning based on the neighborhood assumption of the time series manifold. UP2ME (Zhang et al., 2024) uses task-agnostic univariate pre-training that involves generating univariate instances by varying window lengths and decoupling channels. These instances are then used for pre-training a masked autoencoder, focusing only on temporal dependencies without considering cross-channel dependencies, which allows it to perform tasks immediately after pre-training by simply formulating them as specific mask-reconstruction problems. Senane et al. (2024) propose a time series diffusion embedding model. It combines a diffusion process with a novel imputation-interpolation-forecasting mask to enhance the learning of time-series representations by segmenting time series into observed and masked parts.
5.2.3. Customized Pretext Tasks
Wu et al. (2018) propose random warping series, a positive-definite kernel based on DTW. By computing DTW between original time series and a distribution of random time series, the model generates vector representations, thus learning the structure and patterns of the time-series data. Sqn2Vec (Nguyen et al., 2018) extracts sequential patterns (SP) satisfying gap constraints from the given sequence data and includes the process of learning vectors for each sequence considering these SPs. This enables learning meaningful low-dimensional continuous vectors for sequences and obtaining representations that reflect temporal dependencies. Lee et al. (2022b) obtain universal and robust representations by learning the association between sound and image representations using both weakly-paired and unpaired data. Fang et al. (2023) propose a method to capture the spatial relation shared across all instances. A prior graph is learned to minimize the distance between dynamic graph. Moreover, to capture the instance-specific spatial relation, a dynamic graph is learned to maximize the distance between the prior graph. T-Rep (Fraikin et al., 2023) leverages the time-embedding module in its pretext tasks which enable the model to learn fine-grained temporal dependencies, giving the latent space a more coherent temporal structure than existing methods.
Sun et al. (2023) introduce TEST to convert time series into text data, enabling the learning of time-series representations using LLM’s text understanding capability. TEST tokenizes the time-series data and converts it into text-form embeddings through an encoder that LLM can comprehend. Timer (Liu et al., 2024) adopts a GPT-style, decoder-only architecture that is pre-trained using next-token prediction. This approach is similar to how LLMs are trained and enables the model to learn complex temporal dependencies within the data. Bian et al. (2024) adapt LLMs for time series by treating time-series forecasting as a self-supervised, multi-patch prediction task. This strategy allows the model to better capture temporal dynamics within time series. The first stage involves training the LLMs on various time-series datasets with a focus on next-patch prediction to synchronize the LLMs’ capabilities with the specific characteristics of time series. In the second stage, the model is fine-tuned specifically for multi-patch prediction tasks in targeted time-series contexts. This fine-tuning helps the model learn more detailed and contextually relevant temporal patterns. During decoding, this framework uses a patch-wise decoding layer instead of the conventional sequence-level decoding to allow each patch to be decoded independently into temporal sequences. Dong et al. (2024) introduce TimeSiam, which is tasked with reconstructing the masked portions of the current subseries based on the past observations. This helps the model learn robust time-dependent representations. To capture temporal diversity, TimeSiam introduces lineage embeddings to differentiate between the temporal distances of sampled series, allowing the model to better understand and learn from diverse temporal correlations.
Wang et al. (2020) use a pretext task that classifies the pace of a clip into five categories: super slow, slow, normal, fast, and super fast. Haresh et al. (2021) use the soft-DTW loss for temporal alignment between two videos. Wang et al. (2021) propose a new pretext task using motion-aware curriculum learning. Several spatial partitioning patterns are employed to encode rough spatial locations instead of exact spatial Cartesian coordinates. Liang et al. (2022) introduce three pretext tasks, including clip continuity prediction, discontinuity localization, and missing section approximation, based on video continuity by cutting a clip into three consecutive short clips. CACL (Guo et al., 2022) captures temporal details of a video by learning to predict an approximate Edit distance between a video and its temporal shuffle. In addition, contrastive learning is performed by generating four positive samples from two different encoders, 3D CNN and video transformer, as well as two different augmentations. Duan et al. (2022) propose TransRank with a pretext task aiming to assess the relative magnitude of transformations. It allows the model to capture inherent characteristics of the video, such as speed, even when the challenges of matching the transformations across videos vary. CSTP (Zhang et al., 2022a) uses a pretext task called spatio-temporal overlap rate prediction that considers the intermediate of contrastive learning. With a joint optimization combining pretext tasks with contrastive learning, CSTP enhances the spatio-temporal representation learning for downstream tasks.
5.3. Contrasting Losses for Consistency Learning
5.3.1. Subseries Consistency
Franceschi et al. (2019) employ the triplet loss (i.e., T-Loss) inspired by word2vec (Goldberg and Levy, 2014) for learning scalable representations of multivariate time series. It considers a sub-segment belonging to the input time segment as a positive sample to explore sub-series consistency. Somaiya et al. (2022) propose TS-Rep by combining the T-Loss (Franceschi et al., 2019) with the use of nearest neighbors to diversify positive samples. The method is particularly effective in handling the complexity and variability inherent in robot sensor data. Behrmann et al. (2021) separate the representation space into stationary and non-stationary characteristics through contrastive learning from long and short views, enhancing both the video-level and temporally fine-grained tasks. Qian et al. (2022) propose a windowing based learning by sampling a long clip from a video and a short clip that lies inside the duration of the long clip. These long and short clips become a positive pair for contrastive learning, and other long clips become negative instances. When conducting contrastive learning, final vectors are made in two different embedding spaces. The first one is a fine-grained space and each embedding is made at each timestamp. The second embedding space is a persistent embedding space and each timestamp embedding is global average pooled for contrastive learning. Wang et al. (2020) select positive pairs from a pair of clips in the same video and negative pair from a pair of clips in different videos. CVRL (Qian et al., 2021) uses the temporally consistent spatial augmentation and clip selection strategy, where each frames are spatially augmented. Here, two clips in the same video are positive, while two clips in the different videos are negative.
5.3.2. Temporal Consistency
TNC (Tonekaboni et al., 2021) leverages temporal neighborhoods as positive samples through the ADF test. By incorporating sample weight adjustment into contrastive loss, the sampling bias problem is alleviated. Eldele et al. (2021) incorporate a novel temporal contrasting module to learn temporal dependencies by designing a hard cross-view prediction task that uses past latent features of one augmentation to predict the future of another augmentation for a certain time step. This operation forces the model to learn robust representation by a harder prediction task against any perturbations introduced by different time steps and augmentations. Hajimoradlou et al. (2022) introduce similarity distillation along the temporal and instance dimensions for pre-training universal representations. Yang et al. (2022b) propose TimeCLR, which enables a feature extractor to learn invariant representations by minimizing the similarity between two augmented views of the same sample. Wang et al. (2024b) integrate the FC graph construction with the moving-pooling GNN. The model is capable of learning high-level features that represent both the spatial and temporal aspects of multivariate time series. This dual focus ensures that the temporal consistency is preserved while also accounting for spatial correlations, which is critical for downstream tasks.
Morgado et al. (2020) use audio-visual spatial alignment as a pretext task with 360° video data. This pretext task involves spatially misaligned audio and video clips, treated as negative examples for contrastive learning. Besides the non-contrastive loss, Haresh et al. (2021) jointly use a temporal regularization term (i.e., Contrastive-IDM) to encourage two different frames to be mapped to different points in the embedding space. Chen et al. (2022) propose sequence contrastive loss to sample two subsequences with an overlap for each video. The overlapped timestamps are considered positives, while the clips from other videos are negatives. Two timestamps neighboring each other also become positive pairs with the Gaussian weight proportional to the temporal distance. A recent study (Yang et al., 2023a) proposes TempCLR to explore temporal dynamics in video-paragraph alignment, leveraging a novel negative sampling strategy based on temporal granularity. By focusing on sequence-level comparison using DTW, TempCLR captures temporal dynamics more effectively. Zhang et al. (2023b) model videos as stochastic processes by enforcing an arbitrary frame to agree with a time-variant Gaussian distribution conditioned on the start and end frames.
5.3.3. Contextual Consistency
TimeAutoML (Jiao et al., 2020) adopts the AutoML framework, enabling automated configuration and hyperparameter optimization. Negative samples are created by introducing random noise within the range defined by the minimum and maximum values of the given instances. CARL (Chen et al., 2022) uses a sequence contrastive loss to learn representations by aligning the sequence similarities between augmented video views. This loss function helps maintain temporal coherence and contextual consistency across frames, making the representations robust to variations in video length and content. TS2Vec (Yue et al., 2022) uses randomly overlapped segments to capture multi-scale contextual information through temporal and instance-wise contrastive losses. Zhang et al. (2023f) introduce TS-CoT, which is a co-training algorithm that enhances the global consistency of representations from different views. REBAR (Xu et al., 2023) exploits retrieval-based reconstruction to capture the class-discriminative motifs. This method selects positive samples using the REBAR cross-attention reconstruction trained with a contiguous and intermittent masking strategy. Shin et al. (2023) propose a consistency regularization framework based on two overlapping windows and leverage a merged soft label from those two windows as a shared target. Zhong et al. (2023) propose a novel residual loss to ensure that the model’s decomposition leaves only white noise as residual, further contributing to the contextual consistency of the analysis by reducing information loss during the decomposition. PrimeNet (Chowdhury et al., 2023) generates augmented samples based on the observation density and employs a reconstruction task to facilitate the learning of irregular patterns.
5.3.4. Transformation Consistency
TS-TCC (Eldele et al., 2021) transforms a given time series into two different yet correlated views by weak and strong augmentations. Then, it learns discriminative representations by maximizing the similarity among different contexts of the same sample while minimizing similarity among contexts of different samples, leading to high efficiency in few-labeled data and transfer learning scenarios. RSPNet (Chen et al., 2021b) solves the relative speed perception task and appearance-focused video instance discrimination task using the triplet loss and InfoNCE loss. Jenni and Jin (2021) propose a time-equivariant model using a pair of clips as the unit of contrastive learning. If two pairs have same temporal transformation within each pair, then the output of each clip is concatenated for each pair and the concatenated outputs become similar. Auxiliary tasks are also used, e.g., classifying the speed difference between two clips: clip A as 2x and clip B as slow speed.
5.3.5. Hierarchical and Cross-Scale Consistency
Yue et al. (2022) propose TS2Vec, which is based on novel contextual consistency learning using two contrastive policies over two augmented time segments with different contexts from randomly overlapped subsequences. Unlike other methods, TS2Vec performs contrastive learning in a hierarchical way over the augmented context views, making a robust contextual representation for each timestamp. Its overall representation can be obtained by max pooling over the corresponding timestamps. By using multi-scale contextual information with granularity, this approach can capture multi-scale contextual information with temporal and instance-wise contrastive losses for the given time series and generate fine-grained representations for any granularity. Nguyen et al. (2023) design a new loss function by combining ideas from hierarchical and triplet losses. CSL (Liang et al., 2023b) proposes to learn time-series representations shapelets with multi-grained contrasting and multi-scale alignment for capturing information in various time ranges. COMET (Wang et al., 2023a) incorporates four-level contrastive losses for medical time series. The overall loss has hyper-coefficients about each loss to compromise the multiple hierarchical contrastive losses. By applying contrastive learning across these multiple levels, the framework can create more robust and comprehensive representations of the data. Duan et al. (2024) leverage a hierarchical structure that captures the different frequencies present in the data and embeds subseries of the time series into two groups based on adjacent frequencies. By enforcing consistency between these groups, it create robust representations that are useful across various frequencies.
CCL (Kong et al., 2020) uses the inclusive relation of a video and frames for a contrastive learning strategy. The video and frames of the inclusive relation are learned to be close to each other in the embedding spaces. Qing et al. (2022) learn representations on untrimmed video to reduce the amount of labor required for manual trimming and use the rich semantics of untrimmed video. Hierarchical contrastive learning teaches clips that are near in time and topic to be similar. Zhang and Crandall (2022) introduce a model trained with decoupled learning objectives into two contrastive sub-tasks, which are hierarchically spatial and temporal contrast. With graph learning, TCGL (Liu et al., 2022) uses a spatial-temporal knowledge discovering module for motion-enhanced spatial-temporal representations. It introduces intra- and inter-snippet temporal contrastive graphs to explicitly model multi-scale temporal dependencies via a hybrid graph contrastive learning strategy. TCLR (Dave et al., 2022) is a model for video understanding trained with novel local–local and global–local temporal contrastive losses.
5.3.6. Cross-Domain and Multi-Task Consistency
FEAT (Kim et al., 2023a) jointly learns feature-based and temporal consistencies by using hierarchical temporal contrasting, feature contrasting, and reconstruction losses. Choi and Kang (2023) introduce uncertainty weighting approach to weigh multiple contrastive losses by considering homoscedastic uncertainty of multiple tasks, including contextual, temporal, and transformation consistencies. FOCAL (Liu et al., 2023) enforces the modality consistency to learn the features shared across modalities and the transformation consistency to learn the modality-specific feature. To also accommodate sporadic deviations from locality due to periodic patterns, temporally close sample pairs and distant sample pairs are constrained by a loose ranking loss. Zhang et al. (2022b) argue that time-based and frequency-based representations learned from the same time series should be closer to each other in the time-frequency latent space than representations of different time series. Thus, they introduce time-frequency consistency modeling that aims to minimize the distance between time-based and frequency-based embeddings using a novel consistency loss in the same latent space. TimesURL (Liu and Chen, 2023) constructs double Universums as a hard negative, and introduces a joint optimization objective with contrastive learning to capture both segment-level and instance-level information. Lee et al. (2024b) propose SoftCLT with soft contrastive losses for more nuanced learning. It uses soft assignments for an instance-wise contrastive loss based on the distance between time series instances, where a soft assignment determines the degree to which different instances should be contrasted and a temporal contrastive loss that focuses on the differences in timestamps to handle temporal correlations. Li et al. (2024) reveal a positive correlation between representation bias and spectral distance in time series, resulting in a variation of contrasting losses optimized to reduce the bias from data augmentation. These losses include the spectral characteristics of time-series data, ensuring that the embeddings generated are more robust and generalizable. The bias is quantified by comparing the difference between the embeddings of the original time series and the augmented versions to reduce the discrepancy.
MemDPC (Han et al., 2020) involves a predictive attention mechanism applied to a collection of compressed memories. This training paradigm ensures that any subsequent states can be synthesized consistently through a convex combination of the condensed representations. Zeng et al. (2021) improve generalization by learning spatially-local/temporally-global and spatially-global/temporally-local features from audio-visual modalities to capture global and local information in a video. This method enables the capturing of both slowly changing patch-level information and fast changing frame-level information. DCLR (Ding et al., 2022) presents a dual contrastive formulation by decoupling the input RGB video sequence into two complementary modes, static scene and dynamic motion, to avoid static scene bias.
6. Experimental Design
This section describes the typically used experimental design for comparing universal representation learning methods for time series. We describe widely-used protocol and introduce publicly available benchmark datasets with evaluation metrics according to the downstream tasks.
Given a set of $N$ time series $\{(\mathbf{X}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$ and $J$ pre-trained representation learning models $\{f_{e,j}\}_{j=1}^{J}$, this section describes how we evaluate each model to determine the best one. As discussed in Section 1, representations of time series play a vital role in solving time-series analysis tasks. We expect that the learned representations by $f_{e}$ generalize to unseen downstream tasks. Accordingly, the most common evaluation method is how learned representations help solve downstream tasks.
Additionally, we need a function $g_{d}$ that maps a representation (feature) space to a label space. For example, $g_{d}(f_{e}(\mathbf{X})):\mathbb{R}^{R\times F}\to\mathbb{R}^{|C|}$ for classification or $g_{d}(f_{e}(\mathbf{X})):\mathbb{R}^{R\times F}\to\mathbb{R}^{H}$ for forecasting. This is because $f_{e}$ is designed to extract feature representations, not to solve the downstream task. Commonly, $g_{d}$ is implemented as a simple function, such as linear regression, support vector machines, or shallow neural networks because it is enough to solve the downstream task if the learned representations already capture meaningful and discriminative features.
6.1. Evaluation Procedure
Let $\mathcal{D}=\{(\mathbf{X}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$ denote the downstream data. We then compare the encoders $f_{e}$ by using the task-specific evaluation metrics of the downstream task. The evaluation procedures are as follows.
  1. (1)
    Train $f_{e}$ and $g_{d}$ on the downstream dataset $\mathcal{D}$ with the pre-trained encoder $f_{e,j}$ for each $j$.
  2. (2)
    Compare task-specific evaluation metric values, e.g., accuracy for classification or mean square error for regression.
In the first step, there are two common protocols to evaluate the encoders: frozen and fine-tuning.
6.1.1. Frozen Protocol
As we expect $f_{e}$ to learn meaningful representations for downstream tasks, we do not update the pre-trained $f_{e}$, i.e., freezing its weights. Training $g_{d}$ uses less computation budget and converges faster than the training of $f_{e}$ and $g_{d}$ from scratch on the downstream dataset. The standard choice of $g_{d}$ is a linear model (e.g., linear regression and logistic regression) or non-parametric method (e.g., k-nearest neighbors). This evaluation approach is usually referred as linear probing. To further improve performance with additional computing cost, we can also implement $g_{d}$ as a nonlinear model, such as shallow neural networks with a nonlinear activation function.
6.1.2. Fine-Tuning Protocol
In fine-tuning protocol, we train both pre-trained $f_{e}$ and $g_{d}$ as a single model to obtain more performance gain of the downstream task or fill the gap between representation learning and downstream tasks. $g_{d}$ is also known as a task-specific projection head in the representation learning network. This protocol usually uses small learning rate to optimize the model to preserve the original representation quality.
The combination of linear probing and fine-tuning protocol is also possible, especially for out-of-distribution as the encoder’s performance on data sampled from out-of-distribution degrades after fine-tuning. Even though the fine-tuning protocol requires more computing budget than the frozen protocol, it empirically performs better than the frozen protocol (Nozawa and Sato, 2022).
End-to-End Protocol. This protocol is a special case for evaluating a representation learning framework, where both $f_{e}$ and $g_{d}$ are trained jointly from scratch for each downstream task without pre-training in (1). Notable examples include TimesNet (Wu et al., 2023), MSD-Mixer (Zhong et al., 2023), and WHEN (Wang et al., 2023c).
6.2. Benchmark Datasets and Metrics for Downstream Tasks
We summarize widely-used benchmark datasets and evaluation metrics according to the downstream tasks. Some of these datasets are single-purpose datasets for a particular downstream task, and some are general-purpose time-series datasets that we can use for model evaluation across different tasks. Table 3 presents useful information of the reviewed datasets, including dataset names, dimensions, sizes, application domains, and reference sources.
6.2.1. Forecasting and Imputation
Since the output of forecasting and imputation tasks are numerical sequences, most studies uses the same benchmark datasets for both tasks. Commonly used datasets are from several application domains and services, including electricity (e.g., ETT (Zhou et al., 2021)), transportation (e.g., Traffic (Wu et al., 2023)), meteorology (e.g., Weather (Wu et al., 2023)), finance (e.g., Exchange (Lai et al., 2018)) and control systems (e.g., MoJoCo (Jhin et al., 2022)). To facilitate the evaluation of forecasting models, Godahewa et al. (2021) also introduce a publicly accessible archive for time-series forecasting. Given the numerical nature of the predicted results, the most commonly used metrics are mean squared error (MSE) and mean absolute error (MAE).
6.2.2. Classification and Clustering
As the classification and clustering tasks both aim to identify the real category to which a time-series sample belongs, existing studies usually use the same set of benchmark datasets to evaluate the model performance. Curated benchmarks comprising heterogeneous time series from various application domains, such as UCR (Dau et al., 2019) and UEA (Bagnall et al., 2018), are the most widely used because they can provide a comprehensive evaluation regarding the generalization of the model being evaluated. Many researchers also use human activity (e.g., HAR (Anguita et al., 2013)) and health (e.g., Sepsis (Reyna et al., 2020)) related datasets due to their practicality for real-world applications.
Regarding the evaluation metrics, while we can evaluate the classification task with accuracy, precision, recall, and F1 score, we usually assess the clustering task with Silhouette score, adjusted rand index (ARI), and normalized mutual information (NMI) to assess the inherent clusterability due to the absence of label instances. For classification, we may also use the area under the precision-recall curve (AUPRC) to handle the class imbalance cases.
6.2.3. Regression
Compared to the forecasting and classification tasks, time-series regression, particularly with DL, remains relatively underexplored. Only a handful of public benchmark datasets (e.g., heart rate monitoring data (Demirel and Holz, 2023) and air quality (Zhang et al., 2023a)) exist. The TSER archive, introduced by Tan et al. (2021), is the most comprehensive benchmark for time-series regression. Like forecasting and imputation tasks, the metrics for regression are MSE and MAE. Additional metrics, such as root mean squared error (RMSE) and R-squared ($R^{2}$), are also commonly used.
6.2.4. Segmentation
Likewise, time-series segmentation with DL is also relatively underexplored. There are two standard curated benchmarks: UTSA (Gharghabi et al., 2017) and TSSB (Ermshaus et al., 2023). To assess the segmentation performance, F1 and covering scores are typically used. The F1 score emphasizes the importance of detecting the correct timestamps at which the underlying process changes. In contrast, the covering score focuses on splitting a time series into homogeneous segments and reports a measure for the overlaps of predicted versus labeled segments.
6.2.5. Anomaly Detection
Anomaly detection is one of the most popular research topics in time series. There are several benchmarks publicly available, as listed in Table 3. However, as argued by recent studies (Wu and Keogh, 2021; Wagner et al., 2023), most existing benchmarks are deeply flawed and cannot provide a meaningful assessment of the anomaly detection models. Therefore, we recommend using newly proposed datasets, such as ASD (Li et al., 2021), TimeSeAD (Wagner et al., 2023), and TSB-UAD (Paparrizos et al., 2022b).
Concerning the evaluation metrics, point-adjust F1 score (Xu et al., 2018) is the most widely used metric for time-series anomaly detection. Nevertheless, this metric is also found to have an overestimation problem, which cannot give a reliable performance evaluation. Accordingly, recent studies (Yang et al., 2023b; Nam et al., 2023) have started to adopt more robust evaluation metrics, e.g., VUS (Paparrizos et al., 2022a), PA%K (Kim et al., 2022), and eTaPR (Hwang et al., 2022).
6.2.6. Retrieval
Although we find that a few studies of the reviewed articles use particular datasets for the retrieval task (e.g., EK-100 (Damen et al., 2022), Howto100M (Miech et al., 2019), and MUSIC (Zhao et al., 2018)), the task itself can be evaluated with any benchmark dataset because it is basically based on arbitrary query time series. For time-series retrieval (Chen et al., 2021a), benchmark datasets for classification (e.g., UCR) are commonly used. For the evaluation, the top-$k$ recall rate (higher is better) is the standard metric to examine the overlap percentage of the top-$k$ results and the ground truth. $k$ is usually set to 5, 10, and 20.
6.3. Additional Metric for Inherent Representation Quality
Besides the task-specific metrics, recent studies (Wu et al., 2023; Dong et al., 2023) also evaluate the inherent quality of the learned representation by calculating the centered kernel alignment (CKA) similarity between representations from the first and the last layers. A higher CKA similarity indicates more similar representations. As the bottom layer representations usually contain low-level or detailed information, a smaller similarity means the top layer contains different information from the bottom layer and indicates the model tends to learn high-level representations or more abstract information.
In general, better forecasting and anomaly detection accuracy is related to higher CKA similarity, while better imputation and classification results corresponds to the lower CKA similarity. Specifically, the lower CKA similarity means that the representations are distinguishing among different layers, thus being hierarchical representations. These results also indicate the property of representations that each task requires. Thus, we can use this metric to evalaute whether an encoder learn appropriate representations for different tasks.
7. Open Challenges and Future Research Directions
In this section, we discuss open challenges and outline promising future research directions that have the potential to enhance the existing universal time-series representation learning methods.
7.1. Time-Series Active Learning
The complexity and length of time-series data significantly increase the cost of manual labeling. Obtaining annotated time series is challenging due to the domain-specific nature and lack of publicly accessible sources. Even domain experts may struggle to provide consistent annotations because time-series shapes can be perceived differently among annotators. For instance, datasets like Sleep-EDF (PhysioBank, 2000), containing lengthy EEG recordings, require doctors to identify anomalies with a consensus among multiple experts for each annotation. This level of expertise is also demanded by applications such as anomaly detection in smart factory sensors, making the process time-consuming and labor-intensive. Efforts in the DL community have aimed to address label sparsity, notably through active learning (AL) (Shin et al., 2021; Rana and Rawat, 2022). This approach minimizes labeling costs by selecting the most informative unlabeled instances and presenting them to an expert (oracle) for annotation. AL involves developing a query strategy that prioritizes which data points to label next based on predefined criteria, e.g., informativeness and diversity. The process iterates by training a model with the initial labels, selecting the most informative or diverse unlabeled data points for labeling, retraining the model, and repeating this cycle until a certain performance threshold (e.g., annotation budget) is reached. Therefore, developing an effective annotation process with AL for time-series data, enabling a more robust supervision signal, is a promising research direction.
7.2. Distribution Shifts and Adaptation
When a time series model undergoes continuous testing due to the accumulation of time series data over time, various forms of concept drift, such as sudden, gradual, incremental, and recurrent drift, may occur (Agrahari and Singh, 2022). Additionally, a domain shift problem may arise due to differences between the source domain used in the training phase and the target domain used in the test phase. Given that distribution shifts resulting from concept drift and domain shift are factors that degrade model performance, previous studies focus on concept drift adaptation and domain adaptation to address these shifts in specific downstream tasks, such as classification and forecasting (Yuan et al., 2022; Ragab et al., 2023; Jin et al., 2022; Ozyurt et al., 2023). Addressing distribution shifts in the test phase is also crucial for learning representations for various downstream environments. Therefore, researchers should consider future directions to develop distribution shift adaptations for universal representation learning with discrepancy-based and adversarial methods, for example.
7.3. Reliable Data Augmentation
As minor data augmentation can significantly impact time-series properties, determining the appropriate type and degree of augmentation is essential. Various techniques, including jittering, shifting, and warping, have been used for time-series representation learning. However, the reliability of these methods has not been fully explored. Data augmentation becomes more important with new learning paradigms like contrastive learning. It serves not only to expand the size of datasets but also to provide diverse class-invariant features. Unfortunately, current approaches focus more on the evolving role of data augmentation, often overlooking its fundamental task to maintain the integrity of the original data characteristics. Recent studies (Luo et al., 2023; Demirel and Holz, 2023) have focused on improving data augmentation reliability using innovative approaches and adaptive strategies based on more reliable criteria. Many studies still use empirical approaches to determine data augmentation, despite these advances. Therefore, methods for time-series data augmentation that estimates the reliability and efficacy of selecting the optimal augmentation strategy is promising.
7.4. Neural Architecture Search (NAS)
Making efficient deep models for universal representation learning is challenging because it laboriously relies on tedious manual trial-and-error to design the network architectures and select corresponding hyperparameters relevant to the three design elements we discussed. The inefficient hand-crafted design will inevitably involve human bias, leading to sub-optimal architectures (Zhang et al., 2023d). Accordingly, recent endeavors employ NAS to discover an optimal architecture by automatically designing neural architectures and network hyperparameters, e.g., # layers, network type, and embedding dimension. These configurations significantly affect the learned representation quality and the downstream performance. Even though NAS has demonstrated its successes in diverse tasks (White et al., 2023), NAS for time series is still underexplored. There are only NAS methods for a specific time-series analysis task, such as forecasting (Shah et al., 2021; Lai et al., 2023) and classification (Rakhshani et al., 2020; Ren et al., 2022), thus having very limited generalizability. NAS for universal time-series representation learning that perform well across downstream tasks is an important future direction, especially for industry-scale time series having high dimensions and large volumes newly generated every day.
7.5. Large Language and Foundation Models
Nowadays, many large language models (LLMs) across various domains have significantly transformed the landscape of natural language processing and computer vision. Integrating LLMs into time-series representation learning will enable the model to capture rich meanings embedded in these time-dependent patterns. The synergy between linguistic context and temporal dependencies additionally allows the model to identify complex relationships within the data, enabling more robust and fine-grained representations. The applications of LLMs, such as zero-shot learning, few-shot learning, and fine-tuning, have demonstrated performance comparable to existing deep learning methods. We note ongoing attempts to leverage LLMs for time-series analysis, yet mostly limited to forecasting tasks (Jin et al., 2023b). Thus, using LLMs in time-series representation learning is expected to enhance the embedding quality by accurately capturing time-dependent patterns in the time series. We also expect future research on aligning time-series representations with language embeddings to provide valuable insights, not only in the context of single-modal time series but also in the realm of multi-modal or multivariate time series.
7.6. Representations of Irregularly-Sampled Time Series (ISTS)
Irregular time series naturally appear in real-world scenarios across various fields, including finance, healthcare, and environmental observation. This irregularity usually stems from sporadic observations and system failures. However, managing ISTS effectively has proven challenging because existing deep learning models are devised initially for regularly-spaced natural language data. Early attempts like GRU-D (Che et al., 2018) modify existing deep models to handle ISTS. Another line of work uses Neural ODEs (Chen et al., 2018) that learn continuous vector space to represent ISTS. While the concept of NDEs aligns well with ISTS, these methods still have limitations. First, the training time of NDEs may be instable due to the use of the Runge-Kutta method for integration calculations. Second, the requirement that the hidden state of NDEs should be differentiable limits their ability to represent high-frequency time series. Compared to NDEs, approaches modifying attention (Shukla and Marlin, 2021) or employing interpolation (Shukla and Marlin, 2019) to address ISTS also show promising results.
Besides the model-centric approaches, we expect a potential research direction for ISTS to be a data-centric approach. Unlike regular time series, ISTS can contain diverse causes of irregularities often overlooked by previous studies. For instance, medical data may exhibit irregularities from sensor malfunctions, while financial data (e.g., credit card transactions) may contain irregularities due to its sporadic event-based nature. By integrating the understanding of the causes of irregularities into the learning process, we can obtain more precise representations of ISTS.
7.7. Multi-Modal and Multi-View Representation Learning
Vision-language models such as CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021) have recently shown remarkable performance in zero-shot learning and fine-tuning for various vision-related downstream tasks by utilizing the semantics of human languages. If time-series data entail semantics understandable to humans, we can make multi-view representations of time series by using human languages as an additional modality to annotate them. As a result, the learned representations from time series-language representation learning will become more expressive and fine-grained in semantics. For example, human activity data (Chen et al., 2021c) or time-series anomalies can be described in human languages based on human movement and the context of abnormal situations. However, unlike video representation learning, which can leverage a vision-language pre-trained model for the annotation, time-series data annotation with human language can be laborious and highly domain-specific. Thus, building a large multi-modal time series-text dataset will be a promising future direction.
Table 3. Summary of public datasets widely used for time-series representation learning. $T$ and $V$ indicate the varying time-series length (or # frames) and number of variables (or video resolution) per sample, respectively.
Downstream Tasks Dataset Name Size Dimension Domain Modality Reference Source
Forecasting & Imputation ETTh 14,307 7 Electric Power time series (Zhou et al., 2021)
ETTm 57,507 7 Electric Power time series (Zhou et al., 2021)
Electricity 26,304 321 Electricity Consumption time series (Wu et al., 2023)
Traffic 17,451 862 Transportation time series (Wu et al., 2023)
PEMS-BAY 16,937,179 325 Transportation spatiotemporal (Li et al., 2018)
METR-LA 6,519,002 207 Transportation spatiotemporal (Li et al., 2018)
Weather 52,603 21 Climatological Data time series (Wu et al., 2023)
Exchange 7,588 8 Daily Exchange Rate time series (Lai et al., 2018)
ILI 861 7 Illness time series (Wu et al., 2023)
Google Stock 3,773 6 Stock Prices time series (Yoon et al., 2019)
Monash TSF $30\times T$ $V$ Multiple time series (Godahewa et al., 2021)
LOTSA $105\times T$ $V$ Multiple time series (Woo et al., 2024)
Solar 52,560 137 Solar Power Production time series (Lai et al., 2018)
MoJoCo 10,000 x 100 14 Control Tasks time series (Jhin et al., 2022)
USHCN-Climate 386,068 5 Climatological Data time series (Schirmer et al., 2022)
Classification & Clustering UCR $128\times T$ 1 Multiple time series (Dau et al., 2019)
UEA $30\times T$ $V$ Multiple time series (Bagnall et al., 2018)
PhysioNet Sepsis $40,336\times T$ 34 Medical Data time series (Reyna et al., 2020)
PhysioNet ICU $12,000\times T$ 36 Medical Data time series (Silva et al., 2012)
PhysioNet ECG $12,186\times T$ 1 Medical Data time series (Clifford et al., 2017)
HAR 10,299 9 Human Activity time series (Anguita et al., 2013)
EMG 163 1 Medical Data time series (PhysioBank, 2000)
Epilepsy 11,500 1 Brain Activity time series (Andrzejak et al., 2001)
Waveform 76,567 2 Medical Data time series (Zhang et al., 2023f)
Gesture 440 3 Hand Gestures time series (Liu et al., 2009)
MOD 39,609 2 Moving Object time series (Liu et al., 2023)
PAMAP2 9,611 10 Human Activity time series (Liu et al., 2023)
Sleep-EEG 371,005 1 Sleep Stages time series (Dong et al., 2023)
RealWorld-HAR 12,887 9 Human Activity time series (Liu et al., 2023)
Speech Commands 5,630,975 20 Spoken Words audio (Jhin et al., 2022)
LRW 13,050,000 $64\times 64$ Lib Reading video (Chung and Zisserman, 2017)
ESC50 10,000 1 Environmental Sound audio (Piczak, 2015)
UCF101 333,000 $320\times 240$ Human Activity video (Soomro et al., 2012)
HMDB51 $6,849\times T$ $V_{width}\times V_{height}$ Human Activity video (Kuehne et al., 2011)
Kinetics-400 7,656,125 $V_{width}\times V_{height}$ Human Activity video (Kay et al., 2017)
AD 1,527,552 16 Medical Data time series (Wang et al., 2023a)
PTB 18,711,000 15 Medical Data time series (Wang et al., 2023a)
TDBrain 3,035,136 33 Brain Activity time series (Wang et al., 2023a)
MMAct
36,764
(only number of instances)
N/A Human Activity multi-modality (Kong et al., 2019)
PennAction $2,326\times T$ $640\times 480$ Human Activity video (Chen et al., 2022)
FineGym $4,883\times T$ $V_{width}\times V_{height}$ Human Activity video (Chen et al., 2022)
Pouring $84\times T$ $V_{width}\times V_{height}$ Human Activity video (Chen et al., 2022)
Something-Something 2,650,164 $84\times 84$ Human Activity video (Kim et al., 2023c)
Regression TSER Archive $19\times T$ $V$ Multiple time series (Tan et al., 2021)
Neonate $79\times T$ 18 Neonatal EEG Recordings time series (Stevenson et al., 2019)
IEEE SPC $22\times T$ 5 Heart Rate Monitoring time series (Demirel and Holz, 2023)
DaLia $15\times T$ 11 Heart Rate Monitoring time series (Demirel and Holz, 2023)
IHEPC 2,075,259 1 Electricity Consumption time series (Franceschi et al., 2019)
AEPD 19,735 29 Appliances Energy time series (Zhang et al., 2023a)
BMAD 420,768 6 Air Quality time series (Zhang et al., 2023a)
SML2010 4,137 18 Smart home time series (Zhang et al., 2023a)
Segmentation TSSB $75\times T$ 1 Multiple time series (Ermshaus et al., 2023)
UTSA $32\times T$ 1 Multiple time series (Gharghabi et al., 2017)
Anomaly Detection FD-A 8,184 1 Mechanical System time series (Lessmeier et al., 2016)
FD-B 13,640 1 Mechanical System time series (Lessmeier et al., 2016)
KPI 5,922,913 1 Server Machine time series (Yue et al., 2022)
TODS $T$ $V$ Synthetic Data time series (Lai et al., 2021)
SMD 1,416,825 38 Server Machine time series (Su et al., 2019)
ASD 154,171 19 Server Machine time series (Li et al., 2021)
PSM 220,322 26 Server Machine time series (Abdulaal et al., 2021)
MSL 130,046 55 Spacecraft time series (Hundman et al., 2018)
SMAP 562,800 25 Spacecraft time series (Hundman et al., 2018)
SWaT 944,919 51 Infrastructure time series (Mathur and Tippenhauer, 2016)
WADI 1,221,372 103 Infrastructure time series (Ahmed et al., 2017)
Yahoo 572,966 1 Multiple time series (Yue et al., 2022)
TimeSeAD $21\times T$ $V$ Multiple time series (Wagner et al., 2023)
TSB-UAD $1,980\times T$ 1 Multiple time series (Paparrizos et al., 2022b)
UCR-TSAD $250\times T$ 1 Multiple time series (Wu and Keogh, 2021)
UCFCrime $1,900\times T$ $V_{width}\times V_{height}$ Surveillance video (Sultani et al., 2018)
Oops! $20,338\times T$ $V_{width}\times V_{height}$ Human Activity video (Epstein et al., 2020)
DFDC $128,154\times T$ $256\times 256$ Deepfake video (Dolhansky et al., 2020)
Retrieval EK-100 89,977 $V_{width}\times V_{height}$ Human Activity video (Damen et al., 2022)
HowTo100M 136,600,000 $V_{width}\times V_{height}$ Human Activity video (Miech et al., 2019)
MUSIC 714 $V_{width}\times V_{height}$ Musical Instrument multi-modality (Zhao et al., 2018)
8. Conclusions
This article introduces a universal time-series representation learning research and its importance for downstream time-series analysis. We present a comprehensive and up-to-date literature review of universal representation learning for time series by categorizing the recent advancements from design perspectives. Our main goal is to answer how each fundamental design element—training data, neural architectures, and learning objectives—of state-of-the-art time-series representation learning methods contributes to the improvement of the learned representation quality, resulting in a novel structured taxonomy with 26 subcategories. Although most state-of-the-art studies consider all design elements in their methods, only one or two elements are newly proposed. Given the current review of the selected studies, we find that decomposition, transformation, and sample selection methods in the data-centric approaches are still underexplored. In addition, we provide a practical guideline about standard experimental setups and widely used time-series datasets for particular downstream tasks, together with discussions on various open challenges and future research directions related to time-series representation learning. Ultimately, we expect this survey to be a valuable resource for practitioners and researchers interested in a multi-faceted understanding of the universal representation learning methods for time series.
Acknowledgements.
This work was supported by Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. 2022-0-00157, Robust, Fair, Extensible Data-Centric Continual Learning, 50% and No. 2020-0-00862, DB4DL: High-Usability and Performance In-Memory Distributed DBMS for Deep Learning, 50%).